### v2

# Inference on many jumps in nonparametric panel regression models

Abstract
We investigate the significance of change-point or jump effects within fully nonparametric regression contexts, with a particular focus on panel data scenarios where data generation processes vary across individual or group units, and error terms may display complex dependency structures. In our setting the threshold effect depends on a specific covariate, and we permit the true nonparametric regression to vary based on additional latent variables. We propose two uniform testing procedures: one to assess the existence of change-point effects and another to evaluate the uniformity of such effects across units. Even though the underlying data generation processes are neither independent nor identically distributed, our approach involves deriving a straightforward analytical expression to approximate the variance-covariance structure of change-point effects under general dependency conditions. Notably, when Gaussian approximations are made to these test statistics, the intricate dependency structures within the data can be safely disregarded owing to the localized nature of the statistics. This finding bears significant implications for obtaining critical values. Through extensive simulations, we demonstrate that our tests exhibit excellent control over size and reasonable power performance in finite samples, irrespective of strong cross-sectional and weak serial dependency within the data. Furthermore, applying our tests to two datasets reveals the existence of significant nonsmooth effects in both cases.
Keywords: high-dimensional time series; temporal and cross-sectional dependence; changepoint analysis; Gaussian approximation; simultaneous tests.
1 Introduction
Recently, there has been a notable surge in the study and application of changepoint analysis in high-dimensional time series data (see, e.g., Cho and Fryzlewicz, (2015), Cao et al., (2018), Chen et al., (2022), and Li et al., (2024), along with the references cited therein). The focus of changepoint analysis has shifted from the traditional study of temporal changepoints to exploring changepoints within nonlinear regression models. Detecting changepoints in regression models is akin to clustering data points based on similar characteristics or behaviors. Chang et al., (2015) reviews several intriguing applications of changepoint regression. In the medical domain, changepoint analyses often indicate shifts in the impact of certain risk factors on outcomes (see, e.g., Pastor and Guallar, (1998) and Muggeo, (2003)). Additionally, there is significant interest in detecting and evaluating changepoint regressions within econometrics. For example, Lee, (2008) studies the effect of party incumbency in U.S. House Elections. The purpose of these studies is to estimate not only where the threshold effect exists but also the magnitude of the effects.
There is extensive literature on parametric high-dimensional changepoint regression analysis. Wang et al., (2021), Cho et al., (2024), and Rinaldo et al., (2021) investigate changepoint analysis of coefficients in high-dimensional linear regression models using penalization methods. Lee et al., (2016) examines a high-dimensional regression model with a changepoint due to a covariate threshold in a cross-sectional study. These studies primarily focus on parametric models with high-dimensional covariates. Similar to Lee et al., (2016), this paper also considers changepoint regression due to a covariate threshold. However, our focus is on nonparametric regression with a high-dimensional changepoint effect driven by heterogeneity in the panel data structure. Specifically, we investigate the heterogeneous nonparametric panel data with jumps in the conditional mean functions:
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt},$ (1.1)
$\mathbb{E}(e_{jt}|X_{jt})=0,$ (1.2)
where $Y_{jt}$ and $X_{jt}$ are the outcome and covariate variables for entity $j\in[N]:=\{1,\ldots,N\}$ at time point $t\in[T],$ $f_{j}(\cdot)$ is a continuous function, $c_{0j}$ is a threshold value that is either known or unknown (if $\gamma_{j}=0$, then without loss of generality set $c_{0j}=0$ for the unknown case), $\gamma_{j}$ is the jump effect at $c_{0j}$ and $e_{jt}$ is the error term. When $c_{0j}$’s are known, the above model is linked to a panel version of the regression discontinuity model that is widely studied in an i.i.d. setup. Our primary focus is to test whether changepoints exist and if the jump effects are consistent across different entities under general assumptions about $e_{jt}$. Specifically, we consider $e_{jt}$ as a reduced-form error term with a complex structure that accounts for latent variables $U_{jt}$, which can be arbitrarily correlated with $X_{jt}$. The detailed settings for $e_{jt}$ are discussed in Section 2.1.
Our paper is related to four strands of literature in statistics and econometrics. First, we add to the extensive body of work on changepoint analysis and threshold regression models. Given the vast literature in this area, we only review briefly cases involving high-dimensional data. For example, Li et al., (2016) consider estimation of panel data models with interactive fixed effects (IFEs) and multiple structure changes, and Manner et al., (2019) consider testing for structural breaks in factor copula models. Miao et al., 2020a and Miao et al., 2020b consider the estimation of a panel threshold regression model with IFEs and latent group structures, respectively; Massacci and Trapani, (2022) study inference for high-dimensional threshold regressions with common stochastic trends; Barassi et al., (2023) study a heterogeneous panel threshold regression model with IFEs. Second, our paper is related to the literature on testing homogeneity in a panel data setting. For example, Phillips and Sul, (2003), Pesaran and Yamagata, (2008) and Su and Chen, (2013) propose tests for slope coefficient heterogeneity in a linear regression model with or without cross-sectional dependence. In the presence of slope heterogeneity, Su et al., (2016) and Su and Ju, (2018) model the slope heterogeneity via latent group structures. Third, in terms of methodology, our paper aligns with research on nonparametric trends in multivariate or high-dimensional time series. Inference on smooth trend functions in nonparametric models for time series data is a well-established field. For example, Zhang and Wu, (2012) study inference of time-varying regression models, Karmakar et al., (2022) study simultaneous inference for time-varying models, and Gao et al., (2024) consider estimation and inference for time-varying multivariate causal processes. Extending the above works to high-dimensional settings, Chen and Wu, (2018) consider statistical inference for trends in high-dimensional time series. Finally, our paper is connected to the extensive literature on regression discontinuity design (RDD). Our approach can be extended to a panel RDD framework with specific design assumptions, particularly addressing individual effects that may cause biased RDD estimators or result in efficiency loss when estimating homogeneous treatment effects in finite samples. Different from RDD, our work focuses on inference of jumps in the nonparametric regression functions for high-dimensional data models and is thus also related to a set of literature regarding inference of jumps in nonparametric trends. However, the existing literature concentrates only on the cross-sectional case without dependency; see, e.g., Qiu, (1998), Müller and Stadtmüller, (1999) and Spokoiny, (1998) who focus on the estimation of discontinuous regression functions in the one-dimensional case based on cross-sectional data. Another distinct feature is that we explicitly handle the case with missing covariates or models with misspecified variables. Theoretically, validating such an inference procedure requires working with a reduced-form model that has specific variance-covariance structures. This aspect of theoretical validity has not been explored in the time-varying trend literature. Establishing the validity of uniform inference is a major theoretical contribution that sets our work apart in the study of varying trends.
The methodological contributions are twofold. First, we develop algorithms to test the overall significance or heterogeneity of threshold effects and examine the asymptotic properties of the resulting test statistics. When $N$ is large, assuming homogeneous threshold effects $\gamma_{j}$ across all $j$ is restrictive, and inferences based on this assumption can be misleading if the threshold effects are heterogeneous. Therefore, it is important to test for heterogeneous threshold effects. Given the well-documented cross-sectional and serial dependence in panel data, it is crucial to develop tests that account for both types of dependence. Specifically, we develop a supremum-type test that can distinguish between a constant or zero average changepoint effect and the alternative that different individuals exist for different individuals/groups. The test is built upon the supremum of $N$ kernel-estimator-based test statistics, each of which is constructed from a time-series local linear regression. Due to the use of local smoothing, neither the weak dependence along the time dimension nor the weak or strong dependence along the individual/group dimension enters the asymptotic distribution of the final statistic. Interestingly, we find strong cross-sectional dependence is allowed in our error terms. As a practical recommendation, we find that under fairly weak assumptions, if there are clustering dependency patterns in $e_{jt}$, the correlation in the estimated statistics can be safely ignored, and only adjustments for heteroskedasticity are needed.
Second, we extend existing Gaussian approximation (GA) results to accommodate latent variables in the error terms. The theoretical challenges arise when the number of groups $N$ is much larger than $Tb_{j}$, the effective sample size, where $b_{j}$ represents a certain bandwidth parameter. This high dimensionality makes it difficult to apply the standard central limit theorem. Our theoretical results extend the existing GA results from Chernozhukov et al., (2015), Chernozhukov et al., 2017a , Zhang and Wu, (2017) and Chernozhukov et al., (2019). Our test setup is similar to that in Chernozhukov et al., (2019), but the main difference is that our test statistic is nonparametric, involving the differences between two kernel estimators. The primary technical challenge is accounting for underlying latent variables in our data generating process (DGP). To address this, we introduce conditioning arguments to derive new high-dimensional GA results that accommodate general spatial and temporal dependency structures in the error terms. These GA results are crucial for determining the critical values of our supremum test.
On the empirical side, we employ our methods in two empirical applications. First, we detect jumps in the news impact curve for constituents in the S&P 500 stock index with an unknown threshold. In the second application, we study the effect of party incumbency as determined by the 50% threshold in the two-party vote share in the US House Elections. Our findings confirm the presence of a significant positive incumbency effect as studied by Lee, (2008), and we detect significant heterogeneity in the threshold effect across the US states.
The rest of the paper is organized as follows. Section 2 provides the model setup and the practical steps of the testing procedure. Section 3 presents the assumptions and delivers the main theorems. In Section 4 we provide some Monte Carlo simulation results. Section 5 concludes. The Online Supplement contains the proofs of all results, some additional simulation results, and empirical applications.
Notation. For a vector $v=(v_{1},\ldots,v_{d})\in\mathbb{R}^{d}$ and $q>0$, we denote $|v|_{q}=(\sum_{i=1}^{d}|v_{i}|^{q})^{1/q}$ and $|v|_{\infty}=\max_{1\leq i\leq d}|v_{i}|$. For a matrix $A=(a_{i,j})_{1\leq i\leq m,1\leq j\leq n}$, we define the max norm $|A|_{\text{max}}=\max_{i,j}|a_{i,j}|$. For $s>0$ and a random vector $X$, we say $X\in\mathcal{L}^{s}$ if $\lVert X\rVert_{s}=[\mathbb{E}(|X|^{s})]^{1/s}<\infty$. For two positive number sequences $(a_{n})$ and $(b_{n})$, we say $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$ (resp. $a_{n}\asymp b_{n}$) if there exists $C>0$ such that $a_{n}/b_{n}\leq C$ (resp. $1/C\leq a_{n}/b_{n}\leq C$) for all large $n$, and say $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\rightarrow 0$ as $n\rightarrow\infty$. We set $(X_{n})$ and $(Y_{n})$ to be two sequences of random variables. Write $X_{n}=O_{\mathbb{P}}(Y_{n})$ if for any $\epsilon>0$, there exists $C>0$ such that $\mathbb{P}(X_{n}/Y_{n}\leq C)>1-\epsilon$ for all large $n$, and say $X_{n}=o_{\mathbb{P}}(Y_{n})$ if $X_{n}/Y_{n}\rightarrow 0$ in probability as $n\rightarrow\infty$. We denote the centered random variable $X$ by $\mathbb{E}_{0}(X)$, i.e., $\mathbb{E}_{0}(X)=X-\mathbb{E}(X)$.
2 Model Setup and Test Procedure
In this section, we present the model, the hypotheses and estimators, and the test procedure. Subsections 2.1 and 2.2 are concerned with the known threshold case. In Subsection 2.3 we focus on the case in which the threshold locations $c_{0j}$ are unknown. The algorithms for our testing procedures are listed in Subsection 2.4.
2.1 Model Setup
In this subsection, we formulate our model. The data $(Y_{jt},X_{jt})_{j\in[N],\ t\in[T]}$ are collected for individual/group $j$ at time $t$, where $Y_{jt}$ is the outcome variable and $X_{jt}$ is a continuous covariate variable. We work with the following full model:
$Y_{jt}=h_{j}(X_{jt},U_{jt})+\tau_{j}(X_{jt},U_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt},$ (2.1)
where $U_{jt}$ is a latent random variable that can be arbitrarily correlated with $X_{jt},$ $\epsilon_{jt}$ is a mean zero term that is independent of $\left(X_{jt},U_{jt}\right)$. Further, $h_{j}(\cdot,\cdot)$, $\tau_{j}(\cdot,\cdot)$ and $\sigma_{j}(\cdot,\cdot)$ are continuous functions of $X_{jt}$ and $U_{jt}$ for the baseline effect, the jump effect and the volatility, respectively. The motivation for considering such a model with a latent variable $U_{jt}$ is that, in practice, the outcome is often affected by unobserved variables and factors. For example, in our application (Section E.1 in the online supplement), we analyze the threshold effect of asset returns. $Y_{jt}$ is the stock volatility, $X_{jt}$ is the lagged stock return, and $U_{jt}$ are unobserved risk factors. Modeling $U_{jt}$ is important in this circumstance as the volatility of a single stock is most likely affected by other risk factors such as news events or unobserved market factors.
This paper aims to explore methods for conducting simultaneous inference on threshold effect estimators in a large $N$ and large $T$ setup. In particular, we allow $N\gg T.$ Our primary focus is on the threshold effect, starting with the case where the threshold value $c_{0j}$ is known and then extending the analysis to the unknown case in Subsections 2.3 and 3.4.
Assuming that $X_{jt}$ is a continuous random variable, the jump effect can be identified as a “gap” in the conditional expectations of the dependent variable at the threshold:
$\gamma_{j}=\lim_{x_{j}\downarrow c_{0j}}\mathbb{E}(Y_{jt}|X_{jt}=x_{j})-\lim_{x_{j}\uparrow c_{0j}}\mathbb{E}(Y_{jt}|X_{jt}=x_{j}).$ (2.2)
Let $\tilde{\tau}_{j}(X_{jt})=\mathbb{E}[\tau_{j}(X_{jt},U_{jt})|X_{jt}]$ and $\tilde{h}_{j}(X_{jt})=\mathbb{E[}h_{j}(X_{jt},U_{jt})|X_{jt}]$. Then the conditional expectation of the observed outcome given $X_{jt}$ is $\mathbb{E}(Y_{jt}|X_{jt})=\tilde{h}_{j}(X_{jt})+\tilde{\tau}_{j}(X_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}.$  Note that although model (2.1) takes a complicated structure it can be rewritten into model (1.1) after some simple transformation. Specifically denote $\gamma_{j}=\tilde{\tau}_{j}(c_{0j}),$
$f_{j}(X_{jt})=\tilde{h}_{j}(X_{jt})+[\tilde{\tau}_{j}(X_{jt})-\tilde{\tau}_{j}(c_{0j})]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$\varepsilon_{j}(X_{jt},U_{jt})=h_{j}(X_{jt},U_{jt})-\tilde{h}_{j}(X_{jt})+[\tau_{j}(X_{jt},U_{jt})-\tilde{\tau}_{j}(X_{jt})]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}\text{ and}$
$e_{jt}=\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}.$ (2.3)
Therefore model (2.1) is our full model and can be reduced to the working model in Equation (1.1). The only technical difficulty this introduces to our framework is the need to handle the special variance structure of the error term. It is easy to verify that $\mathbb{E}[\varepsilon_{j}(X_{jt},U_{jt})|X_{jt}]=0$ which further leads to $\mathbb{E}[e_{jt}|X_{jt}]=0.$ Note that $f_{j}\left(\cdot\right)$ is continuous provided $\tilde{h}_{j}\left(\cdot\right)$ and $\tilde{\tau}_{j}(\cdot)$ are. So $f_{j}(X_{jt})$ and $\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$ correspond to the continuous and discrete part of $\mathbb{E}(Y_{jt}|X_{jt}),$ respectively. The noise term $e_{jt}$ has a complex structure, influenced both by an unobserved random variable and by the heterogeneous volatility of the observations. This intricate structure, combined with the temporal and cross-sectional dependencies of the noise, makes the inference process particularly challenging.
Remark 1.
(Additive time and individual effects) Our model can account for both individual fixed effects $\alpha_{j}$ and time fixed effects $\nu_{t}$. Specifically, the working model (1.1) becomes
$Y_{jt}=\alpha_{j}+\nu_{t}+f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}.$ (2.4)
If $\mathbb{E}(\nu_{t}|X_{jt})=d(X_{jt})$ for some function $d(\cdot)$ under a covariance-stationarity condition along the $t$-dimension, then by defining $\tilde{e}_{jt}=e_{jt}+[\nu_{t}-\mathbb{E}\left(\nu_{t}|X_{jt}\right)]$ and $\tilde{f}_{j}(X_{jt})=\alpha_{j}+\mathbb{E}\left(\nu_{t}|X_{jt}\right)+f_{j}(X_{jt}),$ the above model can be reformulated as follows:
$Y_{jt}=\tilde{f}_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\tilde{e}_{jt}.$ (2.5)
In the absence of covariance-stationarity, the estimation would become much more involved and we have to leave it for future research.            $\Box$
Remark 2.
(Adding more covariates to fit the threshold model) Suppose that we want to include more covariates $Z_{jt}\in\mathbb{R}^{d}$ and $\theta\in\mathbb{R}^{d}$, we can impose a partial linear structure. Then the main model in (1.1) can be of the following form:
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}+Z_{jt}^{\top}\theta.$ (2.6)
Our methodology can be extended to the model above by adapting the approach used in Li et al., (2022).            $\Box$
2.2 Hypotheses and Estimators
In this section, we introduce the hypotheses and estimators. We would like to conduct some simultaneous inferences on $\gamma_{j}$ for $j\in\left[N\right]$. The importance of studying individual/group-specific threshold effects is discussed in Zimmert and Lechner, (2019).
First, we are interested in testing the existence of the threshold effects. The null and alternative hypotheses are defined as follows:
$H_{0}^{\left(1\right)}:\gamma_{j}=0\ \forall\text{ }j\in\left[N\right],\quad\text{and }H_{a}^{\left(1\right)}:\gamma_{j}\neq 0\ \text{for some }j\in\left[N\right].$ (2.7)
Additionally, we aim to test the homogeneity of the threshold effects. In this case, the null and alternative hypotheses are:
$H_{0}^{\left(2\right)}:\gamma_{j}=\gamma_{0}\ \text{for some }\gamma_{0}\text{ }\forall\text{ }j\in\left[N\right],\quad$
$\text{and}\quad H_{a}^{\left(2\right)}:\text{there is no constant}\gamma_{0}\text{ such that }\gamma_{j}=\gamma_{0}\text{ }\forall\ j\in\left[N\right].$
Rejecting of $H_{0}^{\left(2\right)}$ suggests the presence of heterogeneous threshold effects, in which case it is useful to extract some meaningful information from the consistent estimates $\hat{\gamma}_{j}$’s of $\gamma_{j}$’s. For example, the quantiles of $\hat{\gamma}_{j}$’s give us information on how the threshold effects spread among individuals. Providing simultaneous confidence intervals for the threshold effects is also tempting. If we attempt to study the asymptotics of $\hat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{N})^{\top}$, the standard central limit theorem may fail due to the high dimensionality $N$ when $N\rightarrow\infty$. One of the key theoretical contributions of the paper is to provide a general framework for making uniform inferences on heterogeneous threshold effects.
Let $\mu_{j}(x):=\mathbb{E}(Y_{jt}|X_{jt}=x)=f_{j}(x)+\gamma_{j}\mathbf{1}_{\{x\geq c_{0j}\}}.$ Then the threshold effect for individual/group $j$ (denoted as $\gamma_{j}$) can be identified as the difference between the left and right limits of $\mu_{j}$ at point $c_{0j}:$ $\gamma_{j}=\mu_{j}(c_{0j}+)-\mu_{j}(c_{0j}-),$ where $\mu_{j}(c_{0j}-)=\lim_{x\uparrow c_{0j}}\mu_{j}(x)$ and $\mu_{j}(c_{0j}+)=\lim_{x\downarrow c_{0j}}\mu_{j}(x).$ To estimate $\gamma_{j}$, we need to estimate both $\mu_{j}(c_{0j}+)$ and $\mu_{j}(c_{0j}-).$ To this aim, it is natural to adopt local linear estimators. Let $\partial_{+}\mu_{j}(\cdot)$ (resp. $\partial_{-}\mu_{j}(\cdot)$) be the right (resp. left) derivative of function $\mu_{j}(\cdot)$ and denote the estimators:
$\big{(}\hat{\mu}_{j}(c_{0j}+),{\partial_{+}\hat{\mu}_{j}}(c_{0j})\big{)}=$ $\mathrm{argmin}_{\beta_{j,0},\beta_{j,1}}\sum_{t=1}^{T}(Y_{jt}-\beta_{j,0}-\beta_{j,1}X_{jt})^{2}K_{jt}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}},$ (2.8)
$\big{(}\hat{\mu}_{j}(c_{0j}-),{\partial_{-}\hat{\mu}_{j}}(c_{0j})\big{)}=$ $\mathrm{argmin}_{\beta_{j,0},\beta_{j,1}}\sum_{t=1}^{T}(Y_{jt}-\beta_{j,0}-\beta_{j,1}X_{jt})^{2}K_{jt}\mathbf{1}_{\{X_{jt}<c_{0j}\}},$ (2.9)
where $K_{jt}:=K((X_{jt}-c_{0j})/b_{j}),$ $b_{j}$ is a bandwidth parameter, and $K(\cdot)$ is a kernel function. For $l=0,1,2$, define the weights as
$w_{j,b}^{+}(u,v)$ $=\frac{K((v-u)/b_{j})[S_{j2,b}^{+}(u)-(v-u)S_{j1,b}^{+}(u)]}{S_{j2}^{+}(u)S_{j0}^{+}(u)-S_{j1,b}^{+2}(u)}\mathbf{1}_{\{v\geq u\}},$
$S_{jl,b}^{+}(u)$ $=\sum_{t=1}^{T}(X_{jt}-u)^{l}K((X_{jt}-u)/b_{j})\mathbf{1}_{\{X_{jt}\geq u\}},$ (2.10)
and $w_{j,b}^{-}(u,v)$ (resp. $S_{jl,b}^{-}(u)$) is the same as $w_{j,b}^{+}(u,v)$ (resp. $S_{jl,b}^{+}(u)$) with $+$ and $\geq$ therein replaced by -- and <,<,< , respectively. We can thus evaluate the threshold effect near the cut-off point and define our local linear estimator of the average threshold effect for individual/group $j$ as
$\hat{\gamma}_{j}=\hat{\mu}_{j}(c_{0j}+)-\hat{\mu}_{j}(c_{0j}-)=\sum_{t=1}^{T}w_{jt,b}^{+}Y_{jt}-\sum_{t=1}^{T}w_{jt,b}^{-}Y_{jt},$ (2.11)
where the weights take the following form $w_{jt,b}^{+}=w_{j,b}^{+}(c_{0j},X_{jt})$ and $w_{jt,b}^{-}=w_{j,b}^{-}(c_{0j},X_{jt})$.
Under $H_{0}^{\left(1\right)}:$ $\gamma_{j}=0$ for all $j\in\left[N\right]$, we are exclusively testing for the overall significance of threshold effects. Since the function $f_{j}(\cdot)$ is continuous, we would expect that
$\sum_{t=1}^{T}w_{jt,b}^{+}f_{j}(X_{jt})\approx f_{j}(c_{0j}+)=f_{j}(c_{0j}-)\approx\sum_{t=1}^{T}w_{jt,b}^{-}f_{j}(X_{jt}).$
Hence by (1.1), the difference between our estimator and the parameter can be approximately expressed in terms of a weighted average of the error sequence,
$\hat{\gamma}_{j}\approx\gamma_{j}+\sum_{t=1}^{T}(w_{jt,b}^{+}e_{jt}-w_{jt,b}^{-}e_{jt}),$ (2.12)
where recall that $e_{jt}=\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}$ denotes the composite error in (1.1). Under $H_{0}^{\left(1\right)}$, one would expect $\left|\hat{\gamma}_{j}\right|$ to be small for all $j\in\left[N\right]$. This motivates us to consider the test statistic
$\mathcal{I}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}/v_{j}|,$ (2.13)
where
$v_{j}^{2}=v_{j}^{2}(c_{0j})=(Tb_{j})\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}$ (2.14)
is the conditional variance of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$ for standardization purposes. For now, we assume $v_{j}^{2}$ is given, in general, it is unknown and has to be replaced by its estimate. We discuss the consistent estimation of $v_{j}^{2}$ in Subsection 2.4. Under the assumptions in Section 3, it can be shown that all asymptotic results based on $v_{j}^{2}$ remain valid when it is replaced by the consistent estimate $\hat{v}_{j}^{2}$.
Similarly, under $H_{0}^{\left(2\right)}$, one would expect $\left|\hat{\gamma}_{j}-\overline{\hat{\gamma}}\right|$ to be small for all $j\in\left[N\right],$ where $\overline{\hat{\gamma}}=\frac{1}{N}\sum_{j=1}^{N}\hat{\gamma}_{j}$. This motivates us to consider the test statistic
$\mathcal{Q}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}-\overline{\hat{\gamma}}|/\tilde{v}_{j},$ (2.15)
where $\tilde{v}_{j}^{2}=(1-1/N)^{2}v_{j}^{2}+\sum_{i\neq j}^{N}v_{i}^{2}/N^{2}$. For practical implementation, we introduce a feasible version of $\mathcal{Q}$ in Algorithm 2 in Section C of the supplementary material.
2.3 The Unknown Threshold Case
It shall be noted that in case the breakpoints $c_{0j}$’s are not known, we shall develop an algorithm to estimate them. We therefore generalize our methodology to estimate unknown breakpoints that vary over ${\normalsize j}$. To this end, we search over a grid $[c_{1},c_{2},\ldots,c_{K}]$ for each individual $j\in[N]$. Let
$w_{jt,b}(c_{i})=w_{j,b}^{+}(c_{i},X_{jt})-w_{j,b}^{-}(c_{i},X_{jt})$ (2.16)
and
$\hat{\gamma}_{j}(c_{i})=\sum_{t=1}^{T}w_{jt,b}(c_{i})Y_{jt}.$ (2.17)
Further, let $v_{j}^{2}(c_{i})=(Tb_{j})\sum_{t=1}^{T}w_{jt,b}^{2}(c_{i})\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}$ and define
$\mathcal{I}^{C}=\max_{1\leq i\leq K}\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/v_{j}(c_{i})|.$ (2.18)
We propose to estimate $c_{0j}$ by $\hat{c}_{j}:=\arg\max_{1\leq i\leq K}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$for $j\in\left[N\right],$ where $\hat{v}_{j}(c_{i})$ is a consistent estimator of $v_{j}(c_{i})$ which will be introduced in Algorithm 3 in Section C of the supplementary material.
2.4 Test Procedure
In this subsection, we present the steps of the test procedures. To proceed, we begin by discussing the estimation of the error variance, $\sigma_{e,j}^{2},$ of $e_{jt}$ when $c_{0j}$’s are known.
  1. 1.
    Recall the definition of $\hat{\gamma}_{j}$ as in Equation (2.11). Regress $Y_{jt}-\hat{\gamma}_{j}$ on $X_{jt}$ using the local linear method with kernel $K\left(\cdot\right)$ and bandwidth $b_{j}$.
  2. 2.
    Calculate the regression residuals $\hat{e}_{jt}$.
  3. 3.
    Estimate $\sigma_{e,j}^{2}$ by
    $\hat{\sigma}_{e,j}^{2}=\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}\hat{e}_{jt}^{2}}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}}$ (2.19)
The estimator $\hat{\sigma}_{e,j}^{2}$ is used to form $\hat{v}_{j}^{2}$ in the Algorithms 1 and 2 in the case of known $c_{0j}$, which are applied to conduct simultaneous tests for $H_{0}^{\left(1\right)}$ and $H_{0}^{\left(2\right)}$. These algorithms are supported by the theoretical GA results in Section 3. The following Algorithm 1 is for testing $H_{0}^{(1)}$.
Algorithm 1 Testing for existence of jumps: i.e., $\gamma_{j}=0$ for all $j$.
1:for $j=1,\ldots,N$ do
2:     Estimate $\gamma_{j}$ for each group/individual $j$ using equation (2.11).
3:     Estimate the variance $v_{j}^{2}$ of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$ by $\hat{v}_{j}^{2}=Tb_{j}\sum_{t=1}^{T}\left(w_{jt,b}^{+}-w_{jt,b}^{-}\right)^{2}\hat{\sigma}_{e,j}^{2}$, where $\hat{\sigma}_{e,j}^{2}$ is a consistent estimator for the error variance using the observations close to the threshold $c_{0j}$ as in (2.19).
4:end for
5:Calculate the feasible test statistic ^$\hat{\mathcal{I}}:=\mbox{max}_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}|/\hat{v}_{j}$.
6:Reject the null hypothesis if ^$\hat{\mathcal{I}}>q_{\alpha}$, where $q_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}|Z_{j}|$ with $Z_{j}$’s being independent standard normal variables.
Algorithm 2 in Section C in the supplement is used for testing homogeneity $H_{0}^{(2)}$. For the case where the threshold is unknown, the corresponding algorithm is detailed in Algorithm 3 in the same Section.
[…]
From (1.1), we see that the reduced-form error $e_{jt}$ inherits the dependency structure in $\epsilon_{jt}$. However, this dependence does not affect the calculation of critical values, even in the case of fairly strong cross-sectional dependence (e.g., a factor structure). This is mainly due to the localized nature of our kernel-type estimators $\hat{\gamma}_{j}$’s in our test statistic. In fact, we can show that under some mild conditions, as long as the joint density of $X_{it}$ and $X_{jt}$ is not degenerate for any $\left(i,j\right)$ pair (see Assumption 3.6 in Section 3), the covariance terms between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ will be of lower order, $O(1/T)$, compared to the variance terms which are $O(1/(Tb))$. For a more detailed discussion on the key intuition on the impact of the dependence structure in a simplified setting, we refer to Section F of the Supplement materials. For a discussion on the issue of bandwidth selection, we refer to Remark 7 in the next section.
3 Main Theorems
In this section, we state the assumptions for the asymptotic analyses and then study the asymptotic properties of the test statistics. Sections 3.2 and 3.3 present theoretical results on the known threshold case, providing theoretical support for the procedures detailed in Algorithms 1 and 2. Section 3.4 addresses the case of unknown thresholds and examines the validity of Algorithm 3.
3.1 Assumptions
Recall $\mathcal{I}$ in (2.13), we aim to study its asymptotic properties. To do this, we require a time series version of a non-centered high-dimensional Gaussian approximation. To account for dependence along both the $j$ and $t$ dimensions, we need to impose certain assumptions.
As mentioned above, we will postulate a general spatial and temporal dependence assumption. As for the covariate variable $X_{jt}$ and the confounding variable $U_{jt}$, we assume that they take the form
$X_{jt}=H_{j}(e_{t},e_{t-1},\ldots)\quad\text{and}\quad U_{jt}=\tilde{H}_{j}(X_{jt},\tilde{e}_{t}),$ (3.1)
where $\left\{e_{t}\right\}_{t\in\mathbb{Z}}$ is a sequence of i.i.d. random innovations, $\left\{H_{j}\right\}_{j\in[N]}$ are functions so that $X_{jt}$ are well defined, the innovations $\left\{\tilde{e}_{t}\right\}_{t\in\mathbb{Z}}$ are i.i.d. and independent of $\left\{e_{t}\right\}_{t\in\mathbb{Z}}$, and $\{\tilde{H}_{j}\}_{j\in[N]}$ are measurable functions. Denote $\mathcal{F}_{t}=(e_{t},e_{t-1},\ldots)$ and ~$\tilde{\mathcal{F}}_{t}=(\tilde{e}_{t},\tilde{e}_{t-1},\ldots)$. For the innovation part, the dependence among $\epsilon_{t}:=(\epsilon_{1t},\ldots,\epsilon_{Nt})^{\top}$is allowed to be strong along the cross-sectional dimension (e.g., with a factor structure) and it has to be weak along the time dimension, as we only exploit the sample size in the time series dimension for the estimation of heterogeneous threshold effects. Specifically, we assume that $\epsilon_{t}$ follows an MA($\infty$) process as follows:
$\epsilon_{t}=\sum_{k\geq 0}A_{k}\eta_{t-k},$ (3.2)
where $\eta_{t}=(\eta_{1t},\ldots,\eta_{\tilde{N}t})^{\top}\in\mathbb{R}^{\tilde{N}},$ $t\in\mathbb{Z},$ are i.i.d. random vectors with zero mean and identity covariance matrix $I_{\tilde{N}}$, and $A_{k}\in\mathbb{R}^{N\times\tilde{N}}$ are real-valued matrices with $\tilde{N}\leq cN$ for some constant $c>0$. In particular, we allow $\tilde{N}=N$ so that the $A_{k}$’s become square matrices. We assume that $\left\{e_{t},\tilde{e}_{t}\right\}_{t\in\mathbb{Z}}$ are independent of $\left\{\eta_{t}\right\}_{t\in\mathbb{Z}}.$
Note that we allow for the dependence between $X_{jt}$ and $U_{jt}$. Equation (3.1) is a fairly general setup, where almost all kinds of dependence structures between $X_{jt}$ and $U_{jt}$ can be fulfilled for the following simple reasoning. For any two continuous random variables $X$ and $U$ with the conditional cumulative distribution function of $U$ given $X$ as $F_{U|X}(\cdot|X)$. Let $\tilde{e}=F_{U|X}(U|X)$. Then by Lemma F.1 in Appendix F in Rio, (2017), $\tilde{e}$ is a uniformly distributed random variable on $[0,1]$ and is independent of $X$. Moreover, $U=F_{U|X}^{-1}(\tilde{e}|X).$ It is worth noting that our model and assumption can be extended to multivariate latent variable $U_{jt}$ without introducing additional theoretical insights.
Let $g_{j}(\cdot)$ denote the density function of $X_{jt}.$ For $t\geq s+1,$ let $g_{j,t}(x|\mathcal{F}_{s}):=d\mathbb{P}(X_{jt}\leq x|\mathcal{F}_{s})/dx.$ For any constant $r>0$, we denote
$v_{j,r}(x)=\mathbb{E}(|\varepsilon_{j}(X_{jt},U_{jt})|^{r}|X_{jt}=x).$ (3.3)
We first impose some conditions on the elements of $\eta_{t}.$
Assumption 3.1.
(Moment) $\eta_{jt}$’s are i.i.d. with zero mean and unit variance across $j$ and $t$. Either one of the following two conditions is satisfied:
  1. (i)
    (Finite moments) For some constant $q>2,$ $\eta_{11}$ has finite $q$th moment.
  2. (ii)
    (Sub-exponential) For some constant $\lambda_{0}>0$, $\eta_{1}=(\eta_{11},\ldots,\eta_{\tilde{N}1})^{\top}$ is sub-exponential with $a_{0}:=\sup_{|v|_{2}\leq\lambda_{0}}\mathbb{E}(e^{|v^{\top}\eta_{1}|})<\infty.$
Remark 3.
Assumption 3.1 imposes moment conditions on the underlying noise terms. We allow for two possibilities, $(i)$ only some finite order of moments exist for $\eta_{11}$, and $(ii)$ $\eta_{1}$ is sub-exponential. We shall provide conclusions for different sets of moment assumptions.            $\Box$
The following assumption imposes some conditions on the temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}.$ Denote $\mathcal{C}_{j}$ as a set of search values for the time series $j$ containing at least the threshold values $c_{0j}$. If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in subsections 3.2 and 3.3. We will specify $\mathcal{C}_{j}$ later in the theorems for the unknown case.
Assumption 3.2.
(Dependence) (i) For any $x\in\mathcal{C}_{j},$ assume $g_{j,t}(x|\mathcal{F}_{t-1})$ has finite $p$th moment for some $p>2$. Denote
$\theta_{k,p}=\Big{\|}\sup_{1\leq j\leq N,1\leq t\leq T,x\in\mathcal{C}_{j}}\big{|}g_{j,t}(x|\mathcal{F}_{t-1})-g_{j,t}(x|\mathcal{F}_{t-1,\{t-1-k\}})\big{|}\Big{\|}_{p}.$
Assume $\sup_{m\geq 0}m^{\alpha}\sum_{k\geq m}\theta_{k,p}<\infty$ for some $\alpha>0.$
(ii) For some $C>0$ and $\beta>0,$ we have $\max_{1\leq j\leq N}\sum_{k\geq m}|A_{k,j,\cdot}|_{2}\leq C(m\vee 1)^{-\beta},\quad$where $m\geq 0$ and $A_{k,j,\cdot}\in\mathbb{R}^{\tilde{N}}$ is the $j$th row of matrix $A_{k}.$
Remark 4.
Assumption 3.2 $(i)$-$(ii)$ essentially assumes weakly temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}$ for any $j\in\left[N\right],$ respectively. No weak dependence conditions are imposed on the cross-sectional dimension. Therefore we can allow for strong cross-sectional dependence over the $j$ dimension. We shall also note that $\beta$ indicates the strength of the dependency structure of the error processes. In the case of $\beta\rightarrow\infty$, the error process becomes temporally independent.            $\Box$
Assumption 3.3.
(Boundedness) For each $j\in\left[N\right]$, $\varepsilon_{j}(\cdot,\cdot)$ is a bounded function.
Due to the construction of $\varepsilon_{j}(\cdot,\cdot)$ in (2.1), if functions $h_{j}(\cdot,\cdot)$ and $\tau_{j}(\cdot,\cdot)$ are both bounded, then the above assumption naturally holds. Let $\bar{b}=\max_{1\leq j\leq N}b_{j}$ and $\underline{b}=\min_{1\leq j\leq N}b_{j}.$ The next assumption imposes some smoothness conditions.
Assumption 3.4.
(Smoothness) (i) The density functions $g_{j}(\cdot)$ of $X_{jt}$ are lower bounded on $\mathcal{C}_{j},$ that is for any $x\in\mathcal{C}_{j},$ assume $\min_{1\leq j\leq N}g_{j}(x)\geq c$ for some constant $c>0$. The conditional densities $g_{j,t}(\cdot|\mathcal{F}_{s})$ satisfy that $\max_{j,t}|g_{j,t}(\cdot|\mathcal{F}_{t-1})|$ $=O(1)$ and that
$\max_{1\leq j\leq N,1\leq t\leq T}\sup_{x\in\mathcal{C}_{j}}\sup_{|y-x|\leq b_{j}}|g_{j,t}(y|\mathcal{F}_{t-1})-g_{j,t}(x|\mathcal{F}_{t-1})|=O(\bar{b}).$
(ii) For all $j\in[N],$ assume that for any $x\in\mathcal{C}_{j}$, function $f_{j}(\cdot)$ satisfies
$f_{j}(y)=\begin{cases}f_{j}(x)+\partial_{+}f_{j}(x)(y-x)+O((y-x)^{2})\quad&\text{if}\ y\in[x,x+b_{j}],\\ f_{j}(x)+\partial_{-}f_{j}(x)(y-x)+O((y-x)^{2})&\text{if}\ y\in[x-b_{j},x],\end{cases}$
where $\partial_{+}f_{j}(x)$ (resp. $\partial_{-}f_{j}(x)$) is the right (resp. left) derivative of function $f_{j}\left(\cdot\right)$ at point $x$, and the constants in $O(\cdot)$ are independent of $j,b_{j}$.
(iii) The standard deviation functions $\sigma_{j}(\cdot,\cdot)$ satisfy $\inf_{u\in\mathbb{R},x\in\mathcal{C}_{j}}\sigma_{j}(x,u)>0$ with
$\max_{1\leq j\leq N}\sup_{u\in\mathbb{R},x\in\mathcal{C}_{j},|y-x|\leq b_{j}}|\sigma_{j}(y,u)/\sigma_{j}(x,u)-1|=1+O(\bar{b}).$
Remark 5.
Assumption 3.4 $(i)$ imposes a lower bound condition for the density functions $g_{j}(\cdot)$. Condition $(ii)$ imposes some smoothness conditions of the mean function $f_{j}(\cdot).$ Note that we do not require $f_{j}(\cdot)$ to be differentiable at point $x$; it suffices for $f_{j}(\cdot)$ to have bounded left and right second derivatives near $x$. Assumption $(iii)$ is for the smoothness of the variance function $\sigma_{j}^{2}(\cdot,\cdot)$. If $\sigma_{j}(x,u)$ has bounded partial left and right derivatives with respect to $x$, for any $x\in\mathcal{C}_{j}$ uniformly over $u,$ then assumption $(iii)$ is fulfilled. Note that we do not assume the left and right derivatives of $f_{j}(\cdot)$ to be the same, meaning the slope around the threshold can differ. To be general, we do not impose any differentiability assumptions on $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$.            $\Box$
The following assumption imposes some regularity conditions on the kernel function and bandwidth parameter.
Assumption 3.5.
[…]
$1<\frac{K_{2}^{+}K_{0}^{+}}{K_{1}^{+2}},\frac{K_{2}^{-}K_{0}^{-}}{K_{1}^{-2}}\leq c$ (3.4)
for some constant $c>1.$
(ii) As $\left(N,T\right)\rightarrow\infty,$ assume $\bar{b}\rightarrow 0$ and $\underline{b}T/\mathrm{log}(NT)\rightarrow\infty.$ Additionally, assume that $b\asymp\underline{b}\asymp\bar{b}$.
Remark 6.
Assumption 3.5$(i)$-$(ii)$ imposes certain conditions on the kernel function $K\left(\cdot\right)$and bandwidth parameters $b_{j}$’s, respectively. If one chooses $b_{j}$ to be the same across $j,$ Assumption 3.5$(ii)$ reduces to the standard conditions on the bandwidth parameter in local linear kernel estimation with one-dimensional covariate.            $\Box$
Let $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1})$ be the conditional joint density of $X_{j_{1}t}$ and $X_{j_{2}t}$ given $\mathcal{F}_{t-1}.$ The following assumption imposes a condition on $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1}).$
Assumption 3.6.
(Joint density) There is no perfect or asymptotically perfect collinearity in $X_{j_{1}t}$ and $X_{j_{2}t}$ for any $j_{1}\neq j_{2}.$ The conditional densities $g_{j_{1},j_{2}}(x_{1},x_{2}|\mathcal{F}_{t-1})$ are uniformly upper bounded, that is, $\max_{1\leq j_{1},j_{2}\leq N}\sup_{x_{1},x_{2}\in\mathbb{R}}|g_{j_{1},j_{2}}(x_{1},x_{2}|\mathcal{F}_{t-1})|<\infty.$
3.2 Gaussian Approximation Results
If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in subsections 3.2 and 3.3. We first consider the GA result for $\mathcal{I}$ defined in (2.13). Define $d_{j}=\gamma_{j}/v_{j}$ and $\underline{d}=(d_{1},d_{2},\ldots,d_{N})^{\top}$. Further define $Z+\underline{d}$ as a Gaussian vector centered around $\underline{d}$, where $Z$ is a vector of $N$ standard normally distributed random variables, and $\underline{d}$ denotes a bias term which is a zero vector under the null $H_{0}^{\left(1\right)}$. The following theorem states that the infeasible statistic $\mathcal{I}$ can be approximated well by the maximum of Gaussian random variables. Denote
$\Delta=\mathrm{log}^{7/6}(NT)(\underline{b}T)^{-1/6}+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}+T^{-(\alpha p)\wedge(p/2-1)}+\bar{b}^{1/3}\mathrm{log}^{2/3}(N).$
Define $\mathcal{R}_{NT}=\Delta+(\underline{b}T^{1-2/q})^{-1/3}\mathrm{log}(NT)+\mathrm{log}(NT)N^{1/q}T^{-\beta}$ under Assumption 3.1$(i)$, and $\mathcal{R}_{NT}=\Delta+\mathrm{log}(NT)T^{-\beta}$ under Assumption 3.1$(ii)$.
Theorem 3.1 (Gaussian Approximation 1).
Let Assumptions 3.1-3.6 hold. Then we have
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}\leq u)-\mathbb{P}(|Z+\underline{d}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{NT}.$ (3.5)
Now we list the related rate requirement under different moment conditions. Under Assumption 3.1$(i)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $\underline{b}^{-1}T^{2/q-1}\mathrm{log}^{3}(NT)\rightarrow 0$ and $\mathrm{log}(NT)N^{1/q}T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$. Under Assumption 3.1$(ii)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $(T\underline{b})^{-1}\mathrm{log}^{7}(NT)\rightarrow 0,$ $\bar{b}\mathrm{log}^{3}(N)\rightarrow 0$ and $\mathrm{log}(NT)T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$.
Theorem 3.1 lays down the foundation for testing the null hypothesis $H_{0}^{\left(1\right)}.$ The proof of Theorem 3.1 is technically challenging and involved for two reasons. First, we allow for both temporal and cross-sectional dependence among $\left\{\left(X_{jt},\epsilon_{jt}\right)\right\}.$ Second, the latent variable $U_{jt}$ makes the error term $e_{jt}$ in (1.1) highly complex. Such complicated structure makes it impossible to apply any existing GA results or their proof strategies directly to prove the above result. In Section A.1 of the online supplement, we outline the main idea used in the proof of the above theorem, highlighting that the key step is proving the GA result in Lemma A.1. We note that the term $\mathrm{log}^{7/6}(NT)(\underline{b}T)^{-1/6}+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}$ arises from extending the original GA results in Chernozhukov et al., 2017a . The term $\mathrm{log}(NT)N^{1/q}T^{-\beta}$ is due to the temporal dependence and the moment condition of the innovation term $\eta_{jt}$. The term $\bar{b}^{1/3}\mathrm{log}^{2/3}(N)$ results from the covariance approximation and Gaussian comparison.
Next, we consider the GA result for $\mathcal{Q}$ defined in (2.15). Note that
$\mathcal{Q}$ $=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}\Big{(}(w_{jt,b}^{+}-w_{jt,b}^{-})Y_{jt}-\sum_{j=1}^{N}(w_{jt,b}^{+}-w_{jt,b}^{-})Y_{jt}/N\Big{)}\Big{|}/\tilde{v}_{j}$
$\approx\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}\Big{(}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}-\sum_{j=1}^{N}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}/N\Big{)}/\tilde{v}_{j}+\tilde{d}_{j}\Big{|},$ (3.6)
where $\tilde{d}_{j}=(\gamma_{j}-\frac{1}{N}\sum_{i=1}^{N}\gamma_{i})/\tilde{v}_{j}.$ The approximation is valid due to the smoothness of $f_{j}(\cdot)$ and the remainder is addressed in Subsection B.3. Let $\tilde{\underline{d}}=(\tilde{d}_{1},\tilde{d}_{2},\ldots,\tilde{d}_{N})^{\top}.$ The following theorem states that the infeasible statistic $\mathcal{Q}$ can be approximated well by the maximum of Gaussian random variables.
Theorem 3.2 (Gaussian Approximation 2).
Suppose that Assumptions 3.1-3.6 hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\Big{(}\mathcal{Q}\leq u\Big{)}-\mathbb{P}(|Z+\tilde{\underline{d}}|_{\infty}\leq u)\Big{|}\lesssim\mathcal{R}_{NT}.$ (3.7)
Theorem 3.2 establishes the foundation for testing the null hypothesis $H_{0}^{\left(2\right)}.$ Its proof cannot directly follow from Theorem 3.1 due to the introduction of the additional averaging term of $Y_{jt}$ which might not be of smaller order. However, the steps can be simplified by imposing weak cross-sectional dependence, as the additional averaging term would then be of higher order under this assumption.
Remark 7.
(Allowance of $N$ and $T$ and bandwidth selection) The relative rate of $N$ and $T$ is determined by the moment assumptions on the error $\eta_{11}$ (Assumption 3.1) and the temporal dependence (Assumption 3.2). When we assume exponential moment conditions (i.e., Assumption 3.1 $(ii)$), $N$ can be of exponential order of $T$. Specifically, if we assume that $b\propto T^{-c}$ with $c>1/5$, then $\mathcal{R}_{NT}\rightarrow 0$ when $\mathrm{log}N\propto T^{r}$, where $30<r<(1-c)/7\wedge(5c-1)\wedge\beta\wedge c/$ 3. When we assume finite moment conditions (i.e., Assumption 3.1 $(i)$), we require $N$ to be of polynomial order of $T$. If we assume that $b\propto T^{-c}$ with $c\in(1/5,1-q/2),$ then $\mathcal{R}_{NT}\rightarrow 0$ when $N\propto T^{\tilde{r}}$ for $0<\tilde{r}<\beta q$. To address the bandwidth selection, we can calculate $\hat{b}_{j}$ for each individual following the procedure of Calonico et al., (2014). Sometimes, it is also recommended to aggregate the bandwidth to avoid the computational burden of estimating the bandwidths for each individual/group.            $\Box$
3.3 Asymptotic Validity of the Feasible Testing Procedures
[…]
Theorem 3.3 (Consistency of the variance estimator).
Suppose that Assumptions 3.1-3.6 hold and $N(\underline{b}T)^{-q/4}\rightarrow 0.$ Then we have $\max_{j\in[N]}|\hat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N))$ and $\max_{j\in[N]}|\hat{v}_{j}^{2}-v_{j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N)).$
[…]
Recall that $q_{\alpha}$ is the $1-\alpha$ quantile of the maximum Gaussian random variable $\max_{j\in\left[N\right]}\left|Z_{j}\right|$ in Algorithm 1. The following theorem studies the asymptotic power properties of the two tests.
Theorem 3.4 (Power).
Suppose that Assumptions 3.1-3.6 hold. Let $\bar{\gamma}=\sum_{j=1}^{N}\gamma_{j}/N.$
[…]
$(ii)$ If $\mathcal{R}_{NT}\to 0$ in Theorem 3.2 and $\max_{j\in\left[N\right]}(Tb_{j})^{1/2}|\gamma_{j}-\bar{\gamma}|\gg(\mathrm{log}N)^{1/2},$ then $\lim_{\left(N,T\right)\rightarrow\infty}\mathbb{P}(\mathcal{\hat{Q}}>q_{\alpha})=1.$
[…]
3.4 Theoretical Results for the Unknown Threshold Case
In this section, we present results related to unknown thresholds. The following theorem provides theoretical support for the GA in cases with unknown $c_{0j}$’s. For simplicity of notation, we define $\mathcal{C}_{j}$ as the unified interval $\mathcal{C}_{j}=[c_{\min},c_{\max}]$, where $c_{\min}$ and $c_{\max}$ are constants in $\mathbb{R}$. The theorem still holds even when $\mathcal{C}_{j}$ differs for different $j$. Recall that $[c_{1},\ldots,c_{K}]$ is the grid we search over. Let $c_{1}=c_{\min}$ and $c_{K}=c_{\max}.$ As our statistics aggregate over different threshold grids, the problem becomes more demanding. Let $\Delta_{c,\min}=\min_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$ and $\Delta_{c,\max}=\max_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$.
Recall $w_{jt,b}(c_{i})$ in (2.16) and define ${\gamma}_{j}(c_{i})=\sum_{t=1}^{T}w_{jt,b}(c_{i})\gamma_{j}$. Let
$d_{jK}=(\gamma_{j}(c_{1})/v_{j}(c_{1}),\ldots,\gamma_{j}(c_{K})/v_{j}(c_{K}))^{\top}$
and $\underline{d}^{C}=(d_{1K}^{\top},d_{2K}^{\top},\ldots,d_{NK}^{\top})^{\top}.$ Define $Z^{C}=\left({{Z_{1}^{C\top},...,Z_{N}^{C\top}}}\right)^{\top}$ as an $NK$-vector of mean zero normally distributed random variables with variance-covariance $\Sigma^{C}$ specified below. For each $j\in[N]$ and $1\leq i_{1},i_{2}\leq K$, the covariance between the $i_{1}$- and $i_{2}$-th elements of $Z_{j}^{C}$ is given by
$\mathrm{Cov}(Z_{j,i_{1}}^{C},Z_{j,i_{2}}^{C})=(v_{j}(c_{i_{1}})v_{j}(c_{i_{2}}))^{-1}(Tb_{j})\sum_{t=1}^{T}w_{jt,b}(c_{i_{1}})w_{jt,b}(c_{i_{2}})\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}.$
${\normalsize{{Z_{j_{1}}^{C}}}}$ and $Z_{j_{2}}^{C}$ are independent for $j_{1}\neq j_{2}$, so that the matrix $\Sigma^{C}$ is block diagonal. Note that $\underline{d}^{C}$ denotes a bias term which is a zero vector under the null hypothesis. In addition, when $2\max_{1\leq j\leq N}b_{j}<\Delta_{c,\min},$ one can ensure the matrix to be diagonal. The following theorem establishes the asymptotic property of our test statistics with unknown thresholds.
Theorem 3.5.
(Gaussian Approximation) Suppose that Assumptions 3.1-3.6 hold. Suppose that $\Delta_{c,\max}\rightarrow 0.$ Then
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{(NK)T}.$ (3.8)
where $\mathcal{R}_{(NK)T}$ is $\mathcal{R}_{NT}$ in Theorem 3.1 with $N$ replaced by $NK$.
The proof of the above theorem is a generalization of Theorem 3.1 with a different variance-covariance structure. It should be noted that the conclusion of the above theorem follows the same pattern as Theorem 3.1, except that $N$ is replaced with $NK$. The following theorem provides the theoretical support for the consistency of the threshold estimator. Note that we only focus on the location $j$ where $\gamma_{j}$ is not equal to 00, and we require those breaks to be significant.
Theorem 3.6 (Consistency of the estimation of thresholds).
Suppose that Assumptions 3.1-3.6 hold. Assume $\min_{1\leq j\leq N:\gamma_{j}\neq 0}(Tb_{j})^{1/2}|\gamma_{j}|\gg 1$ and $\Delta_{c,\max}\rightarrow 0.$ Then
$\max_{1\leq j\leq N,\gamma_{j}\neq 0}\gamma_{j}^{2}|\hat{c}_{j}-c_{0j}|=O_{\mathbb{P}}((\mathrm{log}N)/T).$ (3.9)
Under the assumption of a minimum break signal, we can achieve consistency in detecting the break. The rate is expected to depend on both the signal strength $\gamma_{j}$ and the available sample size $T$. Note that the precision of threshold estimation is not affected by the bandwidth $b_{j}$. This rate is in line with the high-dimensional changepoint literature, see for example Theorem 2 in Li et al., (2024).
[…]
4 Monte Carlo Simulations
In this section, we analyze the finite sample performance of our Gaussian approximation results. In particular, we study the size and power of our uniform testing procedure.
4.1 Data Generating Processes
Throughout the simulation study, we consider the following setup:
$Y_{jt}=\cos(X_{jt})+\sin(U_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq 0\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt},$ (4.1)
where the latent variable is generated by $U_{jt}=\tilde{U}_{jt}X_{jt}$ and $\tilde{U}_{jt}\sim i.i.d.\ U[-1,1]$. To study the performance of our uniform testing procedure, we consider three different data generating processes (DGPs) with different specifications of the covariate variable $X_{jt}$, the error term $\epsilon_{jt}$ and the volatility function $\sigma_{j}(\cdot,\cdot)$. For the first one, we consider the independent setting. For the latter two, we include cross-sectional and temporal dependences. We consider either homoskedastic settings (DGP2) in which $\sigma_{j}(\cdot,\cdot)=1$ and heteroskedastic settings (DGP1, DGP3) in which we set
$\sigma_{j}(x,u)=1+\left(\frac{3}{8}-\frac{\left|x\right|}{4}\right)\left(\frac{3}{2}\right)^{2u}.$ (4.2)
  1. DGP1:
    i.i.d. setting with heteroskedasticity. $X_{jt}\sim i.i.d.\ U[-1,1]$, $\epsilon_{jt}\sim i.i.d.\ N(0,1)$.
  2. DGP2:
    Factor structure and homoskedasticity. Both the running variable and error term have a factor structure, $\epsilon_{jt}=\lambda_{\epsilon,j}f_{\epsilon,t}+u_{\epsilon,jt}$ and $X_{jt}=1/4\left(\lambda_{x,j}f_{x,t}+u_{x,jt}\right)$, where $f_{\epsilon,t}$, $u_{\epsilon,jt}$, $f_{x,t}$ and $u_{x,jt}$ are each generated from an MA($\infty$) process with algebraic decay parameter, $\beta=1.5$. The factor loadings, $\lambda_{\epsilon,j}$ and $\lambda_{x,j}$, are i.i.d. draws from a standard normal distribution.
  3. DGP3:
    Factor structure and heteroskedasticity. The setting is the same as in DGP2 except that the error term is heteroskedastic with the conditional standard deviation described in (4.2).
All DGPs, except for the first one, exhibit strong cross-sectional and weak temporal correlation in the covariate variable and the error term. We are interested in testing the existence of the jump effect, namely, $H_{0}^{\left(1\right)}:\gamma_{j}=0$ for all $j\in[N]$. For each DGP, we study both the size and local power performance of our testing procedure. When studying the size performance, we set all $\gamma_{j}=0$. When examining the local power, we set a fraction of the threshold effects to be $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$, and the remaining $\gamma_{j}$’s to be zero. We consider two different choices for the cross-sectional dimension: $N=10,$ 100. When $N=10$, the above fraction is set to $20\%$, whereas in the high-dimensional setting with $N=100,$ the fraction is decreased to $10\%$.
4.2 Simulation Results
Throughout the study, we use a local linear estimator with a uniform kernel and select a mean square error-optimal bandwidth following the procedure described in Calonico et al., (2014). We consider three significance levels, namely, $\alpha=0.1,$ $0.05,$ and $0.01$. The results are based on $1000$ Monte Carlo iterations.
The results of simultaneous jump effects testing for DGP1 through DGP3 are displayed in the Table 1. First, the empirical size closely matches the nominal size across all $N$ values and for large $T$ in all DGPs. For DGP3, we observe slightly oversize when $N=10$ which may be caused by the presence of heteroskedasticity in comparison with results with DGP2. Overall, these simulation results confirm our theoretical findings in Theorem 3.1. In particular, it shows that the test procedure, which is based on critical values neglecting the dependence in the covariance structure, has fairly accurate size control even in the case of strong cross-sectional dependence in the innovations. Second, the local power properties of the test are reasonably well for all DGPs under investigation, and the power increases as either $N$ or $T$ increases. This result confirms that our testing procedure is well-suited to detect sparse alternatives.
In addition, we present comprehensive simulation results in Section D in the supplementary material. First, we achieve comparable simulation performance in terms of size and power results for the test of homogeneity of the threshold effects in Table A2. The results confirm our Gaussian approximation result in Theorem 3.2. For robustness, we also consider three additional settings with even stronger cross-sectional dependence, see DGP4–6 in Section D of the supplementary materials. The results in Table A1 show that size and power are indeed robust to strong cross-sectional dependence. Furthermore, Table A3 demonstrates that detecting the existence of jump effects simultaneously across cross-sectional dimensions in the case of unknown thresholds has size and power properties as good as in the known threshold case, confirming our theoretical results from Theorem 3.5. Finally, we show the good finite sample performance for the estimation accuracy of the threshold location estimator in Table A4, providing Monte Carlo evidence supporting the consistency result of the threshold in Theorem 3.6.
To demonstrate the practical relevance of our testing procedure, we present two real data applications in Section E of the supplementary material. In Subsection E.1, we detect significant threshold effects in the news impact curve (Engle and Ng, , 1993), i.e., in a model where past returns influence present volatility of stocks. It is an illustration of our testing procedure under unknown thresholds. As a second application, we revisit the study of the incumbency effect in US House elections in Subsection E.2, and find significant jump effects as well as heterogeneity across different states. Since the threshold is naturally given in this problem, it functions as an illustration of our testing procedures for the existence of jump effects and the test for homogeneity in the known case.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP1 10 200 0.113 0.059 0.008 0.395 0.293 0.142
10 400 0.133 0.060 0.014 0.428 0.316 0.153
10 800 0.115 0.054 0.005 0.466 0.333 0.172
100 200 0.099 0.051 0.014 0.884 0.797 0.596
100 400 0.115 0.054 0.011 0.926 0.874 0.699
100 800 0.099 0.050 0.008 0.948 0.922 0.786
DGP2 10 200 0.105 0.054 0.013 0.539 0.443 0.269
10 400 0.122 0.061 0.012 0.583 0.489 0.316
10 800 0.115 0.062 0.010 0.647 0.548 0.377
100 200 0.101 0.042 0.007 0.981 0.962 0.892
100 400 0.107 0.043 0.010 0.982 0.976 0.944
100 800 0.096 0.053 0.010 0.996 0.991 0.964
DGP3 10 200 0.123 0.055 0.007 0.371 0.275 0.105
10 400 0.121 0.058 0.017 0.393 0.299 0.152
10 800 0.122 0.066 0.014 0.465 0.380 0.220
100 200 0.115 0.056 0.019 0.848 0.780 0.594
100 400 0.111 0.061 0.011 0.909 0.849 0.689
100 800 0.116 0.064 0.012 0.952 0.902 0.775
Table 1: Empirical size and power results for the uniform test of the existence of a jump effect for DGP1–3.
5 Conclusion
This paper focuses on the estimation and inference of heterogeneous threshold effects in nonparametric high-dimensional data, accounting for both cross-sectional and time dependencies in the covariates and error terms. We propose a test to assess the significance of the threshold effect and another to evaluate whether the threshold effects are homogeneous across individuals or groups. Additionally, we introduce a consistent method for estimating unknown threshold locations. Our tests are based on high-dimensional Gaussian approximation results. Despite the complexity of the underlying dependency structures, we show that the variance-covariance structure of the threshold effect estimators has a simple analytical expression under general dependence conditions. Unlike changepoint analysis in time series, our approach considers a setup involving latent variables, leading to a significantly different development of the Gaussian approximation results. Simulations and applications demonstrate that the proposed testing procedure performs exceptionally well in finite samples.
SUPPLEMENTARY MATERIAL
[…]
1
Appendix A Proofs of the Main Results in the Paper
In this section, we first outline the main idea for the proof of Theorem 3.1 and present the main lemma for Gaussian approximation. Then we prove Theorems 3.13.4 in turn.
A.1 Outline of the Proof of Theorem 3.1
For exposition purpose, we outline the main step in the proof of Theorem 3.1. The intuition of proving Theorem 3.2 is similar. Define the statistic $I_{\epsilon}$ after canceling the smooth component in the main text,
$I_{\epsilon}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})[\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}]+\gamma_{j}\Big{|}/v_{j}.$ (A.1)
By the continuity of $f_{j}\left(\cdot\right)$ in Assumption 3.4 and Lemma A.2 $(i)$ below, we have
$|\mathcal{I}-I_{\epsilon}|\leq\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})f_{j}(X_{jt})\Big{|}/v_{j}=O(T^{1/2}\bar{b}^{5/2}).$ (A.2)
Thus we only need to work with $I_{\epsilon}.$ We prove by an $m$-lag truncation to facilitate us to construct an $m$-dependent sequence. Specifically, for some $m>0,$ consider a truncated version of the error term:
$\epsilon_{t,m}=(\epsilon_{1t,m},\epsilon_{2t,m},\ldots,\epsilon_{Nt,m})^{\top}=\sum_{k=0}^{m-1}A_{k}\eta_{t-k}.$ (A.3)
Let $I_{\epsilon,m}$ be the $m$-dependent approximation of $I_{\epsilon},$ where $I_{\epsilon,m}$ is $I_{\epsilon}$ with $\epsilon_{t}$ therein replaced by $\epsilon_{t,m},$ namely,
$I_{\epsilon,m}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}\big{[}\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}\big{]}+\gamma_{j}\Big{|}/v_{j},$ (A.4)
where $w_{jt,b}:=(w_{jt,b}^{+}-w_{jt,b}^{-})\mathbf{1}_{\mathcal{A}_{T}}$ where $\mathcal{A}_{T}$ is defined in Lemma A.2 $(ii)$ below. For large $m,$ due to the decay of temporal dependence assumption, we expect
$I_{\epsilon}\approx I_{\epsilon,m}.$
Let $Z_{j,m}$’s be normally distributed random variables that are independent across $j$. Denote the $m$-dependent Gaussian counterpart as
$\tilde{I}_{z,m}=\max_{1\leq j\leq N}\Big{|}Z_{j,m}+\frac{(Tb_{j})^{1/2}}{v_{j}}\gamma_{j}\Big{|}.$
Similarly, let $I_{z}$ (resp. $I_{z,m}$) be $I_{\epsilon}$ (resp. $I_{\epsilon,m}$) with $\eta_{t}$ therein replaced by $z_{t}$, where $\left\{z_{t}\right\}_{t\in\mathbb{Z}}$ are i.i.d. Gaussian vectors with zero mean and identity covariance matrix. Since $I_{\epsilon,m}$ can be rewritten as the maximum of a sum of independent vectors conditional on $(X_{jt},U_{jt})$, we can apply the GA theorem from Chernozhukov et al., 2017b to show the GA for the conditional part. For the remaining part, we need to demonstrate GA again, as detailed in Lemma A.1 below, then we have
$I_{\epsilon,m}\overset{\mathbb{P}}{\approx}I_{z,m}.$
We complete the proof by showing that the distributions of $I_{z,m}$ and $|Z+\underline{d}|_{\infty}$ are close, and the continuity of the maximum of a non-centered Gaussian vector. Then the conclusion will follow.
A.1.1 The Lemma on Gaussian Approximation
This subsection presents a Gaussian approximation lemma that utilizes a nonstandard technique compared to the existing literature. The lemma specifically focuses on the $m$-dependent part $I_{\epsilon,m}$ in (A.4). The difference between $I_{\epsilon,m}$ and $I_{\epsilon}$ is effectively controlled in Lemmas A.7 and A.8 below. Recall that $\mathcal{F}_{t}=(e_{t},e_{t-1},\ldots)$ and ~$\tilde{\mathcal{F}}_{t}=(\tilde{e}_{t},\tilde{e}_{t-1},\ldots).$
The non-triviality of the proof of the theorems lies in the fact that the interplay between the two terms in (A.4):
$\frac{(Tb_{j})^{1/2}}{v_{j}}\sum_{t=1}^{T}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}\quad\textrm{and}\quad\frac{(Tb_{j})^{1/2}}{v_{j}}\sum_{t=1}^{T}w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})$
The second term arises from the potential presence of the latent variable $U_{jt}$. If $U_{jt}$ is absent, the second term will disappear, allowing time-dependent Gaussian approximation to be applied to the first term. This second term causes the derivation steps to differ from those in the existing literature. Noting that the second term becomes determined when we condition on $\{X_{jt},U_{jt}\}_{1\leq t\leq T,1\leq j\leq N}$, we will first prove a GA for the first term under this conditioning. Then we shall replace the first term by its Gaussian counterparts $Z_{j,1}=(Tb_{j})^{1/2}v_{j}^{-1}\sum_{t=1}^{T}\tilde{\xi}_{jt,1}$, where $\tilde{\xi}_{jt,1}$ are i.i.d. normal and independent from $\mathcal{F}_{t}$ and ~$\tilde{\mathcal{F}}_{t}$. Then we can show the GA for the summation of this counterpart and the second term.
The following lemma is the key lemma for Gaussian approximation, whose proof can be found in Section B.1.
Lemma A.1.
Let $T/2\geq m\rightarrow\infty.$
$(i)$ Suppose that the conditions in Theorem 3.1 $(i)$ hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(I_{\epsilon,m}\leq u)-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j})^{1/2}\gamma_{j}/v_{j}\big{|}\leq u\Big{)}\Big{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)},$
where the constant in $\lesssim$ is independent of $N,T.$
$(ii)$ Suppose that the conditions in Theorem 3.1 $(ii)$ hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(I_{\epsilon,m}\leq u)-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j})^{1/2}\gamma_{j}/v_{j}\big{|}\leq u\Big{)}\Big{|}$
$\leq$ $(\underline{b}T)^{-1/6}\mathrm{log}(NT)^{7/6}+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)},$
where the constant in $\lesssim$ is independent of $N,T.$
A.2 Proof of Theorem 3.1
To prove Theorem 3.1, we need the following seven lemmas. See Section B.2 for the proofs of Lemmas A.2, A.7, and A.8. It should be noted that, for simplicity in the exposition of $W_{jt,b}$, we define an event $\mathcal{A}_{T}$ in Lemma A.2, which occurs with probability $1-b_{T}$, where $b_{T}\to 0$.
Lemma A.2.
(Properties for weights) Under Assumptions 3.2-3.5, the followings hold:
[…]
(ii) For $l=0,1,2,$ on some set $\mathcal{A}_{T}$ with
$\mathbb{P}(\mathcal{A}_{T})\geq 1-O(T^{-(\alpha p)\wedge(p/2-1)})$ (A.5)
we have
$\max_{1\leq j\leq N}\big{|}S_{jl,b}^{+}/(Tb_{j}^{l+1})-K_{l}^{+}g_{j}(c_{j0})\big{|}\lesssim\sqrt{\mathrm{log}(TN)/(T\underline{b})}+\bar{b},$ (A.6)
where the constants in $O(\cdot)$ and $\lesssim$ are independent of $N,T,\underline{b},$ and $\bar{b}.$ And thus on set $\mathcal{A}_{T}$,
$\max_{1\leq j\leq N}\Big{|}w_{jt,b}^{+}-\frac{K((X_{jt}-c_{j0})/b_{j})}{g_{j}(c_{j0})Tb_{j}}\frac{K_{2}^{+}-K_{1}^{+}(X_{jt}-c_{j0})/b_{j}}{K_{2}^{+}K_{0}^{+}-K_{1}^{+2}}\mathbf{1}_{\{X_{jt}\geq c_{j0}\}}\Big{|}$
$\lesssim$ $\sqrt{\mathrm{log}(TN)/(\underline{b}T)^{3}}+1/T,$ (A.7)
and moreover,
$\sum_{t=1}^{T}|w_{jt,b}^{+}|=O(1)\quad\text{and}\quad\sum_{t=1}^{T}|w_{jt,b}^{-}|=O(1)$ (A.8)
uniformly over $j$.
The following is a high dimensional extension of the concentration inequality in Freedman, (1975).
Lemma A.3.
(Freedman’s inequality) Let $\mathcal{A}$ be an index set with $|\mathcal{A}|<\infty$. For each $a\in\mathcal{A}$, let $\{\xi_{a,i}\}_{i=1}^{n}$ be a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{i}\}_{i=1}^{n}.$ Let $M_{a}=\sum_{i=1}^{n}\xi_{a,i}$ and $V_{a}=\sum_{i=1}^{n}\mathbb{E}[\xi_{a,i}^{2}|\mathcal{F}_{i-1}]$. Then for all $z,u,v>0$
$\mathbb{P}\Big{(}\max_{a\in\mathcal{A}}|M_{a}|\geq z\Big{)}\leq\sum_{i=1}^{n}\mathbb{P}\Big{(}\max_{a\in\mathcal{A}}|\xi_{a,i}|\geq u\Big{)}+2\mathbb{P}\Big{(}\max_{a\in\mathcal{A}}V_{a}\geq v\Big{)}+2|\mathcal{A}|e^{-z^{2}/(2zu+2v)}.$
Lemma A.4 (Burkholder, (1988), Rio, (2009)).
Let $q>1$, $q^{\prime}=\min\{q,2\}$. Let $M_{T}=\sum_{t=1}^{T}\xi_{t},$ where $\xi_{t}\in\mathcal{L}^{q}$ are martingale differences. Then
$\|M_{T}\|_{q}^{q^{\prime}}\leq K_{q}^{q^{\prime}}\sum_{t=1}^{T}\|\xi_{t}\|_{q}^{q^{\prime}},\mbox{ where }K_{q}=\max\big{(}(q-1)^{-1},\sqrt{q-1}\big{)}.$
Lemma A.5 (Nazarov, (2003)).
Let $X=(X_{1},X_{2},\ldots,X_{v})^{\top}$ be a centered Gaussian vector in $\mathbb{R}^{v}.$ Assume $\mathbb{E}(X_{i}^{2})\geq b$ for some $b>0$ and all $1\leq i\leq v.$ Then for any $e>0$ and $d\in\mathbb{R}^{v},$
$\sup_{x\in\mathbb{R}}\mathbb{P}\Big{(}\big{|}|X+d|_{\infty}-x\big{|}\leq e\Big{)}\leq ce\sqrt{\mathrm{log}(v)},$ (A.9)
where $c$ is some constant depending only on $b$.
The following lemma is Lemma 3 in Chen et al., (2022).
Lemma A.6 (Comparison).
Let $X=(X_{1},X_{2},\ldots,X_{v})^{\top}$ and $Y=(Y_{1},Y_{2},\ldots,Y_{v})^{\top}$ be two centered Gaussian vectors in $\mathbb{R}^{v}$ and let $d=(d_{1},d_{2},\ldots,d_{v})^{\top}\in\mathbb{R}^{v}.$ We denote $\Delta=\max_{1\leq i,j\leq v}|\sigma_{i,j}^{X}-\sigma_{i,j}^{Y}|,$ where we define $\sigma_{i,j}^{X}=\mathbb{E}(X_{i}X_{j})$ (resp. $\sigma_{i,j}^{Y}=\mathbb{E}(Y_{i}Y_{j})$). Assume that $Y_{i}$s have the same variance $\sigma^{2}>0.$ Then we have
$\sup_{x\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}|X+d|_{\infty}\leq x\big{)}-\mathbb{P}\big{(}|Y+d|_{\infty}\leq x\big{)}\Big{|}\lesssim\Delta^{1/3}\mathrm{log}(v)^{2/3},$ (A.10)
where the constant involved in $\lesssim$ only depends on $\sigma.$
Lemma A.7 ($m$-dependent approximation for polynomial case).
Assume that the conditions in Theorem 3.1 $(i)$ hold. For some $1\ll m\leq T/2,$ we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{log}^{q}(NT)(T\underline{b})^{1-q/2}m^{-q\beta}z^{-q}+Ne^{-z^{2}m^{2\beta}}+\mathbb{P}(\mathcal{A}_{T}^{c}).$
where $\mathcal{A}_{T}$ is defined in Lemma A.2.
Lemma A.8 ($m$-dependent approximation for sub-exponential case).
Assume that the conditions in Theorem 3.1$(ii)$ hold. For some $1\ll m\leq T/2,$ we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{exp}\Big{\{}-c\frac{z^{2}}{m^{-\beta}(\underline{b}T)^{-1/2}z+m^{-2\beta}}\Big{\}}+\mathbb{P}(\mathcal{A}_{T}^{c}),$
where $c>0$ is some constant and $\mathcal{A}_{T}$ is defined in Lemma A.2.
Proof of Theorem 3.1.
$(i)$ Recall the definition of $I_{\epsilon}$ in (A.1) and $I_{\epsilon,m}$ is the $m$-dependent part of $I_{\epsilon}$. Let $m=T/2.$ For any $l>0,$ we have
$\mathbb{P}\big{(}\mathcal{I}\leq u\big{)}\leq\mathbb{P}\big{(}|\mathcal{I}-I_{\epsilon,m}|\geq l\big{)}+\mathbb{P}\big{(}I_{\epsilon,m}\leq u+l\big{)},$
and similarly
$\mathbb{P}\big{(}I_{z}\leq u+2l\big{)}\geq-\mathbb{P}\big{(}|I_{z}-I_{z,m}|\geq l\big{)}+\mathbb{P}\big{(}I_{z,m}\leq u+l\big{)}.$
Combining the above inequalities leads to
$\sup_{u\in\mathbb{R}}\Big{[}\mathbb{P}\big{(}\mathcal{I}\leq u)-\mathbb{P}(I_{z}\leq u\big{)}\Big{]}$
$\leq$ $\mathbb{P}\big{(}|\mathcal{I}-I_{\epsilon,m}|\geq l\big{)}+\mathbb{P}\big{(}|I_{z}-I_{z,m}|\geq l\big{)}+\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(I_{\epsilon,m}\leq u)-\mathbb{P}(I_{z,m}\leq u)\big{|}$
$+\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}\big{(}I_{z}\leq u+2l\big{)}-\mathbb{P}\big{(}I_{z}\leq u\big{)}\big{|}$
$=$ $:\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}+\mathrm{I}_{4}.$ (A.11)
For part $\mathrm{I}_{1}$, we have $|\mathcal{I}-I_{\epsilon,m}|\leq|\mathcal{I}-I_{\epsilon}|+|I_{\epsilon}-I_{\epsilon,m}|.$ By Assumption 3.4 and Lemma A.2,
$\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}w_{jt,b}^{+}f_{j}(X_{jt})-f_{j}(c_{j0})\Big{|}$
$=$ $\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}w_{jt,b}^{+}\big{[}f_{j}(c_{j0})+\partial_{+}f_{j}(c_{j0})(X_{jt}-c_{j0})+O((X_{jt}-c_{j0})^{2})\big{]}-f_{j}(c_{j0})\Big{|}$
$=$ $O(\bar{b}^{2}),$
where the constant in $O(\cdot)$ here and all the followings are independent of $N,T$. The same bound can be derived with $w_{jt,b}^{+}$ therein replaced by $w_{jt,b}^{-}$, and thus
$|\mathcal{I}-I_{\epsilon}|\leq\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})f_{j}(X_{jt})\Big{|}/v_{j}=O(T^{1/2}\bar{b}^{5/2}).$ (A.12)
We define $l^{\prime}=N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NT)}$. Then by Lemma A.7 with $m=T/2$, we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\geq l^{\prime}\big{)}=O\big{\{}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{\}}.$ (A.13)
Hence for $l=l^{\prime}+cT^{1/2}\bar{b}^{5/2}$, where $c>0$ is some constant large enough, we have
$\mathrm{I}_{1}=O\big{\{}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{\}}.$
Similar argument leads to the same bound for part $\mathrm{I}_{2}$.
For part $\mathrm{I}_{3}$, by Lemma A.1 $(i)$ we have
$\mathrm{I}_{3}\lesssim(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+T^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{4}$, we use Lemma A.5 to obtain
$\mathrm{I}_{4}\leq\sup_{u\in\mathbb{R}}\mathbb{P}\big{(}|I_{z}-u\big{|}\leq 2l\big{)}\lesssim l\mathrm{log}(N)^{1/2}.$
We complete the proof by combining the $\mathrm{I}_{1}$-$\mathrm{I}_{4}$ parts and a similar argument for the other side of the inequality.
$(ii)$ As in part $(i)$, we have the decomposition as in (A.11). By (A.12) and Lemma A.8 with $m=T/2$, we have $\mathrm{I}_{1}=O((NT)^{-q})$ with $l=O(T^{-\beta}\mathrm{log}(NT)^{1/2}+T^{1/2}b^{5/2}).$ The same bound can be derived for part $\mathrm{I}_{2}$. By Lemma A.1 $(ii)$, we have
$\mathrm{I}_{3}\lesssim(bT)^{-1/6}\mathrm{log}^{7/6}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{4}$, we use the same argument as in part $(i)$ to obtain
$\mathrm{I}_{4}\lesssim l\mathrm{log}(N)^{1/2}\lesssim T^{-\beta}\mathrm{log}(NT)+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}.$
The desired result follows by combining the $\mathrm{I}_{1}$-$\mathrm{I}_{4}$ and a similar argument for the other side of the inequality.
A.3 Proof of Theorem 3.2
Denote
$a_{jt,b}=w_{jt,b}\sigma_{j}(X_{jt},U_{jt})$ (A.14)
and let
$\mathcal{Q}_{\epsilon}=\max_{1\leq j\leq N}\Bigg{|}\sum_{t=1}^{T}$ $\Big{(}a_{jt,b}\epsilon_{jt}-\sum_{l=1}^{N}a_{lt,b}\epsilon_{lt}/N+w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})$
$-\sum_{l=1}^{N}w_{lt,b}\varepsilon_{l}(X_{lt},U_{lt})/N\Big{)}+(\gamma_{j}-\bar{\gamma})\Bigg{|}/\tilde{v}_{j}.$ (A.15)
Denote $\mathcal{Q}_{z}$ be $\mathcal{Q}_{\epsilon}$ with $\epsilon_{jt}$ and $U_{jt}$ replaced by Gaussian distributed random variable with the same covariance structure and $\mathcal{Q}_{\epsilon,m}$ be $\mathcal{Q}_{\epsilon}$ with $\epsilon_{jt}$ therein replaced by $\epsilon_{jt,m}.$ To prove Theorem 3.2, we add two lemmas.
Lemma A.9 ($m$-dependent approximation).
Let $20\ll m\leq T/$ 2 .
(i) Under the conditions in Theorem 3.2(i) we have
$\mathbb{P}\big{(}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{log}(NT)^{q}(T\underline{b})^{1-q/2}m^{-q\beta}z^{-q}+Ne^{-z^{2}m^{2\beta}}+\mathbb{P}(\mathcal{A}_{T}^{c});$
(ii) Under the conditions in Theorem 3.2(ii), we have
$\mathbb{P}\big{(}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{exp}\Big{\{}-c\frac{z^{2}}{m^{-\beta}(\underline{b}T)^{-1/2}z+m^{-2\beta}}\Big{\}}+\mathbb{P}(\mathcal{A}_{T}^{c}).$
Lemma A.10.
Let $20<m\leq T/$ 2 .
(i) Assume the conditions in Theorem 3.2 $(i)$ hold. Then we have
$\sup_{u\in\mathbb{R}}|\mathbb{P}(\mathcal{Q}_{\epsilon,m}\leq u)-\mathbb{P}(\mathcal{Q}_{z,m}\leq u)|$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
(ii) Assume the conditions in Theorem 3.2 $(ii)$ hold. Then we have
$\sup_{u\in\mathbb{R}}|\mathbb{P}(\mathcal{Q}_{\epsilon,m}\leq u)-\mathbb{P}(\mathcal{Q}_{z,m}\leq u)|\lesssim(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
Here the constants in $\lesssim$ are independent of $\left(N,T,b\right).$
Proof of Theorem 3.2
$(i)$ Let $m=T/2.$ For any $l>0,$ as in (A.11), we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}(T\underline{b})^{1/2}\mathcal{Q}\leq u)-\mathbb{P}((T\underline{b})^{1/2}\mathcal{Q}_{z}\leq u\big{)}\Big{|}$
$\leq$ $\mathbb{P}\Big{(}(T\underline{b})^{1/2}|\mathcal{Q}-\mathcal{Q}_{\epsilon,m}|\geq l\Big{)}+\mathbb{P}\Big{(}(T\underline{b})^{1/2}|\mathcal{Q}_{z}-\mathcal{Q}_{z,m}|\geq l\Big{)}+\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(\mathcal{Q}_{\epsilon,m}\leq u)-\mathbb{P}(\mathcal{Q}_{z,m}\leq u)\Big{|}$
$+\sup_{u\in\mathbb{R}}\mathbb{P}\big{(}|(T\underline{b})^{1/2}\mathcal{Q}_{z}-u|\leq 2l\big{)}=:\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}+\mathrm{I}_{4}.$ (A.16)
For part $\mathrm{I}_{1}$, by (A.12), we have
$|\mathcal{Q}-\mathcal{Q}_{\epsilon,m}|\leq|\mathcal{Q}-\mathcal{Q}_{\epsilon}|+|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|=O(\bar{b}^{2})+|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|.$
We define $l^{\prime}=N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NT)}$. Then by Lemma A.9 with $m=T/2$, we have
$\mathbb{P}\big{(}(T\underline{b})^{1/2}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq l^{\prime}\big{)}=O\big{(}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{)}.$ (A.17)
Hence for $l=l^{\prime}+cT^{1/2}\bar{b}^{5/2},$ some constant $c>0$ large enough, we have
$\mathrm{I}_{1}=O(\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}).$
Similar argument leads to the same bound for $\mathrm{I}_{2}$ part. For part $\mathrm{I}_{3}$, by Lemma A.10 we have
$\mathrm{I}_{3}\lesssim(bT)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(bT))^{1/3}\mathrm{log}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{4}$, denote $E=(E_{j_{1}j_{2}})_{1\leq j_{1},j_{2}\leq N}$ as the conditional covariance matrix for $(b_{j}T)^{1/2}\sum_{t=1}^{T}(a_{jt,b}\epsilon_{jt}$ $-\sum_{l=1}^{N}a_{lt,b}\epsilon_{lt}/N)$ given $\mathcal{F}_{T}.$ The diagonal entity takes the form
$E_{jj}$ $=b_{j}T\sum_{h=1-T}^{T-1}\sum_{t=1\vee(1-h)}^{T\wedge(T-h)}\mathbb{E}\Big{[}\big{(}a_{jt,b}\epsilon_{jt}-\sum_{l_{1}=1}^{N}a_{l_{1}t,b}\epsilon_{l_{1}t}/N\big{)}\big{(}a_{j(t+h),b}\epsilon_{j(t+h)}-\sum_{l_{2}=1}^{N}a_{l_{2}(t+h),b}\epsilon_{l_{2}(t+h)}/N\big{)}\Big{]}$
$\cdot(1+O(b_{j})).$
Similarly to the proof of Lemma A.1, with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for some constant $c^{\prime}>0$, we have
$\min_{1\leq j\leq N}E_{jj}\geq c^{\prime}.$
Hence by Lemma A.5, we have
$\sup_{u\in\mathbb{R}}\mathbb{P}\Big{(}(T\underline{b})^{1/2}|\mathcal{Q}_{z}-u|<2l\Big{)}\lesssim l\mathrm{log}(N)^{1/2}.$ (A.18)
Hence
$\mathrm{I}_{4}=O\Big{(}\big{(}T^{1/2}\bar{b}^{5/2}+N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NT)}\big{)}\mathrm{log}(N)^{1/2}\Big{)}.$
By Lemma A.6, we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}(\underline{b}T)^{1/2}\mathcal{Q}_{z}\leq u\big{)}-\mathbb{P}\big{(}|Z+\tilde{\underline{d}}|_{\infty}\leq u\big{)}\Big{|}\lesssim(\underline{b}+T^{-\beta})^{1/3}\mathrm{log}(N)^{2/3}.$ (A.19)
[…]
A.4 Proof of Theorems 3.3-3.4
Proof of Theorem 3.3. Let $\bar{Y}_{jt}$ (resp. $\bar{f}_{jt}$ and $\bar{e}_{jt}$) be the local linear estimator of $Y_{jt}$ (resp. $f_{j}(X_{jt})$ and $e_{jt}$) at location $X_{jt}.$ Consider
$Y_{jt}-\bar{Y}_{jt}=f_{j}(X_{jt})-\bar{f}_{jt}+e_{jt}-\bar{e}_{jt}$
Due to the continuity of $f_{j}$, we have $f_{j}(X_{jt})-\bar{f}_{jt}=O(b_{j})$. Random variables $e_{jt}$ are weakly temporal dependent, by Freedman’s inequality in Lemma A.3, we have
$\mathbb{P}\Big{(}\Big{|}\sum_{t=1}^{T}\mathbf{1}_{|X_{jt}-c_{0,j}|\leq b_{j}}\big{(}e_{jt}^{2}-\mathbb{E}(e_{jt}^{2})\big{)}\Big{|}>x\Big{)}\lesssim b_{j}Tx^{-q/2}+e^{-x^{2}/(b_{j}T)}.$
Then by taking $x=c\mathrm{log}^{1/2}(N)$ for some constant $c>0$ large enough, we have
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}(Tb_{j})^{-1/2}\Big{|}\sum_{t=1}^{T}\mathbf{1}_{|X_{jt}-c_{0,j}|\leq b_{j}}\big{(}e_{jt}^{2}-\mathbb{E}(e_{jt}^{2})\big{)}\Big{|}>x\Big{)}$
$\lesssim$ $N(\underline{b}T)^{1-q/4}\mathrm{log}^{-q/4}(N).$
Hence $\max_{1\leq j\leq N}|\widehat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|=O_{\mathbb{P}}(\bar{b}+(\underline{b}T)^{-1/2}\mathrm{log}^{1/2}(N))$. For $v_{j}^{2}$, note that
$|v_{j}^{2}-\hat{v}_{j}^{2}|=$ $(Tb_{j})\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\big{|}\sigma_{e,j}^{2}-\widehat{\sigma}_{e,j}^{2}\big{|}$
$\leq$ $(Tb_{j})\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\max_{1\leq j\leq N}|\widehat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|.$
Since $\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\lesssim(T\underline{b})^{-1}$, we complete the proof. For case $(ii)$, follow a similar argument as in the proof of Lemma A.8, we can show the desired result in view of
$\mathbb{P}\Big{(}\max_{j}(Tb_{j})^{-1/2}\Big{|}\sum_{t=1}^{T}\mathbf{1}_{|X_{jt}-c_{0,j}|\leq b_{j}}\big{(}e_{jt}^{2}-\mathbb{E}(e_{jt}^{2})\big{)}\Big{|}>x\Big{)}\lesssim Ne^{-x^{2}/(x+1)}.$
[…]
$\leq$ $\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\hat{\gamma}_{j}-\gamma_{j}\right|/\hat{v}_{j}+\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\gamma_{j}\right|/\hat{v}_{j}$
$\lesssim$ $\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\hat{\gamma}_{j}-\gamma_{j}\right|/v_{j}+\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\gamma_{j}\right|:=\text{I}_{1}+\text{I}_{2}.$
[…]
A.5 Proof of Theorem 3.5
Proof.
Denote
$\mathcal{I}_{\epsilon}^{C}=\max_{1\leq i\leq K}\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}(c_{i})\big{[}\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}\big{]}/v_{j}(c_{i})\Big{|}.$ (A.20)
Let $\mathcal{I}_{\epsilon,m}^{C}$ be $\mathcal{I}_{\epsilon}^{C}$ with $\epsilon_{jt}$ therein replaced by $\epsilon_{jt,m}.$ Let $m=T/2.$ For any $l>0,$ similar to (A.11), we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}\mathcal{I}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|\leq u\big{)}\Big{|}$
$\leq$ $\mathbb{P}\Big{(}|\mathcal{I}^{C}-\mathcal{I}^{C}_{\epsilon,m}|\geq l\Big{)}+\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(\mathcal{I}_{\epsilon,m}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|\leq u)\Big{|}$
$+\sup_{u\in\mathbb{R}}\mathbb{P}\big{(}|Z^{C}-u|\leq l\big{)}=:\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}.$ (A.21)
For part $\mathrm{I}_{1}$, by (A.12), we have
$|\mathcal{I}^{C}-\mathcal{I}^{C}_{\epsilon,m}|\leq|\mathcal{I}^{C}-\mathcal{I}^{C}_{\epsilon}|+|\mathcal{I}^{C}_{\epsilon}-\mathcal{I}^{C}_{\epsilon,m}|=O(\bar{b}^{2})+|\mathcal{I}^{C}_{\epsilon}-\mathcal{I}^{C}_{\epsilon,m}|.$
We define $l^{\prime}=N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NKT)}$. Then by a similar argument as in Lemma A.9 with $m=T/2$, we have
$\mathbb{P}\big{(}|\mathcal{I}^{C}_{\epsilon}-\mathcal{I}^{C}_{\epsilon,m}|\geq l^{\prime}\big{)}=O\big{(}\mathrm{log}(NKT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{)}.$ (A.22)
Hence for $l=l^{\prime}+cT^{1/2}\bar{b}^{5/2},$ some constant $c>0$ large enough, we have
$\mathrm{I}_{1}=O(\mathrm{log}(NKT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}).$
For part $\mathrm{I}_{2}$, similar to Lemma A.10 and (A.18), we have
$\mathrm{I}_{2}$ $\lesssim(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NKT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NKT)$
$+\big{(}T^{1/2}\bar{b}^{5/2}+(NK)^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NKT)}\big{)}\mathrm{log}(NK)^{1/2}+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{3},$ by Lemma A.5, we have
$\sup_{u\in\mathbb{R}}\mathbb{P}\Big{(}|Z^{C}-u|<2l\Big{)}\lesssim l\mathrm{log}(NK)^{1/2}.$ (A.23)
Desired result follows by combining $\mathrm{I}_{1}$-$\mathrm{I}_{3}$ parts. ∎
A.6 Proof of Theorem 3.6
For any $c_{0j}-b_{j}<c<c_{0j}$, we can decompose the difference into three parts
$\hat{\gamma}_{j}(c)-\hat{\gamma}_{j}(c_{0j})$
$=$ $\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}Y_{jt}$
$=$ $\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}f_{j}(X_{jt})+\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$+\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}e_{jt}=:\mathrm{I}_{1j}+\mathrm{I}_{2j}+\mathrm{I}_{3j}.$
For part $\mathrm{I}_{1j}$, due to the smoothness of the function $f_{j}(\cdot)$,
$\sum_{t=1}^{T}w_{jt,b}(c)f_{j}(X_{jt})\lesssim b_{j}.$
For part $\mathrm{I}_{2j},$ since $\sum_{t=1}^{T}w_{jt,b}^{+}(c)=\sum_{t=1}^{T}w_{jt,b}^{-}(c)=1,$ we have
$\sum_{t=1}^{T}\big{(}w_{jt,b}(c_{0j})-w_{jt,b}(c)\big{)}\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$=$ $\gamma_{j}\sum_{t}\Big{[}w_{jt,b}(c_{0j})-w_{jt,b}(c-c_{0j}+c_{0j})\Big{]}\textbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$\lesssim$ $\gamma_{j}\sum_{t}\Big{[}\frac{1}{Tb_{j}}\frac{|c-c_{0j}|}{b_{j}}\Big{]}\textbf{1}_{\{|X_{jt}-c_{0j}|<b_{j}\}}$
$\asymp$ $\gamma_{j}|c-c_{0j}|/b_{j}.$
For part $\mathrm{I}_{3j},$ consider the observations in region
$\mathcal{A}_{1}:=\Big{\{}t:X_{jt}\in[c_{0j}-b_{j},c)\Big{\}}.$
In this region, both $w_{jt,b}(c)$ and $w_{jt,b}(c_{0j})$ are non-zero. By Lemma A.2, the variance of $e_{jt}$ in this region can be bounded as
$\mathrm{Var}\Big{(}\sum_{t\in\mathcal{A}_{1}}(w_{jt,b}(c)-w_{jt,b}(c_{0j}))e_{jt}\Big{)}\lesssim(Tb_{j})^{-2}(1/g_{j}(c_{0j})-1/g_{j}(c))^{2}Tb_{j}g_{j}(c_{0j})\lesssim(Tb_{j})^{-1}|c-c_{0j}|^{2}b_{j}^{-2}.$
Consider another kind of region
$\mathcal{A}_{2}:=\Big{\{}t:X_{jt}\in[c-b_{j},c_{0j}-b_{j})\Big{\}}.$
In region $\mathcal{A}_{2},$ different from $\mathcal{A}_{1},$ only $w_{jt,b}(c)$ is non-zero. The variance of noise in this region follows
$\mathrm{Var}\Big{(}\sum_{t\in\mathcal{A}_{2}}(w_{jt,b}(c)-w_{jt,b}(c_{0j}))e_{jt}\Big{)}\lesssim(Tb_{j})^{-2}g_{j}(c_{0j})^{-2}|c-c_{0j}|Tg_{j}(c_{0j})\lesssim(Tb_{j})^{-2}|c-c_{0j}|T.$
The rest regions can be similarly bounded. A similar GA can be constructed as in the proof of Theorem 3.1, which can be used to handle the maximum over $j$. Hence we have
$\max_{1\leq j\leq N}|\mathrm{I}_{3j}|=O_{\mathbb{P}}\Big{(}(Tb_{j})^{-1/2}\mathrm{log}^{1/2}(N)|c-c_{0j}|/b_{j}+(Tb_{j})^{-1}\mathrm{log}^{1/2}(N)|c-c_{0j}|^{1/2}T^{1/2}\Big{)}.$
Since $\hat{\gamma}_{j}(\hat{c})$ shall be the largest, for $c=\hat{c},$
$\gamma_{j}|c-c_{0j}|/b_{j}\lesssim(Tb_{j})^{-1/2}\mathrm{log}^{1/2}(N)|c-c_{0j}|/b_{j}+(Tb_{j})^{-1}\mathrm{log}^{1/2}(N)|c-c_{0j}|^{1/2}T^{1/2},$
which can only be achieved if the following bound uniformly over $j$,
$|c-c_{0j}|=O_{\mathbb{P}}(\mathrm{log}(N)/(\gamma_{j}^{2}T)).$
Appendix B Proofs of the Technical Lemmas in Appendix A
In this section, we prove the technical lemmas in Section A.
B.1 Proofs of Lemma A.1 in Section A.1
Proof of Lemma A.1.
(i). Recall ~$\tilde{\mathcal{F}}_{t}$ in (3.1). Define
$w_{jt,b}=(w_{jt,b}^{+}-w_{jt,b}^{-})\mathbf{1}_{\mathcal{A}_{T}}\text{ and }r_{j}(X_{jt},U_{jt})=\sigma_{j}(X_{jt},U_{jt})/v_{j}.$ (B.1)
Note that $I_{\epsilon,m}$ on $\mathcal{A}_{T}$ can be rewritten as
$I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}$ $=\max_{1\leq j\leq N}\frac{(Tb_{j})^{1/2}}{v_{j}}\Big{|}\sum_{t=1}^{T}w_{jt,b}\big{[}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}+\varepsilon_{j}(X_{jt},U_{jt})\big{]}+\gamma_{j}\Big{|}$
$=\max_{1\leq j\leq N}\frac{(Tb_{j})^{1/2}}{v_{j}}\Big{|}\sum_{k=2-m}^{T}\xi_{jk,1}+\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\Big{|},$ (B.2)
where
$\xi_{jk,1}=\Big{(}\sum_{t=1\vee k}^{T\wedge(k+m-1)}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\Big{)}\eta_{k}\quad\text{and}\quad\xi_{jt,2}=w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt}).$
In the following we shall show that for $B_{T}=c\underline{b}^{-1/2}$ with some constant $c>0$ being large enough:
  1. (a)
    with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for some constant $c^{\prime}>0,$ we have
    $\min_{1\leq j\leq N}(Tb_{j})\sum_{k=2-m}^{T}\mathbb{E}\big{(}\xi_{jk,1}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}\geq c^{\prime};$
  2. (b)
    with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for $l=1,2,$ we have
    $\max_{1\leq j\leq N}T^{1+l}b_{j}^{(2+l)/2}\sum_{k=2-m}^{T}\mathbb{E}(|\xi_{jk,1}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\leq B_{T}^{l}.$
Since $|w_{jt,b}|\lesssim(Tb_{j})^{-1}$, we have $\max_{k}\mathbb{E}(\max_{j}|\xi_{jk,1}|^{q}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\lesssim(Tb_{j})^{-q},$ and thus
$\mathbb{E}\left(\max_{1\leq j\leq N}(Tb_{j}^{1/2}|\xi_{jk,1}|/B_{T})^{q}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\right)\leq 2,$
for all $k.$ Let $\tilde{Z}_{j,1}$, $1\leq j\leq N$, be Gaussian distributed random variables with the same conditional covariance structure as
$(Tb_{j})^{1/2}\sum_{k=2-m}^{T}\xi_{jk,1}/v_{j}$
given $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$. For any fixed $j$, conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T},$ $\xi_{jk,1}$’s are independent for different $k$. Thus combining the above with (a) and (b), by Proposition 2.1 in Chernozhukov et al., 2017b , we have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}\tilde{Z}_{j,1}+(Tb_{j})^{1/2}v_{j}^{-1}\big{(}\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\big{)}\big{|}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT),$
where the conditional variance of $\tilde{Z}_{j,1}$ follows,
$\mathbb{E}(\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}))$
$=$ $(Tb_{j})\mathbb{E}\Big{(}\sum_{t=1}^{T}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}/v_{j}\Big{)}^{2}$
$=$ $\mathbb{E}(v_{j,1}^{2})+(Tb_{j})\sum_{1\leq t_{1}\neq t_{2}\leq T}\mathbb{E}\Big{(}w_{jt_{1},b}\sigma_{j}(X_{jt_{1}},U_{jt_{1}})w_{jt_{2},b}\sigma_{j}(X_{jt_{2}},U_{jt_{2}})/v_{j}^{2}\Big{)}\mathbb{E}\big{(}\epsilon_{jt_{1},m}\epsilon_{jt_{2},m}\big{)}+O(m^{-\beta})$
$=$ $\mathbb{E}(v_{j,1}^{2})+O(b_{j}+m^{-\beta}),$
with
$v_{j,1}^{2}=(Tb_{j})\sum_{t=1}^{T}w_{jt,b}^{2}\mathrm{Var}\big{(}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}|X_{jt}\big{)}/v_{j}^{2}.$ (B.3)
For any $j_{1}\neq j_{2},$ we have $\mathbb{E}(\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j_{2},1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}))=O(\bar{b}).$ Consider the random variable
$Z_{j,1}=(Tb_{j})^{1/2}v_{j}^{-1}\sum_{t=1}^{T}\tilde{\xi}_{jt,1},$ (B.4)
where $\tilde{\xi}_{jt,1}$ are i.i.d. from $N(0,v_{j,1}^{2}v_{j}^{2}/(T^{2}b_{j}))$ and independent from ~$\tilde{\mathcal{F}}_{T}$ and $\epsilon_{jt}$.
By a similar argument as in the proof of part (a) below, with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$ the maximum difference of $\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ (resp. $\mathrm{Var}(Z_{j,1}|\mathcal{F}_{T})$) and $\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j2,1}|$ $\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ (resp. $\mathrm{Cov}(Z_{j_{1},1},Z_{j2,1}|\mathcal{F}_{T})$) from their expectation are bounded by $\bar{b}+\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$ Hence by Lemma A.6, we further have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\Big{|}Z_{j,1}+(Tb_{j})^{1/2}v_{j}^{-1}\big{(}\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\big{)}\Big{|}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$
Then by the construction of $Z_{j,1}$ in (B.4), we have
$Z_{j,1}+(Tb_{j})^{1/2}v_{j}^{-1}\Big{(}\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\Big{)}=(Tb_{j})^{1/2}v_{j}^{-1}\Big{(}\sum_{t=1}^{T}(\tilde{\xi}_{jt,1}+\xi_{jt,2})+\gamma_{j}\Big{)}.$
Given $\mathcal{F}_{T}$, $\tilde{\xi}_{jt,1}+\xi_{jt,2}$ are independent for different $t.$ Recall the definition of $v_{j}^{2}$ in (2.14), we have
$(Tb_{j})v_{j}^{-2}\sum_{t=1}^{T}\mathrm{Var}\big{(}\tilde{\xi}_{jt,1}+\xi_{jt,2}|\mathcal{F}_{T}\big{)}=1.$ (B.5)
Since $|\varepsilon_{j}(\cdot,\cdot)|$ is upper bounded, $\max_{t}\mathbb{E}(|\tilde{\xi}_{jt,1}|+|\xi_{jt,2}||\mathcal{F}_{T})=O((b_{j}T)^{-1}).$ With probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for $l=1,2,$
$\max_{1\leq j\leq N}T^{1+l}b_{j}^{(2+l)/2}\sum_{t=1}^{T}\mathbb{E}\Big{(}|(\tilde{\xi}_{jt,1}+\xi_{jt,2})/v_{j}|^{2+l}|\mathcal{F}_{T}\Big{)}$
$\lesssim$ $\max_{1\leq j\leq N}Tb_{j}^{(2-l)/2}\sum_{t=1}^{T}\mathbb{E}\Big{(}((\tilde{\xi}_{jt,1}+\xi_{jt,2})/v_{j})^{2}|\mathcal{F}_{T}\Big{)}\leq B_{T}^{l}.$ (B.6)
For all $k$, we have
$\mathbb{E}\Big{(}\max_{1\leq j\leq N}(b_{j}^{1/2}Tv_{j}^{-1}|\tilde{\xi}_{jt,1}+\xi_{jt,2}|/B_{T})^{q}\big{|}\mathcal{F}_{T}\Big{)}\leq 2.$
By combining the above results with (B.5) and (B.6), Proposition 2.1 in Chernozhukov et al., 2017b , and Lemma A.6, we have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j})^{1/2}v_{j}^{-1}\gamma_{j}\big{|}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$ (B.7)
Hence we complete the proof.
Proof of part (a). Since $\mathbb{E}(\eta_{jk}^{2})=1,$ we have
$\sum_{k=2-m}^{T}\mathbb{E}(\xi_{jk,1}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ $\geq\sum_{k=1}^{T-m}\Big{|}\sum_{t=k}^{k+m-1}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}\Big{|}_{2}^{2}$
$=\sum_{k=1}^{T-m}\Big{|}\sum_{t=k}^{k+m-1}(M_{jt,1}+D_{jt,1})A_{t-k,j,\cdot}\Big{|}_{2}^{2},$ (B.8)
where
$D_{jt,1}=\mathbb{E}\big{(}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})|\mathcal{F}_{t-1}\big{)}\quad\text{and}\quad M_{jt,1}=w_{jt,b}\sigma_{j}(X_{jt},U_{jk})-D_{jt,1}.$ (B.9)
For the simplicity of notation, denote
$Y_{jk}=\Big{|}\sum_{t=k}^{k+m-1}M_{jt,1}A_{t-k,j,\cdot}\Big{|}_{2}^{2}.$ (B.10)
Since $|D_{jt,1}|\lesssim 1/T,$ (B.8) can be further lower bounded by
$\sum_{k=2-m}^{T}\mathbb{E}(\xi_{jk,1}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\geq\sum_{k=1}^{T-m}\Big{(}Y_{jk}^{1/2}-O(1/T)\Big{)}^{2}\geq\sum_{k=1}^{T-m}Y_{jk}-O\Big{(}\Big{(}\sum_{k=1}^{T-m}Y_{jk}/T\Big{)}^{1/2}\Big{)}.$ (B.11)
We shall bound the difference between $Y_{jk}$ and $\mathbb{E}Y_{jk}$. Note that $\mathbb{E}_{0}Y_{jk}$ can be rewritten as
$\sum_{k=1}^{T-m}\mathbb{E}_{0}Y_{jk}$ $=\sum_{t=1}^{T}\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}(M_{jt,1}L_{k,t,1}+\mathbb{E}_{0}L_{k,t,2})$
$=\sum_{t=1}^{T}\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\big{[}M_{jt,1}L_{k,t,1}+L_{k,t,2}-\mathbb{E}(L_{k,t,2}|\mathcal{F}_{t-1})\big{]}$
$+\sum_{t=1}^{T}\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\big{[}\mathbb{E}(L_{k,t,2}|\mathcal{F}_{t-1})-\mathbb{E}(L_{k,t,2})\big{]}=:\mathrm{I}_{j1}+\mathrm{I}_{j2},$
where
$L_{k,t,1}=2\sum_{s=k}^{t-1}M_{js,1}A_{t-k,j,\cdot}^{\top}A_{s-k,j,\cdot}\quad\text{and}\quad L_{k,t,2}=M_{jt,1}^{2}|A_{t-k,j,\cdot}|_{2}^{2}.$
Let
$\varsigma_{jt}:=\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\big{[}M_{jt,1}L_{k,t,1}+L_{k,t,2}-\mathbb{E}(L_{k,t,2}|\mathcal{F}_{t-1})\big{]}.$
Then $\mathrm{I}_{j1}=\sum_{t=1}^{T}\varsigma_{jt}$, the summation of martingale differences. Since $|M_{jt,1}|=O((Tb_{j})^{-1})$, we have $|\varsigma_{jt}|=O((Tb_{j})^{-2})$ and $\mathbb{E}(\varsigma_{jt}^{2}|\mathcal{F}_{t-1})=O(b_{j}(Tb_{j})^{-4})$. Thus by Freedman’s inequality,
$\mathbb{P}\big{(}b_{j}^{-1/2}(Tb_{j})^{2}|\mathrm{I}_{j1}|\geq z\big{)}\lesssim\mathrm{exp}\Big{\{}-\frac{z^{2}}{b_{j}^{-1/2}z+T}\Big{\}}.$ (B.12)
Hence with probability greater than $1-T^{-p}$ we have
$\max_{1\leq j\leq N}(Tb_{j})|\mathrm{I}_{j1}|\lesssim\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$ (B.13)
For $\mathrm{I}_{j2}$ part, let $\tilde{\sigma}_{j}^{2}(x)=\mathbb{E}(\sigma_{j}^{2}(X_{jt},U_{jk})|X_{jt}=x)$. By Assumption 3.4,
[…]
By Theorem 6.2 in Zhang and Wu, (2017), we have
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}(b_{j}T^{2})|\mathrm{I}_{j2}|\geq z+O(Tb_{j})\Big{)}\lesssim z^{-p}T^{(p/2-\alpha p)\vee 1}\mathrm{log}(N)^{p/2}+e^{-z^{2}/T}.$
Hence with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$ we have
$\max_{1\leq j\leq N}(Tb_{j})|\mathrm{I}_{j2}|\lesssim\sqrt{\mathrm{log}(NT)/T}+\bar{b}.$ (B.15)
For the expectation part, since $M_{jt,1}$ are martingale differences, for $m\leq T/2,$
$\frac{Tb_{j}}{v_{j}^{2}}\sum_{k=1}^{T-m}\mathbb{E}(Y_{jk})=\frac{Tb_{j}}{v_{j}^{2}}\sum_{k=1}^{T-m}\sum_{t=k}^{k+m-1}\mathbb{E}(M_{jt,1}^{2})|A_{t-k,j,\cdot}|_{2}^{2}\asymp c,$
where $c>0$ is some constant independent of $j$. We complete the proof by combining above with (B.13) and (B.15).
Proof of part (b). Recall the definition of $Y_{jk}$ in (B.10). Since $|Y_{jk}|\lesssim(Tb_{j})^{-2}$, we have
$\sum_{k=2-m}^{T}\mathbb{E}(|\xi_{jk,1}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ $\lesssim\sum_{k=2-m}^{T}\Big{|}\sum_{t=1\vee k}^{T\wedge(k+m-1)}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}\Big{|}_{2}^{2+l}$
$\lesssim\sum_{k=2-m}^{T}Y_{jk}^{1+l/2}+\sum_{k=2-m}^{T}\Big{(}\sum_{t=1\vee k}^{T\wedge(k+m-1)}|D_{jt,2}||A_{t-k,j,\cdot}|_{2}\Big{)}^{2+l}$
$\lesssim(Tb_{j})^{-l}\sum_{k=2-m}^{T}Y_{jk}+T^{-1-l}.$ (B.16)
By a similar argument in $(a)$, with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, we have
$\Big{|}\sum_{k=2-m}^{T}Y_{jk}\Big{|}\lesssim(Tb_{j})^{-1}.$ (B.17)
Then part $(b)$ follows from (B.16) and (B.17).
(ii) Since
$\Big{|}\sum_{t=1\vee k}^{T\wedge(k+m-1)}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\Big{|}_{2}=O(1/(b_{j}T)),$
we have
$\mathbb{E}\Big{(}\mathrm{exp}\big{\{}T\sqrt{b_{j}}|\tilde{\xi}_{jt,1}+\xi_{jt,2}|/B_{T}\big{\}}\big{|}\mathcal{F}_{T}\Big{)}\leq 2,$
[…]
B.2 Proofs of Lemmas in Section A.2
Proof of Lemma A.2. $(i)$ The results can be derived by elementary calculus.
$(ii)$ Recall that $\mathbb{E}_{0}(X)=X-\mathbb{E}(X)$. without loss of generality and for notation simplicity, we assume $c_{j0}=0.$ For random variable $X,$ denote $X^{+}=X\mathbf{1}_{\{X\geq 0\}}.$ We shall show that $S_{jl,b}^{+}$ is close to its expectation. To this end, denote
$M_{jt}=X_{jt}^{+l}K(X_{jt}^{+}/b_{j})-\mathbb{E}(X_{jt}^{+l}K(X_{jt}^{+}/b_{j})|\mathcal{F}_{t-1})$
and $D_{jt}=\mathbb{E}(X_{jt}^{+l}K(X_{jt}^{+}/b_{j})|\mathcal{F}_{t-1})-\mathbb{E}(X_{jt}^{+l}K(X_{jt}^{+}/b_{j})).$
Then the difference between $S_{jl,b}^{+}$ and its expectation can be decomposed into two parts
$\mathbb{E}_{0}S_{jl,b}^{+}=\sum_{t=1}^{T}M_{jt}+\sum_{t=1}^{T}D_{jt}=\mathrm{I}_{j1}+\mathrm{I}_{j2}.$
Note that for any $j,$ $M_{jt}$, $t\geq 1,$ are martingale differences with respect to $\mathcal{F}_{t}$. Moreover, we have $M_{jt}\leq|K|_{\infty}b_{j}^{l}$ and
[…]
uniformly over $j$. Thus by Freedman’s inequality (Freedman, (1975)),
$\mathbb{P}\big{(}|\mathrm{I}_{j1}|\geq z\big{)}\lesssim\mathrm{exp}\big{\{}-z^{2}/(b_{j}^{l}z+Tb_{j}^{2l+1})\big{\}},$
where the constants in $\lesssim$ or $O(\cdot)$ here and all the followings of this proof are independent of $N,T,j,b_{j}.$ That is with probability greater than $1-(TN)^{-p}$,
$\max_{1\leq j\leq N}|\mathrm{I}_{j1}/(Tb_{j}^{l+1})|\lesssim\sqrt{\mathrm{log}(TN)/(T\underline{b})}.$ (B.18)
For part $\mathrm{I}_{j2}$, by Taylor’s expansion, we have
[…]
uniformly over $j$. By Theorem 6.2 in Zhang and Wu, (2017), for $z\gtrsim\sqrt{T\mathrm{log}(N)},$
$\mathbb{P}\bigg{(}\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}\mathbb{E}_{0}g_{j,t}(0|\mathcal{F}_{t-1})\Big{|}\geq z\bigg{)}\lesssim T^{(p/2-\alpha p)\vee 1}\mathrm{log}(N)^{p/2}/z^{p}+e^{-z^{2}/T}.$
Hence with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$ we have
$\max_{1\leq j\leq N}|\mathrm{I}_{j2}/(Tb_{j}^{l+1})|\lesssim\bar{b}^{l+1}\sqrt{T\mathrm{log}(TN)}+\bar{b}.$ (B.19)
Combining (B.18) and (B.19), with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, we have
$\max_{1\leq j\leq N}|\mathbb{E}_{0}S_{jl,b}^{+}|/(Tb_{j}^{l+1})\lesssim\sqrt{\mathrm{log}(TN)/(T\underline{b})}+\bar{b}.$
Lastly, we need to approximate the expectation part. Note that
[…]
uniformly over $j$. Hence (A.6) holds. Inequality (A.7) directly comes from (A.6).
[…]
Proof of Lemma A.7. We shall only work on the set $\mathcal{A}_{T}$. Let $w_{jt,b}$ be as defined in (B.1). Then $I_{\epsilon}$ can be rewritten as
$I_{\epsilon}\mathbf{1}_{\mathcal{A}_{T}}=\max_{1\leq j\leq N}\frac{(Tb_{j})^{1/2}}{v_{j}}\Big{|}\sum_{t=1}^{T}w_{jt,b}\big{[}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}+\varepsilon_{j}(X_{jt},U_{jt})\big{]}+\gamma_{j}\Big{|}.$ (B.20)
Recall the definition of $\epsilon_{jt,m}$ in (A.3) and $r_{j}(X_{jt},U_{jt})$ in (B.1), then
$|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\leq$ $\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})(\epsilon_{jt}-\epsilon_{jt,m})\Big{|}$
$=$ $\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})\sum_{k\leq t-m}A_{t-k,j,\cdot}^{\top}\eta_{k}\Big{|}$
$=$ $\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{k\leq T-m}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\eta_{k}\Big{|},$ (B.21)
where $A_{t,j,\cdot}$ is the $j$th row of matrix $A_{t}$. Let
$\xi_{jk}=(Tb_{j})^{1/2}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\eta_{k}.$
Then conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$, $\xi_{jk}$ are independent for different $k,$ and
$|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\leq\max_{1\leq j\leq N}\Big{|}\sum_{k\leq T-m}\xi_{jk}\Big{|}.$
Let
$\sigma^{2}$ $=\max_{1\leq j\leq N}\sum_{k\leq T-m}\mathbb{E}(\xi_{jk}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$
$\lesssim\max_{1\leq j\leq N}Tb_{j}\sum_{k\leq T-m}\Big{(}\sum_{t=1\vee(k+m)}^{T}|w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}|_{2}\Big{)}^{2}.$ (B.22)
Since kernel function $K(\cdot)$ is bounded, by Lemma A.2, $\max_{t,j}(Tb_{j})|w_{jt,b}|=O(1)$ and $\max_{j}\sum_{t=1}^{T}|w_{jt,b}|=O(1),$ where the constants in $O(\cdot)$ and $\lesssim$ here and throughout this proof are independent of $N,T,b,j$. Hence we have
$\max_{j,k}(Tb_{j})\sum_{t=1\vee(k+m)}^{T}|w_{jt,b}A_{t-k,j,\cdot}|_{2}/v_{j}\lesssim m^{-\beta},$ (B.23)
and
$\max_{j}\sum_{k\leq T-m}\sum_{t=1\vee(k+m)}^{T}|w_{jt,b}||A_{t-k,j,\cdot}|_{2}/v_{j}=\max_{j}\sum_{t=1}^{T}\sum_{k\leq t-m}|w_{jt,b}||A_{t-k,j,\cdot}|_{2}/v_{j}\lesssim m^{-\beta}.$ (B.24)
It follows that
$\sigma^{2}\lesssim m^{-2\beta}.$ (B.25)
By Lemma A.3, for any $z>0,$
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\geq z\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$
$\lesssim$ $\sum_{k\leq T-m}\sum_{1\leq j\leq N}\mathbb{P}\Big{(}|\xi_{jk}|\geq u|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}+2N\mathrm{exp}\Big{(}-\frac{z^{2}}{zu+m^{-2\beta}}\Big{)},$ (B.26)
By Markov’s inequality, we have
$\sum_{k\leq T-m}\mathbb{P}\Big{(}|\xi_{jk}|\geq u\Big{)}$ $\leq\sum_{k\leq T-m}u^{-q}\mathbb{E}\big{(}\mathbb{E}(|\xi_{jk}|^{q}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\big{)}$
$\lesssim u^{-q}\sum_{k\leq T-m}(Tb_{j})^{q/2}\mathbb{E}\Big{|}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\Big{|}_{2}^{q}$
$\lesssim(T\underline{b})^{1-q/2}m^{-q\beta}u^{-q}.$ (B.27)
[…]
Proof of Lemma A.8. By (B.21), $|I_{\epsilon}-I_{\epsilon,m}|$ can be bounded by
$|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\leq\max_{1\leq j\leq N}\Big{|}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\Big{|},$
where $c_{jk}=(Tb_{j})^{1/2}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}.$ Then by (B.23) and (B.24), we have
$\max_{j,k}|c_{jk}|_{2}\lesssim m^{-\beta}(\underline{b}T)^{-1/2}\quad\text{and}\quad\max_{j}\sum_{k}|c_{jk}|_{2}^{2}\lesssim m^{-2\beta}.$ (B.28)
Conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$, $c_{jk}^{\top}\eta_{k}$ are independent for different $k$, therefore for $\lambda|c_{jk}|_{2}\leq\lambda_{0}$, where $\lambda_{0}$ is defined in Assumption 3.1. By Markov’s inequality
$\mathbb{P}\Big{(}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\geq z\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}\leq e^{-\lambda z}\prod_{k\leq T-m}\mathbb{E}\big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}.$ (B.29)
Since $\mathbb{E}(\eta_{k}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})=0$, we have
$\mathbb{E}\big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$ $=1+\mathbb{E}\Big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}-1-\lambda c_{jk}^{\top}\eta_{k}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}$
$\leq 1+\mathbb{E}\bigg{(}\frac{e^{|\lambda c_{jk}^{\top}\eta_{k}|}-1-|\lambda c_{jk}^{\top}\eta_{k}|}{\lambda^{2}|c_{jk}|_{2}^{2}}\Big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\bigg{)}\lambda^{2}|c_{jk}|_{2}^{2},$
in view of $e^{x}-x\leq e^{|x|}-|x|$ for any $x.$ Note that for any fixed $x>0,$ $(e^{tx}-tx-1)/t^{2}$ is increasing on $t\in(0,\infty)$. Therefore
$\mathbb{E}\big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$ $\leq 1+\mathbb{E}\bigg{(}\frac{e^{\lambda_{0}|c_{jk}^{\top}\eta_{k}|/|c_{jk}|_{2}}-1-\lambda_{0}|c_{jk}^{\top}\eta_{k}|/|c_{jk}|_{2}}{\lambda_{0}^{2}}\Big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\bigg{)}\lambda^{2}|c_{jk}|_{2}^{2}$
$\leq 1+a_{0}^{\prime}\lambda^{2}|c_{jk}|_{2}^{2}$
$\leq e^{a_{0}^{\prime}\lambda^{2}|c_{jk}|_{2}^{2}},$
where the last inequality is due to $1+x\leq e^{x}$ and $a_{0}^{\prime}=a_{0}/\lambda_{0}^{2}$ with $a_{0}$ defined in Assumption 3.1. Combining the above results with (B.29), we have
$\mathbb{P}\Big{(}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\geq z\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}\leq e^{-\lambda z}\mathbb{E}\Big{(}e^{a_{0}^{\prime}\lambda^{2}\sum_{k\leq T-m}|c_{jk}|_{2}^{2}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}\leq e^{-\lambda z}e^{c\lambda^{2}m^{-2\beta}},$
where the last inequality is due to (B.28) and $c>0$ is some constant. The same argument can be applied for $\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}<-z,$ then we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\geq z\big{)}\leq 2Ne^{-\lambda z+c\lambda^{2}m^{-2\beta}}\lesssim N\mathrm{exp}\Big{\{}-c^{\prime}\frac{z^{2}}{m^{-\beta}(\underline{b}T)^{-1/2}z+m^{-2\beta}}\Big{\}},$
[…]
B.3 Proofs of the Technical Lemmas in Section A.3
Proof of Lemma A.9. $(i)$ Recall the definition of $a_{jt,b}$ in (A.14). We have
$\sum_{t=1}^{T}a_{jt,b}(\epsilon_{jt}-\epsilon_{jt,m})=\sum_{k=-\infty}^{T-m}\sum_{t=1\vee(k+m)}^{T}a_{jt,b}A_{t-k,j,\cdot}^{\top}\eta_{k}.$
Denote
$\xi_{jk}=(Tb_{j})^{1/2}\sum_{t=1\vee(k+m)}^{T}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}^{\top}\eta_{k}/\tilde{v}_{j}.$
Then $\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}$ can be rewritten into
$|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\leq\max_{1\leq j\leq N}\Big{|}\sum_{k\leq T-m}\xi_{jk}\Big{|}.$
Given $\mathcal{F}_{T},\tilde{\mathcal{F}}_{T},$ $\xi_{jk}$ are independent for different $k.$ By the same argument as used to obtain (B.26), we have
$\mathbb{P}\big{(}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq z|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$
$\leq$ $\sum_{k\leq T-m}\mathbb{P}\Big{(}\max_{1\leq j\leq N}|\xi_{jk}|\geq u\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}+2Ne^{-z^{2}/(2zu+2v)},$ (B.30)
where $v$ is the upper bound for
$\sigma^{2}=$ $\max_{1\leq j\leq N}\sum_{k\leq T-m}\mathbb{E}\big{(}\xi_{jk}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$
$=$ $\max_{1\leq j\leq N}\sum_{k\leq T-m}Tb_{j}\Big{|}\sum_{t=1\vee(k+m)}^{T}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}\Big{|}_{2}^{2}/\tilde{v}_{j}^{2}$
$\leq$ $\max_{1\leq j\leq N}Tb_{j}\sum_{k\leq T-m}\Big{(}\sum_{t=1\vee(k+m)}^{T}\Big{(}|a_{jt,b}||A_{t-k,j,\cdot}|_{2}+\sum_{l=1}^{N}|a_{lt,b}||A_{t-k,l,\cdot}|_{2}/N\Big{)}\Big{)}^{2}/\tilde{v}_{j}^{2}.$
As in the proof of (B.25), we have
$\sigma^{2}\lesssim(\bar{b}/\underline{b})m^{-2\beta}\lesssim m^{-2\beta}.$ (B.31)
Since $\sum_{k\geq m}|A_{k,j,\cdot}|_{2}\lesssim m^{-\beta},$ by the same argument as used to derive (B.27), we have
$\sum_{k\leq T-m}\sum_{1\leq j\leq N}\mathbb{P}\Big{(}|\xi_{jk}|\geq u\Big{)}\lesssim N(T\underline{b})^{-q/2+1}m^{-q\beta}u^{-q}.$ (B.32)
Take $u=z/\mathrm{log}(NT)$. Substituting (B.31) and (B.32) into (B.30), we complete the proof of part $(i)$.
$(ii)$ Note that $|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|$ can be bounded by
$|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\leq\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\Big{|}/\tilde{v}_{j},$
where
$c_{jk}=\sum_{t=1\vee(k+m)}^{T}\Big{(}a_{jt,b}A_{t-k,j,i}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,i}/N\Big{)}.$
[…]
Proof of Lemma A.10 $(i)$ Recall the definition of $a_{jt,b}$ in (A.14), then $\mathcal{Q}_{\epsilon,m}$ can be rewritten as
$\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}$ $=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}\Big{(}a_{jt,b}\epsilon_{jt,m}-\sum_{l=1}^{N}a_{lt,b}\epsilon_{lt,m}/N+w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})$
$\quad-\sum_{l=1}^{N}w_{lt,b}\varepsilon_{l}(X_{lt},U_{lt})/N\Big{)}+(\gamma_{j}-\bar{\gamma})\Big{|}/\tilde{v}_{j}$
$=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{k=2-m}^{T}c_{jk}^{\top}\eta_{k}+\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\overline{{\gamma}})\Big{|}/\tilde{v}_{j},$ (B.33)
where
$c_{jk}=\sum_{t=1\vee k}^{T\wedge(k+m-1)}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}$
and
$\xi_{jt}=w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})-\sum_{l=1}^{N}w_{lt,b}\varepsilon_{l}(X_{lt},U_{lt})/N.$
Decompose $a_{jt,b}$ into two parts,
$M_{jt}=a_{jt,b}-\mathbb{E}(a_{jt,b}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})\quad\text{and}\quad D_{jt}=\mathbb{E}(a_{jt,b}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1}).$ (B.34)
Let $\tilde{\sigma}_{j}(X_{jt})=\mathbb{E}(\sigma_{j}(X_{jt},U_{jt})|X_{jt}).$ Since density function $g_{j}$ is bounded,
$|D_{jt}|$ $\lesssim\big{|}\mathbb{E}(\tilde{\sigma}_{j}(X_{jt})K(X_{jt}/b_{j})/(b_{j}T)|\mathcal{F}_{t-1})\big{|}$
$\lesssim\Big{|}\int T^{-1}K(x)\tilde{\sigma}_{j}(b_{j}x)g_{j}(b_{j}x|\mathcal{F}_{t-1})dx\Big{|}\lesssim T^{-1},$
where the constants in $\lesssim$ and $O(\cdot)$ here and all the followings of this proof are independent of $j,t,N,T.$ In the following we shall firstly show that with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$,
$\min_{1\leq j\leq N}(Tb_{j})\sum_{k=2-m}^{T}\mathbb{E}\big{(}(c_{jk}^{\top}\eta_{k})^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}\geq c,$ (B.35)
where $c>0$ is some constant. Note that
$\sum_{k=2-m}^{T}\mathbb{E}\big{(}(c_{jk}^{\top}\eta_{k})^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}\geq\sum_{k=1}^{T-m}\Big{|}\sum_{t=k}^{k+m-1}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}\Big{|}_{2}^{2}.$ (B.36)
Denote $M_{jt}$ and $D_{jt}$ as in (B.34), and let
$Y_{jk}=\Big{|}\sum_{t=k}^{k+m-1}\Big{(}M_{jt}A_{t-k,j,\cdot}-\sum_{l=1}^{N}M_{lt}A_{t-k,l,\cdot}/N\Big{)}\Big{|}_{2}^{2}.$
Since $|D_{jt}|\lesssim 1/T,$ similarly to (B.11), (B.36) can be further lower bounded by
$\sum_{k=2-m}^{T}\mathbb{E}((c_{jk}^{\top}\eta_{k})^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\geq\sum_{k=1}^{T-m}\Big{(}Y_{jk}^{1/2}-O(1/T)\Big{)}^{2}\geq\sum_{k=1}^{T-m}Y_{jk}-O\Big{(}\Big{(}\sum_{k=1}^{T-m}Y_{jk}/T\Big{)}^{1/2}\Big{)}.$
By Assumption 3.6, for $j\neq l$, we have $\mathbb{E}(M_{jt}M_{lt})=O(T^{-2})$ and thus
$\sum_{k=1}^{T-m}\mathbb{E}(Y_{jk})$
$=$ $\sum_{k=1}^{T-m}\sum_{t=k}^{k+m-1}\mathbb{E}\Big{|}M_{jt}A_{t-k,j,\cdot}-\sum_{l=1}^{N}M_{lt}A_{t-k,l,\cdot}/N\Big{|}_{2}^{2}$
$=$ $\sum_{k=1}^{T-m}\sum_{t=k}^{k+m-1}\Big{(}\mathbb{E}(M_{jt}^{2})(1-1/N)^{2}|A_{t-k,j,\cdot}|_{2}^{2}+\sum_{\{l:l\neq j\}}\mathbb{E}(M_{lt}^{2})|A_{t-k,l,\cdot}|_{2}^{2}/N^{2}\Big{)}+O(1/T)$
$\gtrsim$ $(\underline{b}T)^{-1}.$ (B.37)
We shall bound the difference between $Y_{jk}$ and $\mathbb{E}Y_{jk}$. Let $L_{jtk}=M_{jt}A_{t-k,j,\cdot}-\sum_{l=1}^{N}M_{lt}A_{t-k,l,\cdot}/N$ and denote
$\tilde{\xi}_{jt}=\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\Big{(}|L_{jtk}|_{2}^{2}+2L_{jtk}^{\top}\sum_{k\leq s<t}L_{jsk}\Big{)}.$
Then the difference can be decomposed into two parts,
$\sum_{k=1}^{T-m}\mathbb{E}_{0}Y_{jk}=\sum_{t=1}^{T}(\tilde{\xi}_{jt}-\mathbb{E}(\tilde{\xi}_{jt}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1}))+\sum_{t=1}^{T}(\mathbb{E}(\tilde{\xi}_{jt}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})-\mathbb{E}\tilde{\xi}_{jt})=\mathrm{I}_{j1}+\mathrm{I}_{j2},$
Similarly to the argument in (B.12), we have $|\tilde{\xi}_{jt}|=O((b_{j}T)^{-2})$, $\sum_{t=1}^{T}\mathbb{E}(\tilde{\xi}_{jt}^{2}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})=O((b_{j}T)^{-3})$ and thus
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}|\mathrm{I}_{j1}|\geq z/(b_{j}T)^{2}\Big{)}\lesssim N\mathrm{exp}\Big{\{}-\frac{z^{2}}{z+b_{j}T}\Big{\}}.$ (B.38)
Hence with probability greater than $1-T^{-p}$, we have $\max_{1\leq j\leq N}(b_{j}T)|\mathrm{I}_{j1}|=\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$
For $\mathrm{I}_{j2}$ part, similar to (B.14), we have
$\mathbb{E}(M_{jt}^{2}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})=\int b_{j}(b_{j}T)^{-2}K^{2}(x)\tilde{\sigma}_{j}^{2}(c_{j0})g_{j}(c_{j0}|\mathcal{F}_{t-1})\mathrm{d}x(1+O(b_{j})),$
and that
$\mathbb{E}(\tilde{\xi}_{jt}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})=\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\mathbb{E}(|L_{jtk}|_{2}^{2}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1}).$
Hence by Theorem 6.2 in Zhang and Wu, (2017),
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}(b_{j}T^{2})|\mathrm{I}_{j2}|\geq z+O(\bar{b})\Big{)}\lesssim z^{-p}T^{(p/2-\alpha p)\vee 1}\mathrm{log}(N)^{p/2}+e^{-z^{2}/T}.$
Hence combining $\mathrm{I}_{j1}$ and $\mathrm{I}_{j2},$ with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$
$b_{j}T\Big{|}\sum_{k=1}^{T-m}\mathbb{E}_{0}Y_{jk}\Big{|}\lesssim\sqrt{\mathrm{log}(NT)/(\underline{b}T)}+\bar{b}.$ (B.39)
This, in conjunction with (B.37), completes the proof of (B.35).
We shall then show that with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for $B_{T}=c\underline{b}^{-1/2},$ some constant $c>0$ large enough, $l=1,2,$ we have
$\max_{1\leq j\leq N}b_{j}^{(2+l)/2}T^{1+l}\sum_{k=2-m}^{T}\mathbb{E}(|c_{jk}^{\top}\eta_{k}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\leq B_{T}^{l}.$ (B.40)
Since $\eta_{ik}$ are independent for different $i,k$,
$\sum_{k=2-m}^{T}\mathbb{E}(|c_{jk}^{\top}\eta_{k}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\lesssim\sum_{k=2-m}^{T}|c_{jk}|_{2}^{2+l}\lesssim(b_{j}T)^{-l}\sum_{k=2-m}^{T}Y_{jk}+T^{-(1+l)}.$ (B.41)
By a similar argument as used to derive (B.37), we have $b_{j}T\sum_{k=2-m}^{T}\mathbb{E}Y_{jk}=O(1).$ Hence together with (B.39), with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$
$b_{j}T\sum_{k=2-m}^{T}Y_{jk}=O(1).$ (B.42)
Combining (B.41) and (B.42), we complete the proof of (B.40). Let $\tilde{Z}_{j,1}$, $1\leq j\leq N$, be Gaussian distributed random variables with the same conditional covariance structure as $(Tb_{j})^{1/2}\sum_{k=2-m}^{T}c_{jk}^{\top}\eta_{k}$ given $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$. For any $j$ conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T},$ $c_{jk}^{\top}\eta_{k}$ are independent for different $k$. Thus by Proposition 2.1 in Chernozhukov et al., 2017b ,
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\Big{|}\tilde{Z}_{j,1}+(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\bar{\gamma})\Big{)}\Big{|}/\tilde{v}_{j}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT).$
Define
$\tilde{v}_{j,1}^{2}=v_{j}^{2}\big{\{}(1-1/N)^{2}v_{j,1}^{2}+\sum_{l=1}v_{l,1}^{2}/N^{2}\big{\}},$
where $v_{j,1}$ defined in (B.3). Note that
$\mathbb{E}(\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}))=\tilde{v}_{j,1}^{2}+O(b_{j}+m^{-\beta}).$
And for any $j_{1}\neq j_{2}$ we have
$\mathbb{E[}\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j_{2},1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})]=O(\bar{b}).$
Let $Z_{j,1}$ be i.i.d. from $N(0,\tilde{v}_{j,1}^{2})$ and is independent of $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$. With probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$ the maximum difference of $\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ and $\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j2,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ from their expectation are bounded by $\bar{b}+\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$ Hence by Lemma A.6, we further have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\Big{|}Z_{j,1}+(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\bar{\gamma})\Big{)}\Big{|}/\tilde{v}_{j}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$
Random variable $Z_{j,1}=(Tb_{j})^{1/2}\sum_{t=1}^{T}\tilde{\xi}_{jt,1},$ where $\tilde{\xi}_{jt,1}$ are i.i.d. from $N(0,\tilde{v}_{j,1}^{2}/(T^{2}b_{j}))$ and independent from $\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}$. Then
$Z_{j,1}+(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\bar{\gamma})\Big{)}=(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}(\tilde{\xi}_{jt,1}+\xi_{jt})+(\gamma_{j}-\bar{\gamma})\Big{)}.$
Define $Z_{j}^{\diamond}$s as independent Gaussian random variable with mean 00 and variance $\tilde{v}_{j}^{2}$. Given $\mathcal{F}_{T}$, $\tilde{\xi}_{jt,1}+\xi_{jt,2}$ are independent for different $t.$ Similar to the argument in the proof of (B.7), we have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}^{\diamond}+(Tb_{j})^{1/2}(\gamma_{j}-\bar{\gamma})\big{|}/\tilde{v}_{j}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$ (B.43)
Proof of $(ii)$. Since $\max_{j,k}|c_{jk}|_{2}=O(1/(\underline{b}T)),$ for $B_{T}=c\underline{b}^{-1/2}$ some constant $c>0$ large enough, we have $Tb_{j}^{1/2}|c_{jk}|_{2}/B_{T}\leq\lambda_{0}\mathrm{log}_{(a_{0}\vee 2)}(2)$ and thus
$\mathbb{E}\big{(}\mathrm{exp}\{Tb_{j}^{1/2}|c_{jk}^{\top}\eta_{k}|/B_{T}\}\big{)}\leq 2.$
[…]
B.4 Additional Theorems in the Unknown Threshold Case
Let $\hat{e}_{jt}$ be the residual by regressing $Y_{jt}$ on $X_{jt}$ after a local linear regression with a bandwidth $b^{*}\ll\underline{b}$. Let $\phi_{\acute{a}}(x)=\min(\acute{a},|x|)$ with $\acute{a}=c\mathrm{log}^{1/2}(NK).$ Define
$\hat{\sigma}_{e,j}^{2}(c_{i})=\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}\phi_{\acute{a}}(\hat{e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}.$ (B.44)
Theorem B.1 (Consistency of the variance estimator).
Suppose that Assumptions 3.1-3.6 hold and $N(bT)^{-q/4}\rightarrow 0.$ Then we have
$\max_{j\in[N]}\max_{1\leq i\leq K}|\hat{\sigma}_{e,j}^{2}(c_{i})-\sigma_{e,j}^{2}(c_{i})|$ $=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b}+\acute{a}^{-p/2+1}+b^{*2})$ and $\max_{j\in[N]}|\hat{v}_{j}^{2}-v_{j}^{2}|$ $=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b}+\acute{a}^{-p/2+1}+b^{*2})$
Theorem B.2 (Consistency of the tests).
Suppose that Assumptions 3.1-3.6 hold. Let $\bar{\gamma}=\sum_{j=1}^{N}\gamma_{j}/N.$
[…]
Proof.
Recall that $\hat{e}_{jt}=e_{jt}+f_{j}(X_{jt})-\hat{f}_{j,b^{*}}(X_{jt})$ for the observations with $|X_{jt}-c_{i}|\leq b^{*}$ where $b^{*}\ll\underline{b}$. For $c_{i}$ satisfying $\min_{1\leq i\leq K}|c_{i}-c_{0j}|\leq b^{*},$ we have
$\hat{\sigma}_{e,j}^{2}(c_{i})=$ $\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j},|X_{jt}-c_{0j}|\leq b^{*}\}}\phi_{\acute{a}}(\hat{e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}+\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|>b^{*},|X_{jt}-c_{i}|\leq b_{j}\}}\phi_{\acute{a}}(\hat{e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}$
$=$ $:\mathrm{I}_{1ij}+\mathrm{I}_{2ij}.$
By a similar argument as in Lemma A.2, we have for $x\gtrsim\sqrt{T\mathrm{log}(NK)},$
$\mathbb{P}\Big{(}\max_{1\leq i\leq K,1\leq j\leq N}\Big{|}\sum_{t=1}^{T}\big{(}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}-\mathbb{E}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}\big{)}\Big{|}>x\Big{)}$
$\lesssim$ $T^{(p/2-\alpha p)\vee 1}\mathrm{log}(NK)^{p/2}/x^{p}+e^{-x^{2}/T}.$ (B.45)
Same argument can be applied on $\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j},|X_{jt}-c_{0j}|\leq b^{*}\}}$. For part $\mathrm{I}_{1ij},$ since $|\phi_{\acute{a}}|\leq\acute{a}$, by the above bounds we have
$\max_{1\leq i\leq K,1\leq j\leq N}|\mathrm{I}_{1ij}|\leq\max_{1\leq i\leq K,1\leq j\leq N}\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j},|X_{jt}-c_{0j}|\leq b^{*}\}}\acute{a}}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b})$
For $\mathrm{I}_{2ij}$ part, since $\max_{i,j}|f_{j}(X_{jt})-\hat{f}_{j,b^{*}}(X_{jt})|=O(b^{*2}),$ we have $\max_{i,j}|\phi_{\acute{a}}({e}_{jt}^{2})-\phi_{\acute{a}}(\hat{e}_{jt}^{2})|=O(b^{*2})$ and
$\max_{1\leq i\leq K,1\leq j\leq N}|\mathrm{I}_{2ij}|$
$\lesssim$ $b^{*2}+\max_{1\leq i\leq K,1\leq j\leq N}\Big{|}\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|>b^{*},|X_{jt}-c_{i}|\leq b_{j}\}}\phi_{\acute{a}}({e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}-\mathbb{E}(\phi_{\acute{a}}({e}_{jt}^{2})|X_{jt}=c_{i})\Big{|}$
$+\max_{1\leq i\leq K,1\leq j\leq N}\big{|}\mathbb{E}(\phi_{\acute{a}}({e}_{jt}^{2})|X_{jt}=c_{i})-\mathbb{E}({e}_{jt}^{2}|X_{jt}=c_{i})\big{|}=:b^{*2}+\mathrm{II}_{1}+\mathrm{II}_{2}.$
Similar to (B.4), we have $\mathrm{II}_{1}=O_{\mathbb{P}}((Tb_{j})^{-1/2}\mathrm{log}^{1/2}(NK))$. For $\mathrm{II}_{2}$ part, by Markov’s inequality we have $\mathrm{II}_{2}=O(\acute{a}^{-p/2+1}).$
Appendix C Additional Algorithms
We shall list the two additional algorithms, i.e. Algorithm 2 and Algorithm 3.
Algorithm 2 Testing for homogeneity, i.e., $\gamma_{1}=\ldots=\gamma_{N}$.
1:for $j=1,\ldots,N$ do
2:     Estimate $\gamma_{j}$ for each group/individual $j$ using equation (2.11).
3:     Estimate the variance $v_{j}^{2}$ of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$ by $\hat{v}_{j}^{2}=Tb_{j}\sum_{t=1}^{T}\left(w_{jt,b}^{+}-w_{jt,b}^{-}\right)^{2}\hat{\sigma}_{e,j}^{2}$, where $\hat{\sigma}_{e,j}^{2}$ is a consistent estimator for the error variance using the observations close to the threshold $c_{0j}$.
4:end for
5:Calculate the feasible test statistic $\hat{\mathcal{Q}}:=\mbox{max}_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}-\frac{1}{N}\sum_{i=1}^{N}\hat{\gamma}_{i}|/\hat{\tilde{v}}_{j}$, where $\hat{\tilde{v}}_{j}=\left(\hat{v}_{j}^{2}\left(1-\frac{1}{N}\right)^{2}+\frac{1}{N^{2}}\sum_{i\neq j}^{N}\hat{v}_{i}^{2}\right)^{1/2}$.
6:Reject the null hypothesis if $\hat{\mathcal{Q}}>q_{\alpha}$, where $q_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}|Z_{j}|$ with $Z_{j}$’s being independent standard normal variables.
Algorithm 3 Testing for existence of jumps: i.e., $\gamma_{j}=0$ for all $j$ with unknown threshold.
1:for $j=1,\ldots,N$ do
2:     For a given grid of $[c_{1},c_{2},\cdots,c_{K}]$, estimate $\gamma_{j}(c_{i})$, for each threshold location $c_{i}$, ($i=1,\cdots,K$), using equation (2.11).
3:     Estimate the variance $v_{j}^{2}(c_{i})$ of $(Tb_{j})^{1/2}\hat{\gamma}_{j}(c_{i})$ by $\hat{v}_{j}^{2}(c_{i})=Tb_{j}\sum_{t=1}^{T}\left(w_{jt,b}^{+}(c_{i})-w_{jt,b}^{-}(c_{i})\right)^{2}\hat{\sigma}_{e,j}^{2}(c_{i})$ for each $i$, where $\hat{\sigma}_{e,j}^{2}(c_{i})$ is estimated by (B.44).
4:end for
5:Calculate the feasible test statistic ^$\hat{\mathcal{I}}^{C}:=\max_{1\leq i\leq K}\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$.
6:Reject the null hypothesis if ^$\hat{\mathcal{I}}^{C}>q^{C}_{\alpha}$, where $q^{C}_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}\max_{1\leq l\leq K}|Z_{jl}|$ with $Z_{jl}$’s being independent standard normal variables.
7:Return the $\hat{c}_{j}$ which corresponds to the maximum absolute value of $\max_{1\leq i\leq K}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$.
Appendix D Additional Simulation Results
The finite sample size and power properties for the test of homogeneous effects are also examined for DGPs 1-3, as shown in Table A1. The findings are consistent with those of the test for the existence of jump effects, with the empirical size closely matching the nominal size across all the scenarios. However, for DGP3, the test procedure shows a slight tendency to over-reject the null, potentially due to the presence of heteroskedasticity in the error terms. The local power results align with the earlier findings from the significance test for the existence of jump effects.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP1 10 200 0.114 0.070 0.010 0.315 0.226 0.092
10 400 0.114 0.065 0.014 0.374 0.267 0.118
10 800 0.119 0.056 0.010 0.416 0.286 0.113
100 200 0.104 0.058 0.018 0.845 0.741 0.519
100 400 0.118 0.062 0.014 0.897 0.824 0.603
100 800 0.120 0.062 0.013 0.933 0.882 0.709
DGP2 10 200 0.115 0.067 0.011 0.440 0.329 0.160
10 400 0.111 0.059 0.013 0.483 0.382 0.188
10 800 0.107 0.059 0.010 0.563 0.445 0.272
100 200 0.111 0.052 0.012 0.954 0.922 0.816
100 400 0.104 0.051 0.013 0.984 0.960 0.898
100 800 0.115 0.055 0.009 0.995 0.992 0.951
DGP3 10 200 0.115 0.070 0.012 0.315 0.216 0.081
10 400 0.123 0.074 0.012 0.353 0.242 0.094
10 800 0.113 0.059 0.015 0.386 0.269 0.122
100 200 0.125 0.054 0.011 0.801 0.712 0.498
100 400 0.123 0.060 0.009 0.854 0.787 0.572
100 800 0.128 0.058 0.009 0.923 0.873 0.701
Table A1: Empirical size and power results for the uniform test for heterogeneity for DGP1–3.
Now we consider three additional simulation settings with strong cross-sectional and weak temporal dependence.
  1. DGP4:
    Strong cross-sectional dependence. The setting is similar to that in DGP3, but now the factor term in the running variable and error dominate the idiosyncratic part via a nonzero mean and a relatively larger variance,
    $\epsilon_{jt}$ $=\left(\lambda_{\epsilon,j}+2\right)f_{\epsilon,t}+u_{\epsilon,jt}/8,$
    $X_{jt}$ $=\frac{1}{4}\left((\lambda_{x,j}+2)f_{x,t}+u_{x,jt}/8\right).$
  2. DGP5:
    Semi-strong cross-sectional dependence. The setting is similar to that in DGP4, with some more weight put on the idiosyncratic part of the error term and the running variable,
    $\epsilon_{jt}$ $=\left(\lambda_{\epsilon,j}+2\right)f_{\epsilon,t}+u_{\epsilon,jt}/4$
    $X_{jt}$ $=\frac{1}{4}\left((\lambda_{x,j}+2)f_{x,t}+u_{x,jt}/4\right).$
  3. DGP6:
    Semi-strong cross-sectional dependence and time effect. The setting is similar to that in DGP5, with an additional time effect for the error term,
    $\epsilon_{jt}=\left(\lambda_{\epsilon,j}+2\right)f_{\epsilon,t}+\nu_{t}+u_{\epsilon,jt}/4,$
    where $\nu_{t}\sim i.i.d.\ N(0,0.25)$.
We study the size and power properties of the test for the existence of jump effects for DGP4 – DGP6 in Table A2. These DGPs are the ones with the strong cross-sectional dependence we consider in our simulation study. For DGP4, we observe an empirical size slightly below the nominal size for the high-dimensional setting with $N=100$ which is due to the increase in the cross-sectional dependence in the running variable and the error term. The size properties for DGP5 are almost perfect, while the inclusion of an additional time effect in DGP6 increases the empirical size slightly. This time effect adds more noise to the system, so it is not surprising that the empirical power of DGP6 is lower than that of the other DGPs we consider.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP4 10 200 0.104 0.051 0.012 0.311 0.246 0.157
10 400 0.100 0.045 0.008 0.334 0.273 0.193
10 800 0.098 0.048 0.014 0.380 0.314 0.209
100 200 0.084 0.042 0.006 0.755 0.711 0.625
100 400 0.079 0.039 0.009 0.798 0.757 0.671
100 800 0.090 0.047 0.014 0.828 0.792 0.728
DGP5 10 200 0.108 0.058 0.010 0.320 0.240 0.141
10 400 0.114 0.052 0.018 0.368 0.289 0.199
10 800 0.100 0.051 0.011 0.397 0.312 0.199
100 200 0.107 0.053 0.012 0.753 0.702 0.581
100 400 0.106 0.056 0.009 0.816 0.780 0.685
100 800 0.102 0.058 0.009 0.858 0.807 0.713
DGP6 10 200 0.100 0.056 0.017 0.216 0.137 0.053
10 400 0.119 0.059 0.015 0.259 0.174 0.068
10 800 0.108 0.060 0.009 0.270 0.185 0.076
100 200 0.122 0.071 0.013 0.477 0.391 0.229
100 400 0.110 0.056 0.010 0.564 0.468 0.283
100 800 0.117 0.049 0.007 0.647 0.546 0.376
Table A2: Empirical size and power results for the uniform test of the existence of jump effects for DGP4–6.
While the preceding results are concerned with the known threshold case, we now exam the size and power properties of our uniform testing procedure in the case of unknown thresholds. We consider the same model specification as in the known threshold case. Therefore, the true threshold is $c_{0j}=0$ for all $j$. For the practical implementation of our test procedure, we consider grids of threshold locations, $[-0.6,-0.3,0,0.3,0.6]$ for DGP1 and $[-0.3,-0.15,0,0.15,0.3]$ for DGP2 and 3. The results in Table A3 reveal that the test has still good size properties, with the empirical size being smaller but close to the nominal one. As expected, the power of the test decreased compared with the known threshold case.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP1 10 200 0.075 0.036 0.005 0.120 0.067 0.020
10 400 0.092 0.038 0.006 0.154 0.086 0.022
10 800 0.117 0.065 0.012 0.210 0.141 0.052
100 200 0.075 0.038 0.006 0.200 0.127 0.044
100 400 0.061 0.030 0.007 0.393 0.263 0.094
100 800 0.086 0.043 0.008 0.584 0.453 0.206
DGP2 10 200 0.101 0.043 0.010 0.314 0.242 0.120
10 400 0.096 0.040 0.011 0.408 0.314 0.186
10 800 0.093 0.043 0.008 0.481 0.396 0.250
100 200 0.071 0.034 0.006 0.878 0.802 0.618
100 400 0.093 0.047 0.013 0.942 0.913 0.814
100 800 0.095 0.044 0.004 0.979 0.969 0.911
DGP3 10 200 0.081 0.040 0.011 0.205 0.116 0.038
10 400 0.093 0.042 0.008 0.239 0.164 0.065
10 800 0.085 0.040 0.006 0.253 0.167 0.086
100 200 0.062 0.026 0.001 0.523 0.425 0.228
100 400 0.073 0.034 0.011 0.698 0.585 0.360
100 800 0.090 0.042 0.009 0.810 0.732 0.541
Table A3: Empirical size and power results for the uniform test of the existence of jump effects under unknown threshold locations for DGP1–3.
Finally, we check the estimation accuracy of $\hat{c}_{j}$, our estimator of the unknown threshold location, $c_{0j}$. For this purpose, we set all $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}5B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$. We consider a grid of possible threshold locations, i.e. $[-0.3,-0.29,\ldots,0.3]$. The results in Table A4 reveal that the maximum and averaged error can be reduced effectively with increasing $T$.
$N$ $T$ $1/N\sum_{j}|\hat{c}_{j}-c_{0j}|$ $\max_{j}|\hat{c}_{j}-c_{0j}|$
DGP1 10 200 0.0241 0.1399
10 400 0.0140 0.0979
10 800 0.0087 0.0668
10 1600 0.0063 0.0528
50 200 0.0156 0.2260
50 400 0.0069 0.1520
50 800 0.0035 0.1008
50 1600 0.0021 0.0679
DGP2 10 200 0.0085 0.0635
10 400 0.0058 0.0486
10 800 0.0038 0.0332
10 1600 0.0029 0.0272
50 200 0.0048 0.1177
50 400 0.0024 0.0811
50 800 0.0015 0.0564
50 1600 0.0010 0.0381
Table A4: Estimation accuracy of the $\hat{c}_{j}$’s for DGP1–2.
Appendix E Empirical Applications
In this section, we apply our testing procedure to two empirical applications. First, we study possible jumps in the news impact curve for constituents of the S&P 500. As a second application, we study the impact of the incumbency effect in U.S. House elections.
E.1 Application I: Jumps in the News Impact Curve
In this application, we revisit the news impact curve introduced by Engle and Ng, (1993). In particular, we want to identify possible jumps in the relationship between news events in the form of past returns on the volatility of stocks. The application serves as an illustration of our uniform testing procedure when the threshold $c_{0j}$ is unknown. Originally proposed in the context of autoregressive volatility models (e.g., ARCH), Engle and Ng, (1993) and Linton and Mammen, (2005) consider partially nonparametric specifications for the news impact curve which are able to capture possible asymmetries. A limitation of these specifications is the restriction to continuous functions. However, Zakoian, (1994) and Fornari and Mele, (1997) advocate for (parametric) GARCH specifications that allow for discontinuities in the way past returns impact volatility. We therefore want to investigate this issue by testing for the presence of jumps. In our model setup, the dependent variable $Y_{jt}$ is the Garman-Klass volatility of stock $j$ at time $t$, and $X_{jt}$ is the lagged stock return.
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\textbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}.$
We want to test the null hypothesis $H_{0}^{(1)}:\gamma_{i}=0$ for all $i\in[N]$ and all $c_{0j}$, versus the alternative $H_{a}^{(1)}:\gamma_{i}\neq 0$ for some $i$ and some $c_{0j}$. We want to reiterate that our testing procedure is robust to the presence of unobserved risk factors, $U_{jt}$, which are more than likely to occur in our application. Our data includes the daily stock returns and volatility data of the S&P 500 constituents from January 2010 to May 2024, which we obtained from Kaggle  [^1] https://www.kaggle.com/datasets/andrewmvd/sp-500-stocks. Therefore we have a fairly high-dimensional setting with $N=500$ and $T_{j}$ is the number of days of stock $j$ being a member of the stock index. As candidates for the threshold location, we consider $c_{i}\in\{-0.01,-0.005,0,0.005,0.01\}$.
The main findings are summarized in Table A5. We find significant results for $15$ out of the $500$ stocks we consider. For stocks with significant threshold effects, the locations of the estimated thresholds vary, and none of the significant threshold effects occur precisely at zero. This contradicts the rationale of the sign-switching ARCH model of Fornari and Mele, (1997), which models a jump effect based on the sign of the lagged return, and thus expects a significant threshold effect at zero. For $c_{i}=-0.01$, we get significant negative effects at the $1\%$ level for five stocks, including Allstate Corporation and Ecolab. On the other hand, we obtain significant positive effects for the $c_{i}=0.01$ threshold for eight stocks, e.g. AT&T, Booking, and Costco. These results reaffirm the findings of Chen and Ghysels, (2011), who find that both very bad news and very good news have an increasing effect on volatility. Additionally, we get another negatively significant coefficient at the $-0.005$ threshold location and another positively significant coefficient at $c_{i}=0.005$. Overall, we get evidence for the presence of jump effects in the news impact curve.
Symbol $c_{0j}$ $100\cdot\hat{\gamma}_{j}(c_{0j})$ $100\cdot(T_{j}b_{j})^{-1/2}\hat{v}_{j}(c_{0j})$ $(T_{j}b_{j})^{1/2}\hat{\gamma}_{j}(c_{0j})/\hat{v}_{j}(c_{0j})$
ALL –0.01 –0.300 0.059 –5.081
ECL –0.01 –0.224 0.042 –5.374
FI –0.01 –0.217 0.044 –4.958
SJM –0.01 –0.210 0.042 –5.018
WY –0.01 –0.245 0.044 –5.511
NXPI –0.005 –0.534 0.071 –7.476
GE 0.005 0.210 0.044 4.801
T 0.01 0.171 0.032 5.357
BKNG 0.01 0.210 0.043 4.903
COST 0.01 0.154 0.030 5.102
HIG 0.01 0.249 0.048 5.224
MCK 0.01 0.232 0.050 4.641
ROST 0.01 0.209 0.040 5.250
WRB 0.01 0.161 0.032 4.971
WM 0.01 0.192 0.038 5.028
$\hat{\mathcal{I}}$ 7.476
Table A5: Uniform test for the existence of jump effects. This table reports the selected threshold locations, the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. Only the significant results at the $1\%$ level are reported. The last line reports the test statistic of the uniform testing procedure, with critical values $4.611$ at $1\%$ level.
E.2 Application II: Party Incumbency Effects on U.S. House Elections
In the second application, we are interested in estimating the advantage a candidate has in the U.S. House elections when the seat is currently occupied by the party of the candidate. This is called the party incumbency effect. We follow the research design of Lee, (2008). See Caughey and Sekhon, (2011) for an analysis of more recent elections. The dependent variable $Y_{jt}$ is the democratic two-party vote share in year $j$ in district $t$. The coavariate $X_{jt}$ is the difference in the two-party vote share in the previous election. Since the winner of the election is deterministically determined by the 0 threshold, our model setup is appropriate.
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}>0\}}+e_{jt}.$
We want to test the one-sided null hypothesis $H_{0}^{(1)}:\gamma_{i}<0$ for all $i\in[N]$, versus the alternative hypothesis $H_{a}^{(1)}$: $\gamma_{i}\geq 0$ for some $i$. Unlike Lee, (2008) who pools the data together to estimate a single incumbency effect, we conduct a separate analysis for different election years and also a separate analysis for different states by interchanging the roles of $j$ and $t$ and allowing for possible heterogeneity of incumbency effects along either the $j$ or $t$ dimension. In addition, it is easy to see our theories extend to the unbalanced data in which case we can use $T_{j}$ to denote the number of observations associated with individual/group $j$.
We use the data provided by the MIT Election Data and Science Lab, (2017). The data include the vote shares of all candidates in the elections for the US House of Representatives from 1976–2020. Due to redistricting at the start of each decade, we have to exclude years ending with a ‘0’ or ‘2’.
[…]
Now, we are interested in the state-specific effects. We restrict our attention to states with a sufficient number of elections held in the sample. We choose a threshold of 50 elections. This gives us $N=29$ states in our analysis, and $T_{j}$ now denotes the number of elections included for the $j$th state. We argue that this setting constitutes a fairly high-dimensional setting since the effective number of observations is smaller than $N$ for a fairly large proportion of the states. The results are reported in Table A8. Consistent with the previous results, almost all of the estimated coefficients are positive. By using the critical values for the simultaneous tests, we find significant effects for seven states, for Missouri and California at $1\%$ level, and for Pennsylvania, Illinois, Ohio, Iowa, and South Carolina at $5\%$. Again, we can reject the null hypothesis of no jump effect for our uniform testing procedure at a $1\%$ level. The significant states tend to belong to those with the largest number of observations. However, for the other two largest states, New York and Texas, the estimated effect is rather small and far from being significant. The incumbency effect seems to be stronger in some states compared to others. We also run the test for the homogeneity of incumbency effects ($H_{0}^{\left(2\right)}$) over the 29 states. The results are displayed in Table A9 and the test statistic is given by $\hat{\mathcal{Q}}=5.159$, which implies that we can reject the null hypothesis of homogeneous effects at the $1\%$ level. We find heterogeneity in the effects across election periods and states.
Year $\hat{\gamma}_{j}$ $(T_{j}b_{j})^{-1/2}\hat{v}_{j}$ $(T_{j}b_{j})^{1/2}\hat{\gamma}_{j}/\hat{v}_{j}$ Obs EObs
1978 0.174 0.039 4.456 412 127
1984 0.163 0.028 5.826 413 154
1986 0.098 0.039 2.541 408 86
1988 0.161 0.048 3.327 395 79
1994 0.117 0.024 4.810 387 102
1996 0.037 0.022 1.645 394 112
1998 0.112 0.033 3.340 369 132
2004 0.045 0.071 0.633 359 65
2006 0.063 0.077 0.822 403 102
2008 0.063 0.023 2.788 405 142
2014 0.075 0.033 2.238 380 103
2016 0.036 0.046 0.798 386 42
2018 0.037 0.026 1.420 403 37
$\hat{\mathcal{I}}$ 5.826
Table A6: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.405$, $2.657$ and $3.165$, respectively.
Year $\hat{\gamma}_{j}-\hat{\gamma}_{0.5}$ $(T_{j}b_{j})^{-1/2}\hat{\tilde{v}}_{j}$ $(T_{j}b_{j})^{1/2}(\hat{\gamma}_{j}-\hat{\gamma}_{0.5})/\hat{\tilde{v}}_{j}$ Obs EObs
1978 0.083 0.038 2.630 412 127
1984 0.072 0.028 3.123 413 154
1986 0.008 0.038 0.631 408 86
1988 0.070 0.046 1.869 395 79
1994 0.026 0.025 1.661 387 102
1996 –0.054 0.024 –1.596 394 112
1998 0.021 0.033 1.122 369 132
2004 –0.046 0.066 –0.451 359 65
2006 –0.028 0.072 –0.160 403 102
2008 –0.028 0.024 –0.498 405 142
2014 –0.016 0.033 0.000 380 103
2016 –0.054 0.044 –0.876 386 42
2018 –0.054 0.027 –1.396 403 37
$\hat{\mathcal{Q}}$ 2.553 3.123
Table A7: Uniform test for heterogeneity. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure. Critical values at $10\%$, $5\%$ and $1\%$ level are $2.405$, $2.657$ and $3.165$.
State State_ic $\hat{\gamma}_{j}$ $(T_{j}b_{j})^{-1/2}\hat{v}_{j}$ $(T_{j}b_{j})^{1/2}\hat{\gamma}_{j}/\hat{v}_{j}$ Obs EObs
CONNECTICUT 1 0.142 0.054 2.645 72 41
MASSACHUSETTS 3 0.251 0.093 2.713 129 22
NEW JERSEY 12 0.092 0.044 2.082 171 55
NEW YORK 13 0.053 0.032 1.690 387 153
PENNSYLVANIA 14 0.096 0.032 2.984 237 100
ILLINOIS 21 0.114 0.039 2.922 250 85
INDIANA 22 0.080 0.029 2.720 125 62
MICHIGAN 23 0.067 0.036 1.874 206 79
OHIO 24 0.133 0.045 2.977 230 69
WISCONSIN 25 0.128 0.075 1.704 97 31
IOWA 31 0.125 0.040 3.154 64 50
MISSOURI 34 0.258 0.029 8.878 115 37
VIRGINIA 40 0.223 0.102 2.175 102 54
ALABAMA 41 0.212 0.070 3.020 74 21
FLORIDA 43 0.039 0.049 0.798 220 85
GEORGIA 44 0.101 0.048 2.101 126 40
LOUISIANA 45 –0.403 0.228 –1.764 80 16
MISSISSIPPI 46 0.101 0.095 1.057 51 13
NORTH CAROLINA 47 0.027 0.040 0.673 147 77
SOUTH CAROLINA 48 0.352 0.108 3.249 65 25
TEXAS 49 0.016 0.067 0.246 344 117
KENTUCKY 51 0.067 0.057 1.186 70 33
MARYLAND 52 0.126 0.134 0.937 105 29
TENNESSEE 54 0.049 0.059 0.828 101 20
ARIZONA 61 0.131 0.057 2.275 75 34
COLORADO 62 0.070 0.059 1.200 78 44
CALIFORNIA 71 0.088 0.025 3.458 639 189
OREGON 72 0.046 0.047 0.970 64 33
WASHINGTON 73 0.036 0.041 0.878 114 55
$\hat{\mathcal{I}}$ 8.878
Table A8: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.685$, $2.917$ and $3.392,$ respectively. Due to the low effective sample size, the bandwidth is chosen uniformly over all states.
State State_ic $\hat{\gamma}_{j}-\hat{\gamma}_{0.5}$ $(T_{j}b_{j})^{-1/2}\hat{\tilde{v}}_{j}$ $(T_{j}b_{j})^{1/2}(\hat{\gamma}_{j}-\hat{\gamma}_{0.5})/\hat{\tilde{v}}_{j}$ Obs EObs
CONNECTICUT 1 0.046 0.054 0.854 72 41
MASSACHUSETTS 3 0.155 0.090 1.714 129 22
NEW JERSEY 12 –0.004 0.045 –0.085 171 55
NEW YORK 13 –0.043 0.034 –1.271 387 153
PENNSYLVANIA 14 0.000 0.034 0.000 237 100
ILLINOIS 21 0.017 0.040 0.436 250 85
INDIANA 22 –0.016 0.032 –0.512 125 62
MICHIGAN 23 –0.029 0.037 –0.769 206 79
OHIO 24 0.037 0.045 0.810 230 69
WISCONSIN 25 0.031 0.074 0.427 97 31
IOWA 31 0.028 0.041 0.701 64 50
MISSOURI 34 0.161 0.031 5.159 115 37
VIRGINIA 40 0.127 0.100 1.268 102 54
ALABAMA 41 0.116 0.069 1.673 74 21
FLORIDA 43 –0.057 0.049 –1.157 220 85
GEORGIA 44 0.005 0.048 0.099 126 40
LOUISIANA 45 –0.499 0.221 –2.261 80 16
MISSISSIPPI 46 0.005 0.093 0.051 51 13
NORTH CAROLINA 47 –0.069 0.041 –1.682 147 77
SOUTH CAROLINA 48 0.256 0.106 2.427 65 25
TEXAS 49 –0.080 0.066 –1.215 344 117
KENTUCKY 51 –0.029 0.056 –0.513 70 33
MARYLAND 52 0.029 0.130 0.226 105 29
TENNESSEE 54 –0.048 0.058 –0.814 101 20
ARIZONA 61 0.034 0.057 0.603 75 34
COLORADO 62 –0.026 0.058 –0.446 78 44
CALIFORNIA 71 –0.009 0.028 –0.307 639 189
OREGON 72 –0.050 0.048 –1.048 64 33
WASHINGTON 73 –0.060 0.042 –1.431 114 55
$\hat{\mathcal{Q}}$ 5.159
Table A9: Uniform test for heterogeneity. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.685$, $2.917$ and $3.392,$ respectively. Due to the low effective sample size, the bandwidth is chosen uniformly over all states.
Appendix F Key Intuition on the Dependence Structure
Let $\epsilon_{t}=(\epsilon_{1t},\ldots,\epsilon_{Nt})^{\top}.$ The critical question is how the dependency structure in $\epsilon_{t}$ affects the construction of our test statistics. In Section 3 we allow $\left\{\epsilon_{jt}\right\}$ to exhibit some fairly general spatial and temporal dependency structures. One example is a moving average type of model along the $t$-dimension, i.e., $\epsilon_{t}=A_{0}\eta_{t}+A_{1}\eta_{t-1},$ where $A_{0}$ and $A_{1}$ are $N\times N$ matrices, $\eta_{t}=\left(\eta_{1t},\ldots,\eta_{Nt}\right)^{\prime}$ with $\eta_{jt}$’s being i.i.d. across $\left(j,t\right).$ If both $A_{0}$ and $A_{1}$ are diagonal, there is autocorrelation over $t$ but no cross-sectional correlation over $j$. But we allow for general cross-sectional dependence among $\epsilon_{jt}$’s so that neither $A_{0}$nor $A_{1}$ is restricted to be diagonal. Another example is an autoregressive structure along the $t$-dimension: $\epsilon_{t}=A_{0}\epsilon_{t-1}+\eta_{t},$ where, again, $A_{0}$ may or may not be diagonal.
Suppose that there is a cross-sectional dependence in the error term $\epsilon_{jt}$. Then the variance-covariance (VC) matrix of $\epsilon_{t}$ is non-diagonal. Does this have any implication on the choice of the critical value for the simultaneous inferences? In the usual GA setup, a simulated VC matrix is needed; see, e.g., Chernozhukov et al., (2019). Nevertheless, in our algorithm, we do not see any attempt to estimate the covariance of $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$, even though general cross-sectional dependency structures are allowed in $\epsilon_{t}$. This is mainly due to the localized nature of our kernel-type estimators $\hat{\gamma}_{j}$’s. In fact, we can show that under some mild conditions, as long as the joint density of $X_{it}$ and $X_{jt}$ is not degenerate for any $\left(i,j\right)$ pair (see Assumption 3.6 in Section 3), the covariance between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ will be smaller order compared to the variance. Therefore it is fine to standardize just by the standard deviation of $\hat{\gamma}_{j}$ or its consistent estimate without worrying about the covariance.
To understand why the localized statistics have the above favorable properties, we provide some intuitions by focusing on a simplified setting, where both the confounding variable $U_{jt}$ and the individual/group fixed effect $\alpha_{j}$ are absent from the structural model in (2.1), and $\sigma_{j}(X_{jt},U_{jt})=1.$ In this case, $e_{jt}=\epsilon_{jt}.$ For simplicity, we also assume for the moment that $\left(X_{it},\epsilon_{it}\right)$ is i.i.d. over time $t,$ and $X_{it}$’s are independent of $\epsilon_{it}$’s. Heuristically, the weighted statistic $\mbox{max}_{j}(bT)^{1/2}\sum_{t=1}^{T}(w_{jt,b}^{+}\epsilon_{jt}-w_{jt,b}^{-}\epsilon_{jt})$, is of the same order as a weighted statistic adopting a uniform kernel, viz., $\mbox{max}_{j}(bT)^{-1/2}$ $\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{jt}$. The high dimensional GA ensures that we can approximate $\mbox{max}_{j}(bT)^{-1/2}$ $\sum_{t=1}^{T}$ $\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{jt}$ by a maximum of Gaussian random vector with the same VC structure as $\{(bT)^{-1/2}$ $\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{jt}\}_{j=1}^{N}$. Thus the population counterpart of the covariance function is
$(Tb)^{-2}\sum_{t=1}^{T}\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{it}\epsilon_{jt}]=(Tb)^{-2}\sum_{t=1}^{T}\mathbb{E}(\epsilon_{it}\epsilon_{jt})\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}].$
We focus on the case with $\mathbb{E}(\epsilon_{it}\epsilon_{jt})\neq 0$ as cross-sectional dependence is our main concern here. Obviously, $\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}]=\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}]=O(b)$ if $i=j$, and $\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}]=O(b^{2})$ otherwise. Therefore the $i\neq j$ covariance terms are of smaller order than the $i=j$ variance terms and can be ignored asymptotically. In general, such order calculations hold true even when there exists strong cross-sectional correlation among $\left\{(X_{jt},\epsilon_{jt})\right\}_{j=1}^{N}$ as long as we do not have perfect correlation between $X_{it}$ and $X_{jt}$. So we will exclude the case of linear dependence, e.g., $X_{jt}=a_{0}+a_{1}X_{it}$ for some constants $a_{0}$ and $a_{1}\neq 0.$ We will also exclude the case of asymptotic linear dependence: $X_{jt}=a_{1}X_{it}+\eta_{it}$ with Var$(\eta_{it})\rightarrow 0$, which means that the variable $X_{it}$ and $X_{jt}$ are asymptotically linearly dependent. In our simulation study in Section 4, we can even allow strong cross-sectional dependence in $\left\{(X_{jt},\epsilon_{jt})\right\}_{j=1}^{N}$ that is generated via factor structures. As expected, as the cross-sectional dependence becomes stronger, we may lose some information in comparison with the independent case and a large sample size may be needed to achieve the same level of finite sample accuracy as the independent case.
The above explanation is in line with the kernel estimation in the time series literature. It is well known in the nonparametric time series literature that the asymptotic variance of the kernel estimator shares the same asymptotic variance as in the independent case, which means that the covariance terms in the variance calculation are asymptotically negligible under some weak dependence condition. See, e.g., Theorem 5.1 in Fan and Yao, (2003) and Theorems 18.2 and 18.4 in Li and Racine, (2007). Comparatively, our settings are slightly different, since we do not sum over the cross-sectional dimension $j$ and therefore we do not need to impose decay of dependence in the cross-sectional dimension. Even so, we see the covariance between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ is strictly dominated by the variance of $\hat{\gamma}_{j},$ regardless of the strong or weak cross-sectional dependence structure among $\left\{(X_{jt},\epsilon_{jt})\right\}_{j=1}^{N}.$