### v3

# Inference on many jumps in nonparametric panel regression models

Abstract
We investigate the significance of change-points within fully nonparametric regression contexts, with a particular focus on panel data where data generation processes vary across units, and error terms may display complex dependency structures. In our setting the threshold effect depends on one specific covariate, and we permit the true nonparametric regression to vary based on additional (latent) variables. We propose two uniform testing procedures: one to assess the existence of change-points and another to evaluate the uniformity of such effects across units. Our approach involves deriving a straightforward analytical expression to approximate the variance-covariance structure of change-point effects under general dependency conditions. Notably, when Gaussian approximations are made to these test statistics, the intricate dependency structures within the data can be safely disregarded owing to the localized nature of the statistics. This finding bears significant implications for obtaining critical values. Through extensive simulations, we demonstrate that our tests exhibit excellent control over size and reasonable power performance in finite samples, irrespective of strong cross-sectional and weak serial dependency within the data. Furthermore, applying our tests to two datasets reveals the existence of significant nonsmooth effects in both cases.
Keywords: high-dimensional time series; temporal and cross-sectional dependence; changepoint analysis; Gaussian approximation; simultaneous tests; threshold regression.
1 Introduction
Recently, there has been a notable surge in the study and application of changepoint analysis in high-dimensional time series data (see, e.g., Cho and Fryzlewicz, (2015), Cao et al., (2018), Chen et al., (2022), and Li et al., (2024), along with the references cited therein). Detecting changepoints in regression models is important and has wide applications in the field of biology (Pastor and Guallar, , 1998), agricultural sciences (Freeman et al., , 1998), econometrics (Li et al., , 2024), etc. The purpose of these studies is to estimate not only where the threshold effect exists but also the magnitude of the effects.
There is extensive literature on parametric high-dimensional changepoint regression analysis. Li et al., (2016), Manner et al., (2019) Wang et al., (2021), Rinaldo et al., (2021), and Cho et al., (2024) investigate changepoint analysis of coefficients in high-dimensional linear regression models using penalization methods. Lee et al., (2016) examine a high-dimensional regression model with a changepoint due to a covariate threshold in a cross-sectional study. These studies majorly focus on parametric models with high-dimensional covariates. Similar to Lee et al., (2016), this paper also considers changepoint regression due to a covariate threshold. However, our focus is on nonparametric regression with a high-dimensional changepoint effect driven by heterogeneity in the panel data structure. Specifically, we investigate the heterogeneous nonparametric panel data with jumps in the conditional mean functions:
$Y_{jt}=h_{j}(X_{jt},U_{jt})+\tau_{j}(X_{jt},U_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt},$ (1.1)
where $X_{jt}$ is a one-dimensional threshold variable, $Y_{jt}\in\mathbb{R}^{1}$ and $U_{jt}\in\mathbb{R}^{d}$ are the outcome and covariate variables that can be observed or unobserved for individual $j\in[N]:=\{1,\ldots,N\}$ at time point $t\in[T],$ $h_{j}(\cdot,\cdot)$, $\tau_{j}(\cdot,\cdot)$ and $\sigma_{j}(\cdot,\cdot)$ are continuous functions, $c_{0j}$ is a threshold value that is either known or unknown and $\epsilon_{jt}$ is the error term. Our main focus is to test whether changepoint effects related to $\tau_{j}(X_{jt},U_{jt})$ exist and to determine if the jump sizes are identical uniformly across individuals $j$ under general spatial temporal dependence assumptions.
As a motivating example for our model, we explore an application involving stock lagged returns and volatility using threshold regression. Threshold regression, dating back to Chan et al., (1985) and more recently studied in works such as Duffy et al., (2023), is mainly formulated within a linear parametric setting. Our approach naturally extends this framework to the non-parametric setting. Specifically, in our first application, revising the news impact curve of Engle and Ng, (1993), $Y_{jt}$ is the volatility of firm $j$’s stock at time $t$ and $X_{jt}$ is the lagged stock return. A trader would like to discover whether there are asymmetric effects of $X_{jt}$ on $Y_{jt}$, as well as whether and where a discontinuous relationship occurs, for trading and risk management purposes. If there are multiple stocks jump, are they all identical uniformly? In Section 5 in our empirical section, we identify significant threshold effects, demonstrating that our methodology is crucial for predicting stock volatility and effectively managing risk. In another application, we study the effect of party incumbency as determined by the 50% threshold in the two-party vote share in the US House Elections. Our findings confirm the presence of a significant positive incumbency effect as studied by Lee, (2008), and we detect significant heterogeneity in the threshold effect across the US states.
The methodological contributions are twofold. First, we develop algorithms to test for overall threshold effects, examine the potential heterogeneity effects and derive the asymptotic properties of the resulting test statistics. Due to the use of local smoothing, neither the weak dependence along the time dimension nor the weak or strong dependence along the individual/group dimension enters the asymptotic distribution of the final statistic. Interestingly, we find strong cross-sectional dependence is allowed in our error terms. As a practical recommendation, we find that under fairly weak assumptions, if there are clustering dependency patterns in $e_{jt}$, the correlation in the estimated statistics can be safely ignored, and only adjustments for heteroskedasticity are needed. Second, the standard central limit theorem is challenging to apply to high-dimensional time series when $N$ significantly exceeds the effective sample size. Previously, Chernozhukov et al., (2015); Chernozhukov et al., 2017a , Zhang and Wu, (2017), Chernozhukov et al., (2019), and others addressed this issue by introducing high-dimensional Gaussian approximation (GA) methods. However, the existing literature cannot address our problem. The key difficulty is to incorporate possibly non-identically distributed time series under general spatial and temporal dependency structures and more importantly latent variables. In this paper, we propose a new high dimensional GA for weighted statistics that address all these issues. These GA results are essential for determining the critical values of our supremum test.
Our paper is related to three strands of literature in statistics. First, we add to the extensive body of work on threshold regression for high dimensional data. For example, Miao et al., 2020a and Miao et al., 2020b consider the estimation of a panel threshold regression model with interactive fixed effects (IFEs) and latent group structures, respectively; Massacci and Trapani, (2022) study inference for high-dimensional threshold regressions with common stochastic trends. The aforementioned literature does not account for potential heterogeneous threshold effects. Ignoring these effects may bias estimators or lead to efficiency loss when estimating homogeneous threshold effects in finite samples. Second, our paper contributes to the literature on testing homogeneity in a panel data setting. For example, Phillips and Sul, (2003), Pesaran and Yamagata, (2008) and Su and Chen, (2013) propose tests for slope coefficient heterogeneity in a linear regression model with or without cross-sectional dependence. In the presence of slope heterogeneity, Su et al., (2016) and Su and Ju, (2018) model the slope heterogeneity via latent group structures. Barassi et al., (2023) study a heterogeneous panel threshold regression model with IFEs. Above literature all focus on parametric setting. Our work complements these research streams by adopting nonparametric approaches, which address size and power distortions caused by model misspecifications. Third, our nonparametric approach is related with literature in trends estimation. Inference on smooth trend functions in nonparametric models for time series data is a well-established field; see Zhang and Wu, (2012), Karmakar et al., (2022), Gao et al., (2024), Chen and Wu, (2018), among others. Distinctly, while the above literature examines nonlinear dynamic models, they does not address discontinuity in trends, particularly heterogeneous jumps. The nonparametric literature on jumps in trends mostly focuses on independent and identically distributed (i.i.d.) settings, without accounting for heterogeneity or high dimensionality; see, for example, Qiu, (1998), Müller and Stadtmüller, (1999), and Spokoiny, (1998).
Another distinct feature is that we explicitly handle the case with missing covariates or models with misspecified variables. Theoretically, validating such an inference procedure requires working with a reduced-form model that has specific variance-covariance structures. This aspect of theoretical validity has not been explored in the time-varying trend literature. Establishing the validity of uniform inference is a major theoretical contribution that sets our work apart in the study of varying trends.
The rest of the paper is organized as follows. Section 2 provides the model setup and the practical steps of the testing procedure. Section 3 presents the assumptions and delivers the main theorems. In Section 4 we provide simulation results. [^1] The code for our method is available in the R package hdthreshold (https://cran.r-project.org/web/packages/hdthreshold/index.html), with an illustration accessible at https://rpubs.com/Lk1110/1259698. Real data applications are discussed in Section 5, and the paper concludes in Section 6. The Supplementary Material includes additional theorems, proofs of all results, extra simulation outcomes, and tables for the empirical applications.
Notation. For a vector $v=(v_{1},\ldots,v_{d})\in\mathbb{R}^{d}$ and $q>0$, we denote $|v|_{q}=(\sum_{i=1}^{d}|v_{i}|^{q})^{1/q}$ and $|v|_{\infty}=\max_{1\leq i\leq d}|v_{i}|$. For a matrix $A=(a_{i,j})_{1\leq i\leq m,1\leq j\leq n}$, we define the max norm $|A|_{\text{max}}=\max_{i,j}|a_{i,j}|$. For $s>0$ and a random vector $X$, we say $X\in\mathcal{L}^{s}$ if $\lVert X\rVert_{s}=[\mathbb{E}(|X|^{s})]^{1/s}<\infty$. For two positive number sequences $(a_{n})$ and $(b_{n})$, we say $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$ (resp. $a_{n}\asymp b_{n}$) if there exists $C>0$ such that $a_{n}/b_{n}\leq C$ (resp. $1/C\leq a_{n}/b_{n}\leq C$) for all large $n$, and say $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\rightarrow 0$ as $n\rightarrow\infty$. We set $(X_{n})$ and $(Y_{n})$ to be two sequences of random variables. Write $X_{n}=O_{\mathbb{P}}(Y_{n})$ if for any $\epsilon>0$, there exists $C>0$ such that $\mathbb{P}(X_{n}/Y_{n}\leq C)>1-\epsilon$ for all large $n$, and say $X_{n}=o_{\mathbb{P}}(Y_{n})$ if $X_{n}/Y_{n}\rightarrow 0$ in probability as $n\rightarrow\infty$.
2 Model Setup and Test Procedure
In this section, we present the model, the hypotheses and estimators, and the test procedure. Subsections 2.1 and 2.2 are concerned with the known threshold case. In Subsection 2.3 we focus on the case in which the threshold locations $c_{0j}$ are unknown. The algorithms for our testing procedures are listed in Subsection 2.4.
2.1 Model Setup
In this subsection, we formulate our model. Recall the full model in (1.1) where $U_{jt}\in\mathbb{R}^{d}$ is a latent random vector that can be arbitrarily correlated with $X_{jt},$ $\epsilon_{jt}$ is a mean zero term that is independent of $\left(X_{jt},U_{jt}\right)$.
Without loss of generality set $c_{0j}=0$ for the known case. The full model in (1.1) can be rewritten into the following reduced form model:
$Y_{jt}=\tilde{h}_{j}(X_{jt})+\tilde{\tau}_{j}(X_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt},$ (2.1)
where $\tilde{\tau}_{j}(X_{jt})=\mathbb{E}[\tau_{j}(X_{jt},U_{jt})|X_{jt}]$, $\tilde{h}_{j}(X_{jt})=\mathbb{E[}h_{j}(X_{jt},U_{jt})|X_{jt}]$, $\varepsilon_{j}(X_{jt},U_{jt})=h_{j}(X_{jt},U_{jt})-\tilde{h}_{j}(X_{jt})+[\tau_{j}(X_{jt},U_{jt})-\tilde{\tau}_{j}(X_{jt})]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$ and the reduced form error
$e_{jt}=\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}.$ (2.2)
The noise term $e_{jt}$ has a complex structure, influenced by both the latent variable and by the heterogeneous volatility of the observations. This intricate structure, combined with the temporal and cross-sectional dependencies of the noise, makes the inference process particularly challenging.
A few remarks are in order. First, we have $\mathbb{E}(e_{jt}|X_{jt})=0$ and the conditional expectation of the observed outcome given $X_{jt}$ is $\mathbb{E}(Y_{jt}|X_{jt})=\tilde{h}_{j}(X_{jt})+\tilde{\tau}_{j}(X_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}.$ Second, in Section G.2 we show the possibilities of extending our model to add more covariates and incoporating fixed effects. Third, ($X_{jt},U_{jt}$) can be lagged variables. For example, in our application on modeling stock volatility (Section 5.1), $X_{jt}=Y_{j(t-1)}$, functions $h_{j}(\cdot,\cdot)$, $\tau_{j}(\cdot,\cdot)$ and $\sigma_{j}(\cdot,\cdot)$ represent baseline effect, the jump effect and the volatility function, respectively. The $U_{jt}$ are unobserved risk factors or variables we exclude in the nonparametric regression. Modeling $U_{jt}$ is important in this circumstance as the volatility of a single stock is most likely affected by other risk factors such as news events or unobserved market factors.
This paper aims to explore methods for conducting simultaneous inference on threshold effect estimators in a large $N$ and large $T$ setup. In particular, we allow $N\gg T.$ Our focus is on the threshold effect, starting with the case where the threshold value $c_{0j}$ is known and then extending the analysis to the unknown case in Subsections 2.3 and 3.3.
Assuming that $X_{jt}$ is a continuous random variable, the jump effects can be identified as a “gap” in the conditional expectations and their derivatives at the thresholds. Let $\mu_{j}(x):=\mathbb{E}(Y_{jt}|X_{jt}=x)$ and $\mu_{j}(c_{0j}-)=\lim_{x\uparrow c_{0j}}\mu_{j}(x)$ and $\mu_{j}(c_{0j}+)=\lim_{x\downarrow c_{0j}}\mu_{j}(x).$ Consider $\partial_{+}\mu_{j}(\cdot)$ (resp. $\partial_{-}\mu_{j}(\cdot)$) as the right (resp. left) derivative of function $\mu_{j}(\cdot).$ The threshold effect for the individual $j$ is defined as
$\gamma_{j}$ $=$ $\mu_{j}(c_{0j}+)-\mu_{j}(c_{0j}-),$ (2.3)
$\gamma_{j}^{[1]}$ $=$ $\partial_{+}\mu_{j}(c_{0j})-\partial_{-}\mu_{j}(c_{0j}).$ (2.4)
Due to continuity of function $h_{j}(\cdot,\cdot)$ and $\tau_{j}(\cdot,\cdot)$, we have $\gamma_{j}=\tilde{\tau}_{j}(c_{0j}),\gamma_{j}^{[1]}=\tilde{\tau}^{\prime}_{j}(c_{0j}).$ ($\tilde{\tau}^{\prime}_{j}(.)$ denotes the derivative of $\tilde{\tau}_{j}(.)$.)
2.2 Hypotheses and Estimators
In this section, we introduce the hypotheses and estimators. We would like to conduct simultaneous inferences on $\gamma_{j}$ and $\gamma_{j}^{[1]}$ for $j\in\left[N\right]$. The importance of studying individual/group-specific threshold effects is discussed in Zimmert and Lechner, (2019).
First, we are interested in testing the existence of the threshold effects ($\gamma_{j}$ or $\gamma_{j}^{[1]}$). The null and alternative hypotheses are defined as follows:
$H_{0}^{\left(1\right)}:\gamma_{j}=0\ \forall\text{ }j\in\left[N\right],\quad\text{and }H_{a}^{\left(1\right)}:\gamma_{j}\neq 0\ \text{for some }j\in\left[N\right].$ (2.5)
The motivation for considering such a test is crucial. In Figure 1, we present a simple comparison with the pooled threshold test in the literature, for example in Fong et al., (2017). It is clear that the conventional pooled tests exhibit very low power due to signal cancellation, with performance only comparable in cases of strictly positive signals.
Figure 1: Empirical power comparison between our uniform testing procedure (black) for the existence of threshold effects and the test based on the pooled linear threshold model (red) at $5\%$ significance level for different proportions of non-zero coefficients ($\gamma_{j}$) on the x-axis. The non-zero $\gamma_{j}$s are drawn from a standard normal distribution (left panel) and an exponential distribution (right panel). The true process is linear, $f_{j}(x)=0.2x$ (for details see DGP7 in Section E of the supplement.)
When $N$ is large, assuming homogeneous threshold effects $\gamma_{j}$ across all $j$ is restrictive, and inferences based on this assumption can be misleading if the threshold effects are heterogeneous. Therefore, it is important to test for heterogeneous threshold effects. In this case, the null and alternative hypotheses are:
$H_{0}^{\left(2\right)}:\gamma_{j}=\gamma_{0}\ \text{for some }\gamma_{0},\text{ }\forall\text{ }j\in\left[N\right],\quad$
and $H_{a}^{\left(2\right)}:\text{there is no constant }\gamma_{0}\text{ such that }\gamma_{j}=\gamma_{0},\text{ }\forall\ j\in\left[N\right].$
We can similarly define $H_{0}^{(1)[1]}$ and $H_{0}^{(2)[1]}$ for the existence and homogeneity tests of the derivative, respectively. Rejecting of $H_{0}^{\left(2\right)}$ suggests the presence of heterogeneous threshold effects. Moreover, providing simultaneous confidence intervals for the threshold effects is also tempting. If we would like to study the asymptotics of $\hat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{N})^{\top}$, the standard central limit theorem may fail due to the high dimensionality $N$. One of the key theoretical contributions is to provide a general framework for making uniform inferences on heterogeneous threshold effects.
To estimate $\gamma_{j}$, we need to estimate both $\mu_{j}(c_{0j}+)$ and $\mu_{j}(c_{0j}-).$ To this aim, it is natural to adopt local linear estimators, see for example Fan, (2018). Denote the estimators:
$\big{(}\hat{\mu}_{j}(c_{0j}+),{\partial_{+}\hat{\mu}_{j}}(c_{0j})\big{)}=$ $\mathrm{argmin}_{\beta_{j,0},\beta_{j,1}}\sum_{t=1}^{T}(Y_{jt}-\beta_{j,0}-\beta_{j,1}X_{jt})^{2}K_{jt}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}},$ (2.6)
$\big{(}\hat{\mu}_{j}(c_{0j}-),{\partial_{-}\hat{\mu}_{j}}(c_{0j})\big{)}=$ $\mathrm{argmin}_{\beta_{j,0},\beta_{j,1}}\sum_{t=1}^{T}(Y_{jt}-\beta_{j,0}-\beta_{j,1}X_{jt})^{2}K_{jt}\mathbf{1}_{\{X_{jt}<c_{0j}\}},$ (2.7)
where $K_{jt}:=K((X_{jt}-c_{0j})/b_{j}),$ $b_{j}$ is a bandwidth parameter, and $K(\cdot)$ is a kernel function. For $l=0,1,2$, define the weights as
$w_{j,b}^{+}(u,v)$ $=\frac{K((v-u)/b_{j})[S_{j2,b}^{+}(u)-(v-u)S_{j1,b}^{+}(u)]}{S_{j2}^{+}(u)S_{j0}^{+}(u)-S_{j1,b}^{+2}(u)}\mathbf{1}_{\{v\geq u\}},$ (2.8)
$w_{j,b}^{+[1]}(u,v)$ $=\frac{K((v-u)/b_{j})[S_{j1,b}^{+}(u)-(v-u)S_{j0,b}^{+}(u)]}{-S_{j2}^{+}(u)S_{j0}^{+}(u)+S_{j1,b}^{+2}(u)}\mathbf{1}_{\{v\geq u\}}.$ (2.9)
where
$S_{jl,b}^{+}(u)$ $=\sum_{t=1}^{T}(X_{jt}-u)^{l}K((X_{jt}-u)/b_{j})\mathbf{1}_{\{X_{jt}\geq u\}},$ (2.10)
and $w_{j,b}^{-}(u,v)$ (resp. $S_{jl,b}^{-}(u)$, $w_{j,b}^{-[1]}(u,v)$) is the same as $w_{j,b}^{+}(u,v)$ (resp. $S_{jl,b}^{+}(u)$, $w_{j,b}^{+[1]}(u,v)$) with $+$ and $\geq$ therein replaced by -- and <,<,< , respectively.
We can thus evaluate the threshold effect near the cut-off point and define our local linear estimator of the average threshold effect for individual/group $j$ as
$\hat{\mu}_{j}(c_{0j}+)=\sum_{t=1}^{T}w_{jt,b}^{+}Y_{jt}\quad\textrm{and}\quad{\partial_{+}\hat{\mu}_{j}}(c_{0j})=\sum_{t=1}^{T}w_{jt,b}^{+[1]}Y_{jt},$ (2.11)
where the weights take the following form $w_{jt,b}^{+}=w_{j,b}^{+}(c_{0j},X_{jt})$ and $w_{jt,b}^{-}=w_{j,b}^{-}(c_{0j},X_{jt})$. Thus we have,
$\hat{\gamma}_{j}=\sum_{t=1}^{T}w_{jt,b}^{+}Y_{jt}-\sum_{t=1}^{T}w_{jt,b}^{-}Y_{jt}\quad\mbox{and}\quad\hat{\gamma}_{j}^{[1]}=\sum_{t=1}^{T}w_{jt,b}^{+[1]}Y_{jt}-\sum_{t=1}^{T}w_{jt,b}^{-[1]}Y_{jt}.$
We shall focus on $\hat{\gamma}_{j}$ from now on, and the algorithms and theorems concerning the statistical properties of $\hat{\gamma}_{j}^{[1]}$ are developed in Section A.5 in the Appendix.
Under $H_{0}^{\left(1\right)}:$ $\gamma_{j}=0$ for all $j\in\left[N\right]$, we are exclusively testing for the overall significance of threshold effects. Since the functions $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$ are continuous, by (2.1), the difference between our estimator and the parameter can be approximately expressed in terms of a weighted average of the error sequence,
$\hat{\gamma}_{j}\approx\gamma_{j}+\sum_{t=1}^{T}(w_{jt,b}^{+}e_{jt}-w_{jt,b}^{-}e_{jt}),$ (2.12)
where $e_{jt}$ is the composite error in (2.1). Under $H_{0}^{\left(1\right)}$, one would expect $\left|\hat{\gamma}_{j}\right|$ to be small for all $j\in\left[N\right]$. This motivates us to consider the test statistic
$\mathcal{I}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}/v_{j}|,$ (2.13)
where
$v_{j}^{2}=v_{j}^{2}(c_{0j})=(Tb_{j})\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}$ (2.14)
is the conditional variance of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$ for standardization purposes. For now, we assume $v_{j}^{2}$ is given, in general, it is unknown and has to be replaced by its estimate. We discuss a consistent estimator of $v_{j}^{2}$ in Subsection 2.4. Under the assumptions in Section 3, it can be shown that all asymptotic results based on $v_{j}^{2}$ remain valid when it is replaced by the consistent estimate $\hat{v}_{j}^{2}$.
Similarly, under $H_{0}^{\left(2\right)}$, one would expect $\left|\hat{\gamma}_{j}-\overline{\hat{\gamma}}\right|$ to be small for all $j\in\left[N\right],$ where $\overline{\hat{\gamma}}=\frac{1}{N}\sum_{j=1}^{N}\hat{\gamma}_{j}$. This motivates us to consider the test statistic
$\mathcal{Q}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}-\overline{\hat{\gamma}}|/\tilde{v}_{j},$ (2.15)
where $\tilde{v}_{j}^{2}=(1-1/N)^{2}v_{j}^{2}+\sum_{i\neq j}^{N}v_{i}^{2}/N^{2}$. For practical implementation, we introduce a feasible version of $\mathcal{Q}$ in Algorithm SA-2 in Section D of the Supplementary Material.
2.3 The Unknown Threshold Case
It shall be noted that in case the breakpoints $c_{0j}$’s are not known, we shall develop an algorithm to estimate them. We therefore generalize our methodology to estimate unknown breakpoints that vary over ${\normalsize j}$. To this end, we search over a grid $[c_{1},c_{2},\ldots,c_{K}]$ for each individual $j\in[N]$. Let
$w_{jt,b}(c_{i})=w_{j,b}^{+}(c_{i},X_{jt})-w_{j,b}^{-}(c_{i},X_{jt})\quad\mbox{and}\quad\hat{\gamma}_{j}(c_{i})=\sum_{t=1}^{T}w_{jt,b}(c_{i})Y_{jt}.$ (2.16)
Further, let $v_{j}^{2}(c_{i})=(Tb_{j})\sum_{t=1}^{T}w_{jt,b}^{2}(c_{i})\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}$ and define
$\mathcal{I}^{C}=\max_{1\leq i\leq K}\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/v_{j}(c_{i})|.$ (2.17)
We propose to estimate $c_{0j}$ by $\hat{c}_{j}:=\arg\max_{1\leq i\leq K}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$for $j\in\left[N\right],$ where $\hat{v}_{j}(c_{i})$ is a consistent estimator of $v_{j}(c_{i})$ which will be introduced in Algorithm SA-3 in Section D of the Supplementary Material.
2.4 Test Procedure
In this subsection, we present the steps of the test procedures. To proceed, we begin by discussing the estimation of the error variance $\sigma_{e,j}^{2}$ of $e_{jt}$ when $c_{0j}$’s are known.
  1. 1.
    Recall the definition of $\hat{\gamma}_{j}$ as in Equation (2.2). Regress $Y_{jt}-\hat{\gamma}_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$ on $X_{jt}$ using the local linear method with kernel $K\left(\cdot\right)$ and bandwidth $b_{j}$.
  2. 2.
    Calculate the regression residuals $\hat{e}_{jt}$.
  3. 3.
    Estimate $\sigma_{e,j}^{2}$ by
    $\hat{\sigma}_{e,j}^{2}=\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}\hat{e}_{jt}^{2}}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|\leq b_{j}\}}}.$ (2.18)
The estimator $\hat{\sigma}_{e,j}^{2}$ is used to form $\hat{v}_{j}^{2}$ in the Algorithms 1 and SA-2 in the case of known $c_{0j}$, which are applied to conduct simultaneous tests for $H_{0}^{\left(1\right)}$ and $H_{0}^{\left(2\right)}$. These algorithms are supported by the theoretical GA results in Section 3 and Section A in Supplementary Material. The following Algorithm 1 is for testing $H_{0}^{(1)}$.
Algorithm 1 Testing for existence of jumps in conditional means (in slope): i.e., $\gamma_{j}=0$ ( $\gamma_{j}^{[1]}=0$) for all $j$.
1:for $j=1,\ldots,N$ do
2:     Estimate $\gamma_{j}$ for each group/individual $j$ using equation (2.2).
3:     Estimate the variance $v_{j}^{2}$ of $(Tb_{j}^{3})^{1/2}\hat{\gamma}_{j}$ by $\hat{v}_{j}^{2}=Tb_{j}\sum_{t=1}^{T}\left(w_{jt,b}^{+}-w_{jt,b}^{-}\right)^{2}\hat{\sigma}_{e,j}^{2}$ ( $\hat{v}_{j}^{2[1]}=Tb_{j}^{3}\sum_{t=1}^{T}\left(w_{jt,b}^{+[1]}-w_{jt,b}^{-[1]}\right)^{2}\hat{\sigma}_{e,j}^{2}$) , where $\hat{\sigma}_{e,j}^{2}$ is a consistent estimator for the error variance using the observations close to the threshold $c_{0j}$ as in (2.18).
4:end for
5:Calculate the feasible test statistic ^$\hat{\mathcal{I}}:=\mbox{max}_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}|/\hat{v}_{j}$ ( ^$\hat{\mathcal{I}}^{[1]}:=\mbox{max}_{1\leq j\leq N}(Tb_{j}^{3})^{1/2}|\hat{\gamma}_{j}^{[1]}|/\hat{v}_{j}^{[1]}$) .
6:Reject the null hypothesis if ^$\hat{\mathcal{I}}>q_{\alpha}$ ( ^$\hat{\mathcal{I}}^{[1]}>q_{\alpha}$), where $q_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}|Z_{j}|$ with $Z_{j}$’s being independent standard normal variables.
[…]
Given the well-documented cross-sectional and serial dependence in high dimensional data, it is crucial to develop tests that account for both types of dependence. It is surprising that, according to the above algorithm, $q_{\alpha}$ can be obtained by simulating only i.i.d. Gaussian random variables, without considering the complicated dependency structure of $e_{jt}$. Recall from (2.1), we see that the reduced-form error $e_{jt}$ inherits nontrivial dependency structure in $\epsilon_{jt}$. However, this dependence does not affect the calculation of critical values, even in the case of fairly strong cross-sectional dependence (e.g., a factor structure). This is mainly due to the localized nature of our kernel-type estimators $\hat{\gamma}_{j}$’s in our test statistic. In fact, we can show that under some mild conditions, as long as the joint density of $X_{it}$ and $X_{jt}$ is not degenerate for any $\left(i,j\right)$ pair (see Assumption 3.3 in Section 3), the covariance terms between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ will be of lower order compared to the variance terms, $O(1/T)$. For a more detailed discussion on the key intuition on the impact of the dependence structure in a simplified setting, we refer to Section G.1 of the Supplementary Material. For a discussion on the issue of bandwidth selection, we refer to Remark 4 in Section A.2 in the Supplementary Material.
3 Main Theorems
In this section, we present our main results, with a particular focus on Theorem 3.1, which provide Gaussian approximation results that determine the critical values for testing the existence of threshold effects. Theorem 3.2 extends the result to unknown threshold case and Theorem 3.3 shows the consistency of the estimated threshold locations. Theorem A.1 shows GA for testing homogeneity. Theorems A.2 and A.4 derive the consistency of variance to ensure the validity of our test statistics. Theorem A.3 analyzes the power of the tests for both the threshold effects and homogeneity. Results regarding the derivative cases are presented in Subsection A.5.
3.1 Assumptions
In this subsection, we state the assumptions for the asymptotic analyses in Section 3.1 and Section A.1 in the Supplementary Material. First of all, we will pose a general spatial and temporal dependence assumption on the data generating processes. As for the covariate variable $X_{jt}$ and the confounding variable $U_{jt}\in\mathbb{R}^{d}$, we assume that they take the form
$X_{jt}=H_{j}(e_{t},e_{t-1},\ldots)\quad\text{and}\quad U_{jt}=\tilde{H}_{j}(X_{jt},\tilde{e}_{t}),$ (3.1)
where $\tilde{H}_{j}=(\tilde{H}_{j1},\ldots,\tilde{H}_{jd})^{\top}$, $\{\tilde{H}_{ji}\}_{j\in[N],i\in[d]}$ and $\left\{H_{j}\right\}_{j\in[N]}$ are functions so that $X_{jt}$ and $U_{jt}$ are well defined. For $t\in\mathbb{Z}$, $e_{t}\in\mathbb{R}$ denote i.i.d. innovations, and $\tilde{e}_{t}\in\mathbb{R}^{\tilde{d}}$ are i.i.d. random vectors with some constant $\tilde{d}>0$, independent of $\left\{e_{t}\right\}_{t\in\mathbb{Z}}$. Denote
$\mathcal{F}_{t}=(e_{t},e_{t-1},\ldots)\quad\textrm{and}\quad\tilde{\mathcal{F}}_{t}=(\tilde{e}_{t},\tilde{e}_{t-1},\ldots).$ (3.2)
For the innovation part $\epsilon_{t}=(\epsilon_{t1},\ldots,\epsilon_{tN})^{\top}$ as in (1.1), the dependence among $\epsilon_{t}:=(\epsilon_{1t},\ldots,\epsilon_{Nt})^{\top}$is allowed to be strong (e.g., with a factor structure) along the cross-sectional dimension and it has to be weak along the time dimension, as we only exploit the sample size in the time series dimension for the estimation of heterogeneous threshold effects. Specifically, we assume that $\epsilon_{t}$ follows an MA($\infty$) process as follows:
$\epsilon_{t}=\sum_{k\geq 0}A_{k}\eta_{t-k},$ (3.3)
where $\eta_{t}=(\eta_{1t},\ldots,\eta_{\tilde{N}t})^{\top}\in\mathbb{R}^{\tilde{N}},$ $t\in\mathbb{Z},$ are i.i.d. random vectors with zero mean and identity covariance matrix $I_{\tilde{N}}$, and $A_{k}\in\mathbb{R}^{N\times\tilde{N}}$ are real-valued matrices with $\tilde{N}\leq cN$ for some constant $c>0$. In particular, we allow $\tilde{N}=N$ so that the $A_{k}$’s become square matrices. We assume that $\left\{e_{t},\tilde{e}_{t}\right\}_{t\in\mathbb{Z}}$ are independent of $\left\{\eta_{t}\right\}_{t\in\mathbb{Z}}.$
Remark 1.
(Dependence between $X_{jt}$ and $U_{jt}$.) Note that Equation (3.1) is a fairly general setup, where almost all kinds of dependence structures between $X_{jt}$ and $U_{jt}$ can be fulfilled for the following simple reasoning. To see this, take $U_{jt}$ as a scalar random variable as an example. For any two continuous random variables $X$ and $U$ with the conditional cumulative distribution function of $U$ given $X$ as $F_{U|X}(\cdot|X)$. Let $\tilde{e}=F_{U|X}(U|X)$. Then by Lemma F.1 in Appendix F in Rio, (2017), $\tilde{e}$ is a uniformly distributed random variable on $[0,1]$ and is independent of $X$. Moreover, $U=F_{U|X}^{-1}(\tilde{e}|X).$            $\Box$
Next, we impose some conditions on the elements of $\eta_{t}.$
Assumption 3.1.
(Moment) $\eta_{jt}$’s are i.i.d. with zero mean and unit variance across $j$ and $t$. Either one of the following two conditions is satisfied:
  1. (i)
    (Finite moments) For some constant $q>2,$ $\eta_{11}$ has finite $q$th moment.
  2. (ii)
    (Sub-exponential) For some constant $\lambda_{0}>0$, $\eta_{1}=(\eta_{11},\ldots,\eta_{\tilde{N}1})^{\top}$ is sub-exponential with $a_{0}:=\sup_{|v|_{2}\leq\lambda_{0}}\mathbb{E}(e^{|v^{\top}\eta_{1}|})<\infty.$
The following assumption imposes some conditions on the temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}.$ Denote $\mathcal{C}_{j}$ as a set of search values for the time series $j$ containing at least the threshold values $c_{0j}$. If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in subsections 3.2 and A.3. We will specify $\mathcal{C}_{j}$ later in the theorems for the unknown case. Let $g_{j}(\cdot)$ be the density function of $X_{jt}.$ Recall $\mathcal{F}_{s}$ in (3.2), for $t\geq s+1,$ let $g_{j,t}(x|\mathcal{F}_{s}):=d\mathbb{P}(X_{jt}\leq x|\mathcal{F}_{s})/dx.$
Assumption 3.2.
(Dependence) (i) For any $x\in\mathcal{C}_{j},$ assume $g_{j,t}(x|\mathcal{F}_{t-1})$ has finite $p$th moment for some $p>2$. Denote
$\theta_{k,p}=\Big{\|}\sup_{1\leq j\leq N,1\leq t\leq T,x\in\mathcal{C}_{j}}\big{|}g_{j,t}(x|\mathcal{F}_{t-1})-g_{j,t}(x|\mathcal{F}_{t-1,\{t-1-k\}})\big{|}\Big{\|}_{p}.$
Assume $\sup_{m\geq 0}m^{\alpha}\sum_{k\geq m}\theta_{k,p}<\infty$ for some $\alpha>0.$
(ii) For some $C>0$ and $\beta>0,$ we have $\max_{1\leq j\leq N}\sum_{k\geq m}|A_{k,j,\cdot}|_{2}\leq C(m\vee 1)^{-\beta},\quad$where $m\geq 0$ and $A_{k,j,\cdot}\in\mathbb{R}^{\tilde{N}}$ is the $j$th row of matrix $A_{k}.$
Assumption 3.2 $(i)$-$(ii)$ essentially assumes weak temporal dependence for the processes $(X_{jt})_{t\in\mathbb{Z}}$ and $(\epsilon_{jt})_{t\in\mathbb{Z}}$ for any $j\in\left[N\right]$. We allow strong cross-sectional dependence over the $j$ dimension. We shall also note that $\beta$ indicates the strength of the dependency structure of the error processes.
Furthermore, let $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1})$ be the conditional joint density of $X_{j_{1}t}$ and $X_{j_{2}t}$ given $\mathcal{F}_{t-1}.$ Finally, the following assumption imposes a condition on $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1}).$
Assumption 3.3.
(Joint density) There is no perfect or asymptotically perfect collinearity in $X_{j_{1}t}$ and $X_{j_{2}t}$ for any $j_{1}\neq j_{2}.$ The conditional densities $g_{j_{1},j_{2}}(x_{1},x_{2}|\mathcal{F}_{t-1})$ are uniformly upper bounded, that is, $\max_{1\leq j_{1},j_{2}\leq N}\sup_{x_{1},x_{2}\in\mathbb{R}}|g_{j_{1},j_{2}}(x_{1},x_{2}|\mathcal{F}_{t-1})|<\infty.$
Let
$\bar{b}=\max_{1\leq j\leq N}b_{j}\quad\text{and}\quad\underline{b}=\min_{1\leq j\leq N}b_{j}.$ (3.4)
Assumptions SA-1-SA-3 in Section A.1 in the Supplementary Material are concerning standard assumptions on boundedness, smoothness and kernel functions involved.
3.2 Gaussian Approximation Results
If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in Subsections 3.2 and A.3. We first consider the GA result for $\mathcal{I}$ defined in (2.13). Define $d_{j}=(Tb_{j})^{1/2}\gamma_{j}/v_{j}$ and $\underline{d}=(d_{1},d_{2},\ldots,d_{N})^{\top}$. Further define $Z$ as a centered Gaussian random vector with identity covariance matrix. Under the null $H_{0}^{\left(1\right)},$ the bias term $\underline{d}$ becomes a zero vector. The following theorem states that the infeasible statistic $\mathcal{I}$ can be approximated well by the maximum of Gaussian random variables. Denote
$\Delta=\mathrm{log}^{7/6}(NT)(\underline{b}T)^{-1/6}+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}+T^{-(\alpha p)\wedge(p/2-1)}+\bar{b}^{1/3}\mathrm{log}^{2/3}(N).$
Define $\mathcal{R}_{NT}=\Delta+(\underline{b}T^{1-2/q})^{-1/3}\mathrm{log}(NT)+\mathrm{log}(NT)N^{1/q}T^{-\beta}$ under Assumption 3.1$(i)$, and $\mathcal{R}_{NT}=\Delta+\mathrm{log}(NT)T^{-\beta}$ under Assumption 3.1$(ii)$.
Theorem 3.1 (Gaussian Approximation 1).
Let Assumptions 3.1-3.3 and SA-1-SA-3 hold. Then we have
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}\leq u)-\mathbb{P}(|Z+\underline{d}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{NT}.$ (3.5)
Now we list the related rate requirement on $T$, $\bar{b}$ and $N$ under different moment conditions. Under Assumption 3.1$(i)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $\underline{b}^{-1}T^{2/q-1}\mathrm{log}^{3}(NT)\rightarrow 0$ and $\mathrm{log}(NT)N^{1/q}T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$. Under Assumption 3.1$(ii)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $(T\underline{b})^{-1}\mathrm{log}^{7}(NT)\rightarrow 0,$ $\bar{b}\mathrm{log}^{3}(N)\rightarrow 0$ and $\mathrm{log}(NT)T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$.
Theorem 3.1 lays down the foundation for testing the null hypothesis $H_{0}^{\left(1\right)}.$ The proof of Theorem 3.1 is technically challenging and involved for two reasons. First, we allow for both temporal and cross-sectional dependence among $\left\{\left(X_{jt},\epsilon_{jt}\right)\right\}.$ Second, the latent variable $U_{jt}$ makes the error term $e_{jt}$ in (2.1) highly complex. Such complicated structure makes it impossible to apply any existing GA results or their proof strategies directly. In Section B.1.1 of the online supplement, we outline the main idea used in the proof of the above theorem, highlighting that the key step is proving the GA result in Lemma B.1. We note that the term $\mathrm{log}^{7/6}(NT)(\underline{b}T)^{-1/6}+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}$ arises from extending the original GA results in Chernozhukov et al., 2017a . The term $\mathrm{log}(NT)N^{1/q}T^{-\beta}$ is due to the temporal dependence and the moment condition of the innovation term $\eta_{jt}$. The term $\bar{b}^{1/3}\mathrm{log}^{2/3}(N)$ results from the covariance approximation and Gaussian comparison. Here we adopt maximum based test statistic, extending to other type of statistic can also be considered, see for example the statistics in Fan et al., (2015) and Giessing and Fan, (2020).
Next, we consider the GA result for $\mathcal{Q}$ defined in (2.15). Note that
$\mathcal{Q}$ $=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}\Big{(}(w_{jt,b}^{+}-w_{jt,b}^{-})Y_{jt}-\sum_{j=1}^{N}(w_{jt,b}^{+}-w_{jt,b}^{-})Y_{jt}/N\Big{)}\Big{|}/\tilde{v}_{j}$
$\approx\max_{1\leq j\leq N}\Big{|}(Tb_{j})^{1/2}\sum_{t=1}^{T}\Big{(}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}-\sum_{j=1}^{N}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}/N\Big{)}/\tilde{v}_{j}+\tilde{d}_{j}\Big{|},$ (3.6)
where $\tilde{d}_{j}=(Tb_{j})^{1/2}(\gamma_{j}-\frac{1}{N}\sum_{i=1}^{N}\gamma_{i})/\tilde{v}_{j}.$ The approximation is valid due to the smoothness of $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$, the remainder is addressed in Subsection C.3. Let $\tilde{\underline{d}}=(\tilde{d}_{1},\tilde{d}_{2},\ldots,\tilde{d}_{N})^{\top}.$ We deliver the GA results for $\mathcal{Q}$ in Supplementary Material A.2 and the proofs in Section B.2. Furthermore, the theoretical results regarding the derivative are given in Section A.5 and the proofs in Section B.7. In addition, in Section A.3, we establish the validity of the test statistics when variance estimators are utilized.
The most interesting finding is that in the GA result (3.5), neither temporal nor cross-sectional dependencies appear in $|Z+\underline{d}|_{\infty}$, despite allowing strong cross-sectional dependence for the underlying data generating process $X_{jt},U_{jt},\epsilon_{jt}$ within $\mathcal{I}$. The theoretical intuition behind this result is that local smoothing effectively whitens the noise. In practice, this observation suggests that, when the Assumption 3.3 hold for cross-sectional dependence, the correlation in the estimated statistics can be safely ignored, requiring only adjustments for heteroskedasticity.
3.3 Theoretical Results for the Unknown Threshold Case
In this section, we present theoretical results in the case of unknown thresholds. For simplicity of notation, we define $\mathcal{C}_{j}$ as the unified interval $\mathcal{C}_{j}=[c_{\min},c_{\max}]$, where $c_{\min}$ and $c_{\max}$ are constants in $\mathbb{R}$. The theorem still holds even when $\mathcal{C}_{j}$ differs for different $j$. Recall that $[c_{1},\ldots,c_{K}]$ is the grid we search over. Let $c_{1}=c_{\min}$ and $c_{K}=c_{\max}.$ As our statistics aggregate over different threshold grids, the problem becomes more demanding. Let $\Delta_{c,\min}=\min_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$ and $\Delta_{c,\max}=\max_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$.
Recall $w_{jt,b}(c_{i})$ in (2.16) and define ${\gamma}_{j}(c_{i})=\sum_{t=1}^{T}w_{jt,b}(c_{i})\gamma_{j}$. Let
$d_{jK}=(\gamma_{j}(c_{1})/v_{j}(c_{1}),\ldots,\gamma_{j}(c_{K})/v_{j}(c_{K}))^{\top}$
and $\underline{d}^{C}=(d_{1K}^{\top},d_{2K}^{\top},\ldots,d_{NK}^{\top})^{\top}.$ Define $Z^{C}=\left({{Z_{1}^{C\top},...,Z_{N}^{C\top}}}\right)^{\top}$ as an $NK$-vector of mean zero Gaussian distributed random vector with covariance matrix $\Sigma^{C}$ specified below. For each $j\in[N]$ and $1\leq i_{1},i_{2}\leq K$, the covariance between the $i_{1}$- and $i_{2}$-th elements of $Z_{j}^{C}$ is given by
$\mathrm{Cov}(Z_{j,i_{1}}^{C},Z_{j,i_{2}}^{C})=(v_{j}(c_{i_{1}})v_{j}(c_{i_{2}}))^{-1}(Tb_{j})\sum_{t=1}^{T}w_{jt,b}(c_{i_{1}})w_{jt,b}(c_{i_{2}})\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}.$
${\normalsize{{Z_{j_{1}}^{C}}}}$ and $Z_{j_{2}}^{C}$ are independent for $j_{1}\neq j_{2}$, so that the matrix $\Sigma^{C}$ is block diagonal. Note that $\underline{d}^{C}$ denotes a bias term which is a zero vector under the null hypothesis. In addition, when $2\max_{1\leq j\leq N}b_{j}<\Delta_{c,\min},$ one can ensure the matrix to be diagonal. The following theorem establishes the asymptotic property of our test statistics with unknown thresholds.
Theorem 3.2.
(Gaussian Approximation) Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Suppose that $\Delta_{c,\max}\rightarrow 0.$ Then
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{(NK)T},$ (3.7)
where $\mathcal{R}_{(NK)T}$ is $\mathcal{R}_{NT}$ in Theorem 3.1 with $N$ replaced by $NK$.
The proof of the above theorem is a generalization of Theorem 3.1 with a different variance-covariance structure. It should be noted that the conclusion of the above theorem follows the same pattern as Theorem 3.1, except that $N$ is replaced with $NK$. The following theorem provides the theoretical support for the consistency of the threshold estimator. Note that we only focus on the location $j$ where $\gamma_{j}$ does not equal to 00, and we require those breaks to be significant.
Theorem 3.3 (Consistency of the estimation of thresholds).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Assume $\min_{1\leq j\leq N:\gamma_{j}\neq 0}(Tb_{j})^{1/2}|\gamma_{j}|\gg 1$ and $\Delta_{c,\max}\rightarrow 0.$ Suppose $c_{0j}\in\{c_{i},1\leq i\leq K\}$, then
$\max_{1\leq j\leq N,\gamma_{j}\neq 0}\gamma_{j}^{2}|\hat{c}_{j}-c_{0j}|=O_{\mathbb{P}}((\mathrm{log}N)/T).$ (3.8)
If the true break $c_{0j}$ does not fall in the grid, then the difference is controlled by $\max\{\Delta_{c,\max},\mathrm{log}(N)/(T\gamma_{j}^{2})\}.$ Under the assumption of a minimum break signal, we can achieve consistency in detecting the break. The rate is expected to depend on both the signal strength $\gamma_{j}$ and the available sample size $T$. Note that the precision of threshold estimation is not affected by the bandwidth $b_{j}$. This rate is in line with the high-dimensional changepoint literature, see for example Theorem 2 in Li et al., (2024).
[…]
4 Monte Carlo Simulations
In this section, we analyze the finite sample performance of our GA results. In particular, we study the size and power of our uniform testing procedures in different settings. The data generating processes(DGPs) 1-7 are presented in Section E in Supplementary Materials.
4.1 Simulation Results
Throughout the study, we use a local linear estimator with a uniform kernel and select the bandwidth according to the procedure described in Calonico et al., (2014). We consider three significance levels, namely, $\alpha=0.1,$ $0.05,$ and $0.01$. The results are based on $1000$ Monte Carlo iterations. All tables and figures are presented in Section E in Supplementary Materials.
The results of simultaneous jump effects testing for DGP1–DGP3 are displayed in Table A1. First, the empirical size closely matches the nominal size across all $N$ values and for large $T$ in all DGPs. For DGP3, we observe slightly oversize when $N=10$ which may be caused by the presence of heteroskedasticity in comparison with results with DGP2. Overall, these simulation results confirm our theoretical findings in Theorem 3.1. In particular, it shows that the test procedure, which is based on critical values neglecting the dependence in the covariance structure, has fairly accurate size control even in the case of strong cross-sectional dependence in the innovations. Second, the local power properties of the test are reasonably well for all DGPs under investigation, and the power increases as either $N$ or $T$ increases. This result confirms that our testing procedure is well-suited to detect sparse alternatives.
In addition, we achieve comparable simulation performance in terms of size and power results for the test of homogeneity of the threshold effects in Table A2. The results confirm our Gaussian approximation result in Theorem A.1. For robustness, we also consider three additional settings with even stronger cross-sectional dependence for DGP4–6 in Table A3. The results show that size and power are indeed robust to strong cross-sectional dependence. Furthermore, Table A4 demonstrates that detecting the existence of jump effects simultaneously across cross-sectional dimensions in the case of unknown thresholds has size and power properties as good as in the known threshold case, confirming our theoretical results from Theorem 3.2. We additionally show the good finite sample performance for the estimation accuracy of the threshold location estimator in Table A5, providing Monte Carlo evidence supporting the consistency result of the threshold in Theorem 3.3.
For comparison, one direct method is the sup-Wald or sup-likelihood ratio test from the pooled threshold regression literature, neither of which accounts for nonlinearity or heterogeneity (see, for example, Andrews, (1993), Hansen, (2000) and Fong et al., (2017)). Notably, recent studies, such as Barassi et al., (2023), propose linear models with heterogeneous threshold effects in panel settings. However, they do not provide uniform testing methods. To facilitate comparison, we firstly restrict ourself to linear threshold model DGP7. And under the alternative we sample a fraction of coefficients $\gamma_{j}$ from a standard normal distribution. As shown in Table A6, our method has similar performance as the sup-Wald type statistic under the null hypothesis, while it has much better power under the alternative. See also Figure 1 in Section 2. The reasons for this are twofold. First, our uniform testing procedure is robust towards settings with sparse signal, in which a pooled test suffers from dilution of signals. And second, our uniform testing procedure does not have the problem of signal cancellation in case the signals for different $j$’s have opposite signs and cancel each other out. As a second comparison, we consider the nonparametric pooled test statistic under DGP3. Again, we draw the non-zero threshold coefficients from a standard normal. Such a setup without accounting for parameter heterogeneity is covered such as in Spokoiny, (1998), Yang and Song, (2014) and Chiou et al., (2018). As shown in Table A7, similar to the linear case, our method works with better power. The above superior performance is mainly due to the sparsity of the signal under the alternative hypothesis. In addition, our method has better power even for dense alternative comparing to the pooled type test statistic when there might exist signal cancellation as illustrated in Figure 1.
We also consider the test of existence of a threshold effect in the derivative under DGP1–3, to achieve similar size and power performance, oversmoothing is needed, as shown in Table A8.
5 Empirical Applications
In this section, we apply our testing procedure to two empirical applications. First, we study possible jumps in the news impact curve for constituents of the S&P 500. As a second application, we study the impact of the incumbency effect in U.S. House elections.
5.1 Application I: Jumps in the News Impact Curve
In this application, we revisit the news impact curve introduced by Engle and Ng, (1993). In particular, we want to identify possible jumps in the relationship between news events in the form of past returns on the volatility of stocks. The application serves as an illustration of our uniform testing procedure when the threshold $c_{0j}$ is unknown. Originally proposed in the context of autoregressive volatility models (e.g., ARCH), Engle and Ng, (1993) and Linton and Mammen, (2005) consider partially nonparametric specifications for the news impact curve which are able to capture possible asymmetries. A limitation of these specifications is the restriction to continuous functions. However, Zakoian, (1994) and Fornari and Mele, (1997) advocate for (parametric) GARCH specifications that allow for discontinuities in the way past returns impact volatility. We therefore want to investigate this issue by testing for the presence of jumps.
In our model setup, the dependent variable $Y_{jt}$ is the Garman-Klass volatility of stock $j$ at time $t$, and $X_{jt}$ is the lagged stock return. We want to test the null hypothesis $H_{0}^{(1)}:\gamma_{i}=0$ for all $i\in[N]$ and all $c_{0j}$, versus the alternative $H_{a}^{(1)}:\gamma_{i}\neq 0$ for some $i$ and some $c_{0j}$. Our testing procedure is robust to the presence of unobserved risk factors, $U_{jt}$, which are more than likely to occur in our application. Our data includes the daily stock returns and volatility data of the S&P 500 constituents from January 2010 to May 2024, which we obtained from Kaggle  [^2] https://www.kaggle.com/datasets/andrewmvd/sp-500-stocks.. Therefore we have a fairly high-dimensional setting with $N=500$ and $T_{j}$ is the number of days of stock $j$ being a member of the stock index. See Figure 2 for a visualization of the time series of returns and volatility for the stock of AT&T. As candidates for the threshold location, we consider $c_{i}\in\{-0.01,-0.005,0,0.005,0.01\}$.
Figure 2: Time series for daily stock returns (left panel) and daily volatility (right panel) for AT&T. The dashed lines in the left panel indicate the $10\%$ and $90\%$ quantiles of the return distribution.
Figure 3: Local linear fit of a significant threshold effect for AT&T (left panel) and an insignificant effect for Akamai (right panel).
The main findings are summarized in Table A9. We find significant results for $11$ out of the $500$ stocks we consider. For stocks with significant threshold effects, the locations of the estimated thresholds vary, and none of the significant threshold effects occur precisely at zero. This contradicts the rationale of the sign-switching ARCH model of Fornari and Mele, (1997), which models a jump effect based on the sign of the lagged returns, and thus expects a significant threshold effect at zero. For $c_{i}=-0.01$, we get significant negative effects at the $1\%$ level for five stocks, including Allstate Corporation and American Express. On the other hand, we obtain significant positive effects for the $c_{i}=0.01$ threshold for four stocks, e.g. AT&T and McKesson. These results reaffirm the findings of Chen and Ghysels, (2011), who find that both bad and good news have an increasing effect on volatility. Additionally, we get another negatively significant coefficient at the $-0.005$ threshold location and another positively significant coefficient at $c_{i}=0.005$. Overall, we get evidence for the presence of jump effects in the news impact curve. Figure 3 shows a visualization of the significant positive threshold effect for AT&T at $c_{i}=0.01$, and as a comparison we show the fit for Atamai for which we did not find a significant effect at any threshold location.
Furthermore, we are interested in studying the predictive advantage of our method compared to a baseline model without jumps. The baseline model can be written as $Y_{jt}=\tilde{h}_{j}(X_{jt})+e_{jt}$, where $\tilde{h}_{j}(\cdot)$ and $e_{jt}$ are defined in (2.1). For this purpose, we split the data into two parts, with training data from 2010–2019 and test data from 2020–2024. We restrict our analysis to the stocks identified in Table A9 with significant threshold effects at $1\%$ level. The mean squared error of prediction is lower for all but one of the 11 stocks we consider. The same is true if we restrict our analysis to these observations close to the respective threshold, $c_{0j}$. We conclude that including threshold effects can slightly improve the predictive accuracy with a median gain of $1.6\%$ in predictive power.
5.2 Application II: Party Incumbency Effects on U.S. House Elections
In the second application, we are interested in estimating the advantage a candidate has in the U.S. House elections when the seat is currently occupied by the party of the candidate  [^3] We use the data provided by the MIT Election Data and Science Lab, (2017). The data include the vote shares of all candidates in the elections for the US House of Representatives from 1976–2020. Due to redistricting at the start of each decade, we have to exclude years ending with a ‘0’ or ‘2’.. This is called the party incumbency effect. We follow the research design of Lee, (2008). See Caughey and Sekhon, (2011) for an analysis of more recent elections. The dependent variable $Y_{jt}$ is the democratic two-party vote share in year $j$ in district $t$. The covariate $X_{jt}$ is the difference in the two-party vote share in the previous election. Since the winner of the election is determined by the 0 threshold, our model setup is appropriate. We want to test the one-sided null hypothesis $H_{0}^{(1)}:\gamma_{i}<0$ for all $i\in[N]$, versus the alternative hypothesis $H_{a}^{(1)}$: $\gamma_{i}\geq 0$ for some $i$. Unlike Lee, (2008) who pools the data together to estimate a single incumbency effect, we conduct a separate analysis for different election years and also a separate analysis for different states by interchanging the roles of $j$ and $t$ and allowing for possible heterogeneity of incumbency effects along either the $j$ or $t$ dimension. In addition, it is easy to see our theories extend to the unbalanced data in which case we can use $T_{j}$ to denote the number of observations associated with individual/group $j$.
[…]
Now, we are interested in the state-specific effects. We restrict our attention to states with a sufficient number of elections held in the sample. We choose a threshold of 50 elections. This gives us $N=29$ states in our analysis, and $T_{j}$ now denotes the number of elections included for the $j$th state. We argue that this setting constitutes a fairly high-dimensional setting since the effective number of observations is smaller than $N$ for a fairly large proportion of the states. The results are reported in Table A12. Consistent with the previous results, almost all of the estimated coefficients are positive. By using the critical values for the simultaneous tests, we find significant effects for six states, for Missouri at $1\%$ level, and for California, Iowa, Virginia, Alabama and South Carolina at $5\%$. Again, we can reject the null hypothesis of no jump effect for our uniform testing procedure at a $1\%$ level. The significant states tend to belong to those with the largest number of observations. However, for the other two largest states, New York and Texas, the estimated effect is rather small and far from being significant. The incumbency effect seems to be stronger in some states compared to others. See Figure F.4 in the Supplement. for a visualization of the estimated threshold effects for the states of California and Pennsylvania. We also run the test for the homogeneity of incumbency effects ($H_{0}^{\left(2\right)}$) over the 29 states. The results are displayed in Table A13 and the test statistic is given by $\hat{\mathcal{Q}}=3.011$, which implies that we can reject the null hypothesis of homogeneous effects at the $5\%$ level. We find heterogeneity in the effects across election periods and states.
6 Conclusion
This paper focuses on the estimation and inference of heterogeneous threshold effects in nonparametric high-dimensional data, accounting for both cross-sectional and time dependencies in the covariates and error terms. We propose a test to assess the significance of the threshold effect and another to evaluate whether the threshold effects are homogeneous across individuals or groups. Additionally, we introduce a consistent method for estimating unknown threshold locations. Our tests are based on high-dimensional Gaussian approximation results. Despite the complexity of the underlying dependency structures, we show that the variance-covariance structure of the threshold effect estimators has a simple analytical expression under general dependence conditions. Unlike changepoint analysis in time series, our approach considers a setup involving latent variables, leading to a significantly different development of the GA results. Simulations demonstrate significant power enhancement compared to the existing literature under various settings.
SUPPLEMENTARY MATERIALS
Appendix A Additional Main Theorems
A.1 Additional Assumptions
Assumption SA-1.
(Boundedness) For each $j\in\left[N\right]$, $\varepsilon_{j}(\cdot,\cdot)$ is a bounded function.
Due to the construction of $\varepsilon_{j}(\cdot,\cdot)$ in (2.2), if functions $h_{j}(\cdot,\cdot)$ and $\tau_{j}(\cdot,\cdot)$ are both bounded, then the above assumption naturally holds. The next assumption imposes some smoothness conditions.
As in model (2.1), denote
$f_{j}(x)=\tilde{h}_{j}(x)+\tilde{\tau}_{j}(x)\mathbf{1}_{\{x\geq c_{0j}\}}-\tilde{\tau}_{j}(c_{0j}).$ (SA-1)
Assumption SA-2.
(Smoothness) (i) The density functions $g_{j}(\cdot)$ of $X_{jt}$ are lower bounded on $\mathcal{C}_{j},$ that is for any $x\in\mathcal{C}_{j},$ assume $\min_{1\leq j\leq N}g_{j}(x)\geq c$ for some constant $c>0$. The conditional densities $g_{j,t}(\cdot|\mathcal{F}_{s})$ satisfy that $\max_{j,t}|g_{j,t}(\cdot|\mathcal{F}_{t-1})|$ $=O(1)$ and that
$\max_{1\leq j\leq N,1\leq t\leq T}\sup_{x\in\mathcal{C}_{j}}\sup_{|y-x|\leq b_{j}}|g_{j,t}(y|\mathcal{F}_{t-1})-g_{j,t}(x|\mathcal{F}_{t-1})|=O(\bar{b}).$
(ii) For all $j\in[N],$ assume that for any $x\in\mathcal{C}_{j}$, function $f_{j}(\cdot)$ satisfies
$f_{j}(y)=\begin{cases}f_{j}(x)+\partial_{+}f_{j}(x)(y-x)+O((y-x)^{2})\quad&\text{if}\ y\in[x,x+b_{j}],\\ f_{j}(x)+\partial_{-}f_{j}(x)(y-x)+O((y-x)^{2})&\text{if}\ y\in[x-b_{j},x],\end{cases}$
where $\partial_{+}f_{j}(x)$ (resp. $\partial_{-}f_{j}(x)$) is the right (resp. left) derivative of function $f_{j}\left(\cdot\right)$ at point $x$, and the constants in $O(\cdot)$ are independent of $j,b_{j}$.
(iii) The standard deviation functions $\sigma_{j}(\cdot,\cdot)$ satisfy $\inf_{u\in\mathbb{R},x\in\mathcal{C}_{j}}\sigma_{j}(x,u)>0$ with
$\max_{1\leq j\leq N}\sup_{u\in\mathbb{R},x\in\mathcal{C}_{j},|y-x|\leq b_{j}}|\sigma_{j}(y,u)/\sigma_{j}(x,u)-1|=1+O(\bar{b}).$
Remark 2.
Assumption SA-2 $(i)$ imposes a lower bound condition for the density functions $g_{j}(\cdot)$. Condition $(ii)$ imposes some smoothness conditions of the mean function $f_{j}(\cdot).$ Note that we do not require $f_{j}(\cdot)$ to be differentiable at point $x$; it suffices for $f_{j}(\cdot)$ to have bounded left and right second derivatives near $x$. Assumption $(iii)$ is for the smoothness of the variance function $\sigma_{j}^{2}(\cdot,\cdot)$. If $\sigma_{j}(x,u)$ has bounded partial left and right derivatives with respect to $x$, for any $x\in\mathcal{C}_{j}$ uniformly over $u,$ then assumption $(iii)$ is fulfilled. Note that we do not assume the left and right derivatives of $f_{j}(\cdot)$ to be the same, meaning the slope around the threshold can differ. To be general, we do not impose any differentiability assumptions on $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$.            $\Box$
The following assumption imposes some regularity conditions on the kernel function and bandwidth parameter.
Assumption SA-3.
[…]
$1<\frac{K_{2}^{+}K_{0}^{+}}{K_{1}^{+2}},\frac{K_{2}^{-}K_{0}^{-}}{K_{1}^{-2}}\leq c$ (SA-2)
for some constant $c>1.$
(ii) As $\left(N,T\right)\rightarrow\infty,$ assume $\bar{b}\rightarrow 0$ and $\underline{b}T/\mathrm{log}(NT)\rightarrow\infty.$ Additionally, assume that $b\asymp\underline{b}\asymp\bar{b}$.
Remark 3.
Assumption SA-3$(i)$-$(ii)$ imposes certain conditions on the kernel function $K\left(\cdot\right)$and bandwidth parameters $b_{j}$’s, respectively. If one chooses $b_{j}$ to be the same across $j,$ Assumption SA-3$(ii)$ reduces to the standard conditions on the bandwidth parameter in local linear kernel estimation with one-dimensional covariate.            $\Box$
A.2 A Homogeneity Test
Theorem A.1 (Gaussian Approximation 2).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\Big{(}\mathcal{Q}\leq u\Big{)}-\mathbb{P}(|Z+\tilde{\underline{d}}|_{\infty}\leq u)\Big{|}\lesssim\mathcal{R}_{NT}.$ (SA-3)
Theorem A.1 establishes the foundation for testing the null hypothesis $H_{0}^{\left(2\right)}.$ Its proof cannot directly follow from Theorem 3.1 due to the introduction of the additional averaging term of $Y_{jt}$ which might not be of smaller order. However, the steps can be simplified by imposing weak cross-sectional dependence, as the additional averaging term would then be of higher order under this assumption.
Remark 4.
(Allowance of $N$ and $T$ and bandwidth selection) The relative rate of $N$ and $T$ is determined by the moment assumptions on the error $\eta_{11}$ (Assumption 3.1) and the temporal dependence (Assumption 3.2). When we assume exponential moment conditions (i.e., Assumption 3.1 $(ii)$), $N$ can be of exponential order of $T$. Specifically, if we assume that $b\propto T^{-c}$ with $c>1/5$, then $\mathcal{R}_{NT}\rightarrow 0$ when $\mathrm{log}N\propto T^{r}$, where $30<r<(1-c)/7\wedge(5c-1)\wedge\beta\wedge c/$ 3. When we assume finite moment conditions (i.e., Assumption 3.1 $(i)$), we require $N$ to be of polynomial order of $T$. If we assume that $b\propto T^{-c}$ with $c\in(1/5,1-q/2),$ then $\mathcal{R}_{NT}\rightarrow 0$ when $N\propto T^{\tilde{r}}$ for $0<\tilde{r}<\beta q$. To address the bandwidth selection, we can calculate $\hat{b}_{j}$ for each individual following the procedure of Calonico et al., (2014). Sometimes, it is also recommended to aggregate the bandwidth to avoid the computational burden of estimating the bandwidths for each individual/group.            $\Box$
A.3 Asymptotic Validity of the Feasible Testing Procedures
[…]
Theorem A.2 (Consistency of the variance estimator).
Suppose that Assumptions 3.1- 3.3 and SA-1-SA-3 hold and $N(\underline{b}T)^{-q/4}\rightarrow 0.$ Then we have $\max_{j\in[N]}|\hat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N))$ and $\max_{j\in[N]}|\hat{v}_{j}^{2}-v_{j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N)).$
[…]
Recall that $q_{\alpha}$ is the $1-\alpha$ quantile of the maximum Gaussian random variable $\max_{j\in\left[N\right]}\left|Z_{j}\right|$ in Algorithm 1. The following theorem studies the asymptotic power properties of the two tests.
Theorem A.3 (Power).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Let $\bar{\gamma}=\sum_{j=1}^{N}\gamma_{j}/N.$
[…]
$(ii)$ If $\mathcal{R}_{NT}\to 0$ in Theorem A.1 and $\max_{j\in\left[N\right]}(Tb_{j})^{1/2}|\gamma_{j}-\bar{\gamma}|\gg(\mathrm{log}N)^{1/2},$ then $\lim_{\left(N,T\right)\rightarrow\infty}\mathbb{P}(\mathcal{\hat{Q}}>q_{\alpha})=1.$
[…]
A.4 Theorems for the Unknown Threshold Case
Let $\hat{e}_{jt}$ be the residual by regressing $Y_{jt}$ on $X_{jt}$ after a local linear regression with a bandwidth $b^{*}\ll\underline{b}$. Let $\phi_{\acute{a}}(x)=\min(\acute{a},|x|)$ with $\acute{a}=c\mathrm{log}^{1/2}(NK).$ Define
$\hat{\sigma}_{e,j}^{2}(c_{i})=\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}\phi_{\acute{a}}(\hat{e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}.$ (SA-4)
Theorem A.4 (Consistency of the variance estimator).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold and $N(bT)^{-q/4}\rightarrow 0.$ Then we have
$\max_{j\in[N]}\max_{1\leq i\leq K}|\hat{\sigma}_{e,j}^{2}(c_{i})-\sigma_{e,j}^{2}(c_{i})|$ $=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b}+\acute{a}^{-p/2+1}+b^{*2})$ and $\max_{j\in[N]}|\hat{v}_{j}^{2}-v_{j}^{2}|$ $=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b}+\acute{a}^{-p/2+1}+b^{*2})$.
Theorem A.5 (Consistency of the tests).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Let $\bar{\gamma}=\sum_{j=1}^{N}\gamma_{j}/N.$
[…]
A.5 Theorems on Derivatives
In this subsection, we provides regarding deritative estimators. Consider the test statistic
$\mathcal{I}^{[1]}=\max_{1\leq j\leq N}(Tb_{j}^{3})^{1/2}|\hat{\gamma}_{j}^{[1]}/v_{j}^{[1]}|,$ (SA-5)
where
$v_{j}^{[1]2}=v_{j}^{[1]2}(c_{0j})=(Tb_{j}^{3})\sum_{t=1}^{T}(w_{jt,b}^{+[1]}-w_{jt,b}^{-[1]})^{2}\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}.$ (SA-6)
Define $d_{j}^{[1]}=(Tb_{j}^{3})^{1/2}\gamma^{[1]}_{j}/v_{j}^{[1]}$ and $\underline{d}^{[1]}=(d_{1}^{[1]},d_{2}^{[1]},\ldots,d_{N}^{[1]})^{\top}$.
Theorem A.6.
Let Assumptions 3.1-3.3 and SA-1-SA-3 hold. Then we have
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}^{[1]}\leq u)-\mathbb{P}(|Z+\underline{d}^{[1]}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{NT}.$ (SA-7)
Denote $\mathcal{I}^{C[1]}$, $Z^{C[1]}$ and $\underline{d}^{C[1]}$ as the derivative counterparts of $\mathcal{I}^{C[1]}$, $Z^{C}$ and $\underline{d}^{C}$. Let
$d_{jK}^{[1]}=(\gamma_{j}^{[1]}(c_{1})/v_{j}^{[1]}(c_{1}),\ldots,\gamma_{j}^{[1]}(c_{K})/v_{j}^{[1]}(c_{K}))^{\top},$
where $\underline{d}^{C[1]}=(d_{1K}^{[1]\top},d_{2K}^{[1]\top},\ldots,d_{NK}^{[1]\top})^{\top}.$ Let $\Sigma^{C[1]}$ be the variance covariance matrix of $Z^{C[1]}$. For each $j\in[N]$ and $1\leq i_{1},i_{2}\leq K$, the covariance between the $i_{1}$- and $i_{2}$-th elements of $Z_{j}^{C[1]}$ is given by
$\mathrm{Cov}(Z_{j,i_{1}}^{C[1]},Z_{j,i_{2}}^{C[1]})=(v_{j}(c_{i_{1}}^{[1]})v_{j}^{[1]}(c_{i_{2}}))^{-1}(Tb_{j}^{3})\sum_{t=1}^{T}w^{[1]}_{jt,b}(c_{i_{1}})w^{[1]}_{jt,b}(c_{i_{2}})\mathrm{Var}\big{(}e_{jt}|X_{jt}\big{)}.$
${\normalsize{{Z_{j_{1}}^{C[1]}}}}$ and $Z_{j_{2}}^{C[1]}$ are independent for $j_{1}\neq j_{2}$, so that the matrix $\Sigma^{C[1]}$ is block diagonal.
$\mathcal{I}^{C[1]}=\max_{1\leq i\leq K}\max_{1\leq j\leq N}(Tb_{j}^{3})^{1/2}|\hat{\gamma}_{j}^{[1]}(c_{i})/v_{j}^{[1]}(c_{i})|.$ (SA-8)
Theorem A.7.
(Gaussian Approximation) Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Suppose that $\Delta_{c,\max}\rightarrow 0.$ Then
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}^{C[1]}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C[1]}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{(NK)T}.$ (SA-9)
Define $\hat{v}_{j}^{[1]2}=Tb_{j}^{3}\sum_{t=1}^{T}\left(w_{jt,b}^{+[1]}-w_{jt,b}^{-[1]}\right)^{2}\hat{\sigma}_{e,j}^{2}$ and $\hat{\tilde{v}}_{j}^{[1]}=\left(\hat{v}_{j}^{[1]2}\left(1-\frac{1}{N}\right)^{2}+\frac{1}{N^{2}}\sum_{i\neq j}^{N}\hat{v}_{i}^{[1]2}\right)^{1/2}$.
Theorem A.8 (Consistency of the variance estimator).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold and $N(\underline{b}T)^{-q/4}\rightarrow 0.$ Then we have $\max_{j\in[N]}|\hat{v}_{j}^{[1]2}-v_{j}^{[1]2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N)).$
[…]
Theorem A.9 (Power).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Let $\bar{\gamma}^{[1]}=\sum_{j=1}^{N}\gamma_{j}^{[1]}/N.$
[…]
$(ii)$ If $\mathcal{R}_{NT}\to 0$ in Theorem A.1 and $\max_{j\in\left[N\right]}(Tb_{j}^{3})^{1/2}|\gamma_{j}^{[1]}-\bar{\gamma}^{[1]}|\gg(\mathrm{log}N)^{1/2},$ then $\lim_{\left(N,T\right)\rightarrow\infty}\mathbb{P}(\mathcal{\hat{Q}}^{[1]}>q_{\alpha})=1.$
Regarding the unknown threshold case for the derivatives, we expect the algorithms and theorems can be extended in a similar fashion; therefore, we omit the details here. We are aware that the rates of the derivative estimator can be improved by local quartic estimation, see for example Fan, (2018).
Appendix B Proofs of the Main Results in the Paper
B.1 Proof of Theorem 3.1
In this section, we first outline the main idea for the proof of Theorem 3.1 and present the main lemma for Gaussian approximation. Then we prove Theorems 3.1A.3 in turn.
B.1.1 Outline of the Proof of Theorem 3.1
For exposition purpose, we outline the main step in the proof of Theorem 3.1. The intuition of proving Theorem A.1 is similar. Define the statistic $I_{\epsilon}$ after canceling the smooth component in the main text,
$I_{\epsilon}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})[\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}]+\gamma_{j}\Big{|}/v_{j}.$ (SA-1)
By the continuity of $f_{j}\left(\cdot\right)$ in Assumption SA-2 and Lemma B.2 $(i)$ below, we have
$|\mathcal{I}-I_{\epsilon}|\leq\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})f_{j}(X_{jt})\Big{|}/v_{j}=O(T^{1/2}\bar{b}^{5/2}).$ (SA-2)
Thus we only need to work with $I_{\epsilon}.$ We prove by an $m$-lag truncation to facilitate us to construct an $m$-dependent sequence. Specifically, for some $m>0,$ consider a truncated version of the error term:
$\epsilon_{t,m}=(\epsilon_{1t,m},\epsilon_{2t,m},\ldots,\epsilon_{Nt,m})^{\top}=\sum_{k=0}^{m-1}A_{k}\eta_{t-k}.$ (SA-3)
Let $I_{\epsilon,m}$ be the $m$-dependent approximation of $I_{\epsilon},$ where $I_{\epsilon,m}$ is $I_{\epsilon}$ with $\epsilon_{t}$ therein replaced by $\epsilon_{t,m},$ namely,
$I_{\epsilon,m}=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}\big{[}\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}\big{]}+\gamma_{j}\Big{|}/v_{j},$ (SA-4)
where $w_{jt,b}:=(w_{jt,b}^{+}-w_{jt,b}^{-})\mathbf{1}_{\mathcal{A}_{T}}$ where $\mathcal{A}_{T}$ is defined in Lemma B.2 $(ii)$ below. For large $m,$ due to the decay of temporal dependence assumption, we expect
$I_{\epsilon}\approx I_{\epsilon,m}.$
Let $Z_{j,m}$’s be normally distributed random variables that are independent across $j$. Denote the $m$-dependent Gaussian counterpart as
$\tilde{I}_{z,m}=\max_{1\leq j\leq N}\Big{|}Z_{j,m}+\frac{(Tb_{j})^{1/2}}{v_{j}}\gamma_{j}\Big{|}.$
Similarly, let $I_{z}$ (resp. $I_{z,m}$) be $I_{\epsilon}$ (resp. $I_{\epsilon,m}$) with $\eta_{t}$ therein replaced by $z_{t}$, where $\left\{z_{t}\right\}_{t\in\mathbb{Z}}$ are i.i.d. Gaussian vectors with zero mean and identity covariance matrix. Since $I_{\epsilon,m}$ can be rewritten as the maximum of a sum of independent vectors conditional on $X_{jt},U_{jt}$, we can apply the GA theorem from Chernozhukov et al., 2017b to show the GA for the conditional part. For the remaining part, we need to demonstrate GA again, as detailed in Lemma B.1 below, then we have
$I_{\epsilon,m}\overset{\mathbb{P}}{\approx}I_{z,m}.$
We complete the proof by showing that the distributions of $I_{z,m}$ and $|Z+\underline{d}|_{\infty}$ are close, and the continuity of the maximum of a non-centered Gaussian vector. Then the conclusion follows.
B.1.2 A Lemma on Gaussian Approximation
This subsection presents a Gaussian approximation lemma that utilizes a nonstandard technique compared to the existing literature. The lemma specifically focuses on the $m$-dependent part $I_{\epsilon,m}$ in (SA-4). The difference between $I_{\epsilon,m}$ and $I_{\epsilon}$ is effectively controlled in Lemmas B.7 and B.8 below. Recall that $\mathcal{F}_{t}=(e_{t},e_{t-1},\ldots)$ and ~$\tilde{\mathcal{F}}_{t}=(\tilde{e}_{t},\tilde{e}_{t-1},\ldots).$
The non-triviality of the proof of the theorems lies in the fact that the interplay between the two terms in (SA-4):
$\frac{(Tb_{j})^{1/2}}{v_{j}}\sum_{t=1}^{T}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}\quad\textrm{and}\quad\frac{(Tb_{j})^{1/2}}{v_{j}}\sum_{t=1}^{T}w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})$
The second term arises from the potential presence of the latent variable $U_{jt}$. If $U_{jt}$ is absent, the second term will disappear, allowing time-dependent Gaussian approximation to be applied to the first term. This second term causes the derivation steps to differ from those in the existing literature. Noting that the second term becomes determined when we condition on $\{X_{jt},U_{jt}\}_{1\leq t\leq T,1\leq j\leq N}$, we will first prove a GA for the first term under this conditioning. Then we shall replace the first term by its Gaussian counterparts $Z_{j,1}=(Tb_{j})^{1/2}v_{j}^{-1}\sum_{t=1}^{T}\tilde{\xi}_{jt,1}$, where $\tilde{\xi}_{jt,1}$ are i.i.d. normal and independent from $\mathcal{F}_{t}$ and ~$\tilde{\mathcal{F}}_{t}$. Then we can show the GA for the summation of this counterpart and the second term.
The following lemma is the key lemma for Gaussian approximation, whose proof can be found in Section C.1.
Lemma B.1.
Let $T/2\geq m\rightarrow\infty.$
$(i)$ Suppose that the conditions in Theorem 3.1 $(i)$ hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(I_{\epsilon,m}\leq u)-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j})^{1/2}\gamma_{j}/v_{j}\big{|}\leq u\Big{)}\Big{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)},$
where the constant in $\lesssim$ is independent of $N,T.$
$(ii)$ Suppose that the conditions in Theorem 3.1 $(ii)$ hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(I_{\epsilon,m}\leq u)-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j})^{1/2}\gamma_{j}/v_{j}\big{|}\leq u\Big{)}\Big{|}$
$\leq$ $(\underline{b}T)^{-1/6}\mathrm{log}(NT)^{7/6}+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)},$
where the constant in $\lesssim$ is independent of $N,T.$
B.1.3 Main Steps of Proof of Theorem 3.1
To prove Theorem 3.1, we need the following seven lemmas. See Section C.2 for the proofs of Lemmas B.2, B.7, and B.8. It should be noted that, for simplicity in the exposition of $W_{jt,b}$, we define an event $\mathcal{A}_{T}$ in Lemma B.2, which occurs with probability $1-b_{T}$, where $b_{T}\to 0$.
Lemma B.2.
(Properties for weights) Under Assumptions 3.1-3.3 and SA-1-SA-3, the followings hold:
[…]
(ii) For $l=0,1,2,$ on some set $\mathcal{A}_{T}$ with
$\mathbb{P}(\mathcal{A}_{T})\geq 1-O(T^{-(\alpha p)\wedge(p/2-1)})$ (SA-5)
we have
$\max_{1\leq j\leq N}\big{|}S_{jl,b}^{+}/(Tb_{j}^{l+1})-K_{l}^{+}g_{j}(c_{j0})\big{|}\lesssim\sqrt{\mathrm{log}(TN)/(T\underline{b})}+\bar{b},$ (SA-6)
where the constants in $O(\cdot)$ and $\lesssim$ are independent of $N,T,\underline{b},$ and $\bar{b}.$ And thus on set $\mathcal{A}_{T}$,
$\max_{1\leq j\leq N}\Big{|}w_{jt,b}^{+}-\frac{K((X_{jt}-c_{j0})/b_{j})}{g_{j}(c_{j0})Tb_{j}}\frac{K_{2}^{+}-K_{1}^{+}(X_{jt}-c_{j0})/b_{j}}{K_{2}^{+}K_{0}^{+}-K_{1}^{+2}}\mathbf{1}_{\{X_{jt}\geq c_{j0}\}}\Big{|}$
$\lesssim$ $\sqrt{\mathrm{log}(TN)/(\underline{b}T)^{3}}+1/T,$ (SA-7)
and moreover,
$\sum_{t=1}^{T}|w_{jt,b}^{+}|=O(1)\quad\text{and}\quad\sum_{t=1}^{T}|w_{jt,b}^{-}|=O(1)$ (SA-8)
uniformly over $j$.
The following is a high dimensional extension of the concentration inequality in Freedman, (1975).
Lemma B.3.
(Freedman’s inequality) Let $\mathcal{A}$ be an index set with $|\mathcal{A}|<\infty$. For each $a\in\mathcal{A}$, let $\{\xi_{a,i}\}_{i=1}^{n}$ be a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{i}\}_{i=1}^{n}.$ Let $M_{a}=\sum_{i=1}^{n}\xi_{a,i}$ and $V_{a}=\sum_{i=1}^{n}\mathbb{E}[\xi_{a,i}^{2}|\mathcal{F}_{i-1}]$. Then for all $z,u,v>0$
$\mathbb{P}\Big{(}\max_{a\in\mathcal{A}}|M_{a}|\geq z\Big{)}\leq\sum_{i=1}^{n}\mathbb{P}\Big{(}\max_{a\in\mathcal{A}}|\xi_{a,i}|\geq u\Big{)}+2\mathbb{P}\Big{(}\max_{a\in\mathcal{A}}V_{a}\geq v\Big{)}+2|\mathcal{A}|e^{-z^{2}/(2zu+2v)}.$
Lemma B.4 (Burkholder, (1988), Rio, (2009)).
Let $q>1$, $q^{\prime}=\min\{q,2\}$. Let $M_{T}=\sum_{t=1}^{T}\xi_{t},$ where $\xi_{t}\in\mathcal{L}^{q}$ are martingale differences. Then
$\|M_{T}\|_{q}^{q^{\prime}}\leq K_{q}^{q^{\prime}}\sum_{t=1}^{T}\|\xi_{t}\|_{q}^{q^{\prime}},\mbox{ where }K_{q}=\max\big{(}(q-1)^{-1},\sqrt{q-1}\big{)}.$
Lemma B.5 (Nazarov, (2003)).
Let $X=(X_{1},X_{2},\ldots,X_{v})^{\top}$ be a centered Gaussian vector in $\mathbb{R}^{v}.$ Assume $\mathbb{E}(X_{i}^{2})\geq b$ for some $b>0$ and all $1\leq i\leq v.$ Then for any $e>0$ and $d\in\mathbb{R}^{v},$
$\sup_{x\in\mathbb{R}}\mathbb{P}\Big{(}\big{|}|X+d|_{\infty}-x\big{|}\leq e\Big{)}\leq ce\sqrt{\mathrm{log}(v)},$ (SA-9)
where $c$ is some constant depending only on $b$.
The following lemma is Lemma 3 in Chen et al., (2022).
Lemma B.6 (Comparison).
Let $X=(X_{1},X_{2},\ldots,X_{v})^{\top}$ and $Y=(Y_{1},Y_{2},\ldots,Y_{v})^{\top}$ be two centered Gaussian vectors in $\mathbb{R}^{v}$ and let $d=(d_{1},d_{2},\ldots,d_{v})^{\top}\in\mathbb{R}^{v}.$ We denote $\Delta=\max_{1\leq i,j\leq v}|\sigma_{i,j}^{X}-\sigma_{i,j}^{Y}|,$ where we define $\sigma_{i,j}^{X}=\mathbb{E}(X_{i}X_{j})$ (resp. $\sigma_{i,j}^{Y}=\mathbb{E}(Y_{i}Y_{j})$). Assume that $Y_{i}$s have the same variance $\sigma^{2}>0.$ Then we have
$\sup_{x\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}|X+d|_{\infty}\leq x\big{)}-\mathbb{P}\big{(}|Y+d|_{\infty}\leq x\big{)}\Big{|}\lesssim\Delta^{1/3}\mathrm{log}(v)^{2/3},$ (SA-10)
where the constant involved in $\lesssim$ only depends on $\sigma.$
Lemma B.7 ($m$-dependent approximation for polynomial case).
Assume that the conditions in Theorem 3.1 $(i)$ hold. For some $1\ll m\leq T/2,$ we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{log}^{q}(NT)(T\underline{b})^{1-q/2}m^{-q\beta}z^{-q}+Ne^{-z^{2}m^{2\beta}}+\mathbb{P}(\mathcal{A}_{T}^{c}).$
where $\mathcal{A}_{T}$ is defined in Lemma B.2.
Lemma B.8 ($m$-dependent approximation for sub-exponential case).
Assume that the conditions in Theorem 3.1 $(ii)$ hold. For some $1\ll m\leq T/2,$ we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{exp}\Big{\{}-c\frac{z^{2}}{m^{-\beta}(\underline{b}T)^{-1/2}z+m^{-2\beta}}\Big{\}}+\mathbb{P}(\mathcal{A}_{T}^{c}),$
where $c>0$ is some constant and $\mathcal{A}_{T}$ is defined in Lemma B.2.
Proof of Theorem 3.1.
$(i)$ Recall the definition of $I_{\epsilon}$ in (SA-1) and $I_{\epsilon,m}$ is the $m$-dependent part of $I_{\epsilon}$. Let $m=T/2.$ For any $l>0,$ we have
$\mathbb{P}\big{(}\mathcal{I}\leq u\big{)}\leq\mathbb{P}\big{(}|\mathcal{I}-I_{\epsilon,m}|\geq l\big{)}+\mathbb{P}\big{(}I_{\epsilon,m}\leq u+l\big{)},$
and similarly
$\mathbb{P}\big{(}I_{z}\leq u+2l\big{)}\geq-\mathbb{P}\big{(}|I_{z}-I_{z,m}|\geq l\big{)}+\mathbb{P}\big{(}I_{z,m}\leq u+l\big{)}.$
Combining the above inequalities leads to
$\sup_{u\in\mathbb{R}}\Big{[}\mathbb{P}\big{(}\mathcal{I}\leq u)-\mathbb{P}(I_{z}\leq u\big{)}\Big{]}$
$\leq$ $\mathbb{P}\big{(}|\mathcal{I}-I_{\epsilon,m}|\geq l\big{)}+\mathbb{P}\big{(}|I_{z}-I_{z,m}|\geq l\big{)}+\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(I_{\epsilon,m}\leq u)-\mathbb{P}(I_{z,m}\leq u)\big{|}$
$+\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}\big{(}I_{z}\leq u+2l\big{)}-\mathbb{P}\big{(}I_{z}\leq u\big{)}\big{|}$
$=$ $:\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}+\mathrm{I}_{4}.$ (SA-11)
For part $\mathrm{I}_{1}$, we have $|\mathcal{I}-I_{\epsilon,m}|\leq|\mathcal{I}-I_{\epsilon}|+|I_{\epsilon}-I_{\epsilon,m}|.$ By Assumption SA-2 and Lemma B.2,
$\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}w_{jt,b}^{+}f_{j}(X_{jt})-f_{j}(c_{j0})\Big{|}$
$=$ $\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}w_{jt,b}^{+}\big{[}f_{j}(c_{j0})+\partial_{+}f_{j}(c_{j0})(X_{jt}-c_{j0})+O((X_{jt}-c_{j0})^{2})\big{]}-f_{j}(c_{j0})\Big{|}$
$=$ $O(\bar{b}^{2}),$
where the constant in $O(\cdot)$ here and all the followings are independent of $N,T$. The same bound can be derived with $w_{jt,b}^{+}$ therein replaced by $w_{jt,b}^{-}$, and thus
$|\mathcal{I}-I_{\epsilon}|\leq\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})f_{j}(X_{jt})\Big{|}/v_{j}=O(T^{1/2}\bar{b}^{5/2}).$ (SA-12)
We define $l^{\prime}=N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NT)}$. Then by Lemma B.7 with $m=T/2$, we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\geq l^{\prime}\big{)}=O\big{\{}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{\}}.$ (SA-13)
Hence for $l=l^{\prime}+cT^{1/2}\bar{b}^{5/2}$, where $c>0$ is some constant large enough, we have
$\mathrm{I}_{1}=O\big{\{}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{\}}.$
Similar argument leads to the same bound for part $\mathrm{I}_{2}$.
For part $\mathrm{I}_{3}$, by Lemma B.1 $(i)$ we have
$\mathrm{I}_{3}\lesssim(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+T^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{4}$, we use Lemma B.5 to obtain
$\mathrm{I}_{4}\leq\sup_{u\in\mathbb{R}}\mathbb{P}\big{(}|I_{z}-u\big{|}\leq 2l\big{)}\lesssim l\mathrm{log}(N)^{1/2}.$
We complete the proof by combining the $\mathrm{I}_{1}$-$\mathrm{I}_{4}$ parts and a similar argument for the other side of the inequality.
$(ii)$ As in part $(i)$, we have the decomposition as in (SA-11). By (SA-12) and Lemma B.8 with $m=T/2$, we have $\mathrm{I}_{1}=O((NT)^{-q})$ with $l=O(T^{-\beta}\mathrm{log}(NT)^{1/2}+T^{1/2}b^{5/2}).$ The same bound can be derived for part $\mathrm{I}_{2}$. By Lemma B.1 $(ii)$, we have
$\mathrm{I}_{3}\lesssim(bT)^{-1/6}\mathrm{log}^{7/6}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{4}$, we use the same argument as in part $(i)$ to obtain
$\mathrm{I}_{4}\lesssim l\mathrm{log}(N)^{1/2}\lesssim T^{-\beta}\mathrm{log}(NT)+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}.$
The desired result follows by combining the $\mathrm{I}_{1}$-$\mathrm{I}_{4}$ and a similar argument for the other side of the inequality.
B.2 Proof of Theorem A.1
Denote
$a_{jt,b}=w_{jt,b}\sigma_{j}(X_{jt},U_{jt})$ (SA-14)
and let
$\mathcal{Q}_{\epsilon}=\max_{1\leq j\leq N}\Bigg{|}\sum_{t=1}^{T}$ $\Big{(}a_{jt,b}\epsilon_{jt}-\sum_{l=1}^{N}a_{lt,b}\epsilon_{lt}/N+w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})$
$-\sum_{l=1}^{N}w_{lt,b}\varepsilon_{l}(X_{lt},U_{lt})/N\Big{)}+(\gamma_{j}-\bar{\gamma})\Bigg{|}/\tilde{v}_{j}.$ (SA-15)
Denote $\mathcal{Q}_{z}$ be $\mathcal{Q}_{\epsilon}$ with $\epsilon_{jt}$ and $U_{jt}$ replaced by Gaussian distributed random variable with the same covariance structure and $\mathcal{Q}_{\epsilon,m}$ be $\mathcal{Q}_{\epsilon}$ with $\epsilon_{jt}$ therein replaced by $\epsilon_{jt,m}.$ To prove Theorem A.1, we add two lemmas.
Lemma B.9 ($m$-dependent approximation).
Let $20\ll m\leq T/$ 2 .
(i) Under the conditions in Theorem A.1(i) we have
$\mathbb{P}\big{(}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{log}(NT)^{q}(T\underline{b})^{1-q/2}m^{-q\beta}z^{-q}+Ne^{-z^{2}m^{2\beta}}+\mathbb{P}(\mathcal{A}_{T}^{c});$
(ii) Under the conditions in Theorem A.1(ii), we have
$\mathbb{P}\big{(}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq z\big{)}\lesssim N\mathrm{exp}\Big{\{}-c\frac{z^{2}}{m^{-\beta}(\underline{b}T)^{-1/2}z+m^{-2\beta}}\Big{\}}+\mathbb{P}(\mathcal{A}_{T}^{c}).$
Lemma B.10.
Let $20<m\leq T/$ 2 .
(i) Assume the conditions in Theorem A.1 $(i)$ hold. Then we have
$\sup_{u\in\mathbb{R}}|\mathbb{P}(\mathcal{Q}_{\epsilon,m}\leq u)-\mathbb{P}(\mathcal{Q}_{z,m}\leq u)|$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
(ii) Assume the conditions in Theorem A.1 $(ii)$ hold. Then we have
$\sup_{u\in\mathbb{R}}|\mathbb{P}(\mathcal{Q}_{\epsilon,m}\leq u)-\mathbb{P}(\mathcal{Q}_{z,m}\leq u)|\lesssim(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
Here the constants in $\lesssim$ are independent of $\left(N,T,b\right).$
Proof of Theorem A.1
$(i)$ Let $m=T/2.$ For any $l>0,$ as in (SA-11), we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}(T\underline{b})^{1/2}\mathcal{Q}\leq u)-\mathbb{P}((T\underline{b})^{1/2}\mathcal{Q}_{z}\leq u\big{)}\Big{|}$
$\leq$ $\mathbb{P}\Big{(}(T\underline{b})^{1/2}|\mathcal{Q}-\mathcal{Q}_{\epsilon,m}|\geq l\Big{)}+\mathbb{P}\Big{(}(T\underline{b})^{1/2}|\mathcal{Q}_{z}-\mathcal{Q}_{z,m}|\geq l\Big{)}+\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(\mathcal{Q}_{\epsilon,m}\leq u)-\mathbb{P}(\mathcal{Q}_{z,m}\leq u)\Big{|}$
$+\sup_{u\in\mathbb{R}}\mathbb{P}\big{(}|(T\underline{b})^{1/2}\mathcal{Q}_{z}-u|\leq 2l\big{)}=:\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}+\mathrm{I}_{4}.$ (SA-16)
For part $\mathrm{I}_{1}$, by (SA-12), we have
$|\mathcal{Q}-\mathcal{Q}_{\epsilon,m}|\leq|\mathcal{Q}-\mathcal{Q}_{\epsilon}|+|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|=O(\bar{b}^{2})+|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|.$
We define $l^{\prime}=N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NT)}$. Then by Lemma B.9 with $m=T/2$, we have
$\mathbb{P}\big{(}(T\underline{b})^{1/2}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq l^{\prime}\big{)}=O\big{(}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{)}.$ (SA-17)
Hence for $l=l^{\prime}+cT^{1/2}\bar{b}^{5/2},$ some constant $c>0$ large enough, we have
$\mathrm{I}_{1}=O(\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}).$
Similar argument leads to the same bound for $\mathrm{I}_{2}$ part. For part $\mathrm{I}_{3}$, by Lemma B.10 we have
$\mathrm{I}_{3}\lesssim(bT)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(bT))^{1/3}\mathrm{log}(NT)+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{4}$, denote $E=(E_{j_{1}j_{2}})_{1\leq j_{1},j_{2}\leq N}$ as the conditional covariance matrix for $(b_{j}T)^{1/2}\sum_{t=1}^{T}(a_{jt,b}\epsilon_{jt}$ $-\sum_{l=1}^{N}a_{lt,b}\epsilon_{lt}/N)$ given $\mathcal{F}_{T}.$ The diagonal entity takes the form
$E_{jj}$ $=b_{j}T\sum_{h=1-T}^{T-1}\sum_{t=1\vee(1-h)}^{T\wedge(T-h)}\mathbb{E}\Big{[}\big{(}a_{jt,b}\epsilon_{jt}-\sum_{l_{1}=1}^{N}a_{l_{1}t,b}\epsilon_{l_{1}t}/N\big{)}\big{(}a_{j(t+h),b}\epsilon_{j(t+h)}-\sum_{l_{2}=1}^{N}a_{l_{2}(t+h),b}\epsilon_{l_{2}(t+h)}/N\big{)}\Big{]}$
$\cdot(1+O(b_{j})).$
Similarly to the proof of Lemma B.1, with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for some constant $c^{\prime}>0$, we have
$\min_{1\leq j\leq N}E_{jj}\geq c^{\prime}.$
Hence by Lemma B.5, we have
$\sup_{u\in\mathbb{R}}\mathbb{P}\Big{(}(T\underline{b})^{1/2}|\mathcal{Q}_{z}-u|<2l\Big{)}\lesssim l\mathrm{log}(N)^{1/2}.$ (SA-18)
Hence
$\mathrm{I}_{4}=O\Big{(}\big{(}T^{1/2}\bar{b}^{5/2}+N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NT)}\big{)}\mathrm{log}(N)^{1/2}\Big{)}.$
By Lemma B.6, we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}(\underline{b}T)^{1/2}\mathcal{Q}_{z}\leq u\big{)}-\mathbb{P}\big{(}|Z+\tilde{\underline{d}}|_{\infty}\leq u\big{)}\Big{|}\lesssim(\underline{b}+T^{-\beta})^{1/3}\mathrm{log}(N)^{2/3}.$ (SA-19)
[…]
B.3 Proof of Theorems A.2-A.3
Proof of Theorem A.2. Let $\bar{Y}_{jt}$ (resp. $\bar{f}_{jt}$ and $\bar{e}_{jt}$) be the local linear estimator of $Y_{jt}$ (resp. $f_{j}(X_{jt})$ and $e_{jt}$) at location $X_{jt}.$ Consider
$Y_{jt}-\bar{Y}_{jt}=f_{j}(X_{jt})-\bar{f}_{jt}+e_{jt}-\bar{e}_{jt}$
Due to the continuity of $f_{j}$, we have $f_{j}(X_{jt})-\bar{f}_{jt}=O(b_{j})$. Random variables $e_{jt}$ are weakly temporal dependent, by Freedman’s inequality in Lemma B.3, we have
$\mathbb{P}\Big{(}\Big{|}\sum_{t=1}^{T}\mathbf{1}_{|X_{jt}-c_{0,j}|\leq b_{j}}\big{(}e_{jt}^{2}-\mathbb{E}(e_{jt}^{2})\big{)}\Big{|}>x\Big{)}\lesssim b_{j}Tx^{-q/2}+e^{-x^{2}/(b_{j}T)}.$
Then by taking $x=c\mathrm{log}^{1/2}(N)$ for some constant $c>0$ large enough, we have
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}(Tb_{j})^{-1/2}\Big{|}\sum_{t=1}^{T}\mathbf{1}_{|X_{jt}-c_{0,j}|\leq b_{j}}\big{(}e_{jt}^{2}-\mathbb{E}(e_{jt}^{2})\big{)}\Big{|}>x\Big{)}$
$\lesssim$ $N(\underline{b}T)^{1-q/4}\mathrm{log}^{-q/4}(N).$
Hence $\max_{1\leq j\leq N}|\widehat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|=O_{\mathbb{P}}(\bar{b}+(\underline{b}T)^{-1/2}\mathrm{log}^{1/2}(N))$. For $v_{j}^{2}$, note that
$|v_{j}^{2}-\hat{v}_{j}^{2}|=$ $(Tb_{j})\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\big{|}\sigma_{e,j}^{2}-\widehat{\sigma}_{e,j}^{2}\big{|}$
$\leq$ $(Tb_{j})\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\max_{1\leq j\leq N}|\widehat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|.$
Since $\sum_{t=1}^{T}(w_{jt,b}^{+}-w_{jt,b}^{-})^{2}\lesssim(T\underline{b})^{-1}$, we complete the proof. For case $(ii)$, follow a similar argument as in the proof of Lemma B.8, we can show the desired result in view of
$\mathbb{P}\Big{(}\max_{j}(Tb_{j})^{-1/2}\Big{|}\sum_{t=1}^{T}\mathbf{1}_{|X_{jt}-c_{0,j}|\leq b_{j}}\big{(}e_{jt}^{2}-\mathbb{E}(e_{jt}^{2})\big{)}\Big{|}>x\Big{)}\lesssim Ne^{-x^{2}/(x+1)}.$
[…]
$\leq$ $\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\hat{\gamma}_{j}-\gamma_{j}\right|/\hat{v}_{j}+\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\gamma_{j}\right|/\hat{v}_{j}$
$\lesssim$ $\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\hat{\gamma}_{j}-\gamma_{j}\right|/v_{j}+\max_{j\in\left[N\right]}\left(Tb_{j}\right)^{1/2}\left|\gamma_{j}\right|:=\text{I}_{1}+\text{I}_{2}.$
[…]
B.4 Proof of Theorem 3.2
Proof.
Denote
$\mathcal{I}_{\epsilon}^{C}=\max_{1\leq i\leq K}\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}(c_{i})\big{[}\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}\big{]}/v_{j}(c_{i})\Big{|}.$ (SA-20)
Let $\mathcal{I}_{\epsilon,m}^{C}$ be $\mathcal{I}_{\epsilon}^{C}$ with $\epsilon_{jt}$ therein replaced by $\epsilon_{jt,m}.$ Let $m=T/2.$ For any $l>0,$ similar to (SA-11), we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\big{(}\mathcal{I}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|\leq u\big{)}\Big{|}$
$\leq$ $\mathbb{P}\Big{(}|\mathcal{I}^{C}-\mathcal{I}^{C}_{\epsilon,m}|\geq l\Big{)}+\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(\mathcal{I}_{\epsilon,m}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|\leq u)\Big{|}$
$+\sup_{u\in\mathbb{R}}\mathbb{P}\big{(}|Z^{C}-u|\leq l\big{)}=:\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}.$ (SA-21)
For part $\mathrm{I}_{1}$, by (SA-12), we have
$|\mathcal{I}^{C}-\mathcal{I}^{C}_{\epsilon,m}|\leq|\mathcal{I}^{C}-\mathcal{I}^{C}_{\epsilon}|+|\mathcal{I}^{C}_{\epsilon}-\mathcal{I}^{C}_{\epsilon,m}|=O(\bar{b}^{2})+|\mathcal{I}^{C}_{\epsilon}-\mathcal{I}^{C}_{\epsilon,m}|.$
We define $l^{\prime}=N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NKT)}$. Then by a similar argument as in Lemma B.9 with $m=T/2$, we have
$\mathbb{P}\big{(}|\mathcal{I}^{C}_{\epsilon}-\mathcal{I}^{C}_{\epsilon,m}|\geq l^{\prime}\big{)}=O\big{(}\mathrm{log}(NKT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{)}.$ (SA-22)
Hence for $l=l^{\prime}+cT^{1/2}\bar{b}^{5/2},$ some constant $c>0$ large enough, we have
$\mathrm{I}_{1}=O(\mathrm{log}(NKT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}).$
For part $\mathrm{I}_{2}$, similar to Lemma B.10 and (SA-18), we have
$\mathrm{I}_{2}$ $\lesssim(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NKT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NKT)$
$+\big{(}T^{1/2}\bar{b}^{5/2}+(NK)^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NKT)}\big{)}\mathrm{log}(NK)^{1/2}+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{3},$ by Lemma B.5, we have
$\sup_{u\in\mathbb{R}}\mathbb{P}\Big{(}|Z^{C}-u|<2l\Big{)}\lesssim l\mathrm{log}(NK)^{1/2}.$ (SA-23)
Desired result follows by combining $\mathrm{I}_{1}$-$\mathrm{I}_{3}$ parts. ∎
B.5 Proof of Theorem 3.3
For any $c_{0j}-b_{j}<c\leq c_{0j}$, we can decompose the difference into three parts
$\hat{\gamma}_{j}(c)-\hat{\gamma}_{j}(c_{0j})$
$=$ $\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}Y_{jt}$
$=$ $\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}f_{j}(X_{jt})+\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$+\sum_{t=1}^{T}\big{(}w_{jt,b}(c)-w_{jt,b}(c_{0j})\big{)}e_{jt}=:\mathrm{I}_{1j}+\mathrm{I}_{2j}+\mathrm{I}_{3j}.$
For part $\mathrm{I}_{1j}$, due to the smoothness of the function $f_{j}(\cdot)$,
$\sum_{t=1}^{T}w_{jt,b}(c)f_{j}(X_{jt})\lesssim b_{j}.$
For part $\mathrm{I}_{2j},$ since $\sum_{t=1}^{T}w_{jt,b}^{+}(c)=\sum_{t=1}^{T}w_{jt,b}^{-}(c)=1,$ we have
$\sum_{t=1}^{T}\big{(}w_{jt,b}(c_{0j})-w_{jt,b}(c)\big{)}\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$=$ $\gamma_{j}\sum_{t}\Big{[}w_{jt,b}(c_{0j})-w_{jt,b}(c-c_{0j}+c_{0j})\Big{]}\textbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$\lesssim$ $\gamma_{j}\sum_{t}\Big{[}\frac{1}{Tb_{j}}\frac{|c-c_{0j}|}{b_{j}}\Big{]}\textbf{1}_{\{|X_{jt}-c_{0j}|<b_{j}\}}$
$\asymp$ $\gamma_{j}|c-c_{0j}|/b_{j}.$
For part $\mathrm{I}_{3j},$ consider the observations in region
$\mathcal{A}_{1}:=\Big{\{}t:X_{jt}\in[c_{0j}-b_{j},c)\Big{\}}.$
In this region, both $w_{jt,b}(c)$ and $w_{jt,b}(c_{0j})$ are non-zero. By Lemma B.2, the variance of $e_{jt}$ in this region can be bounded as
$\mathrm{Var}\Big{(}\sum_{t\in\mathcal{A}_{1}}(w_{jt,b}(c)-w_{jt,b}(c_{0j}))e_{jt}\Big{)}\lesssim(Tb_{j})^{-2}(1/g_{j}(c_{0j})-1/g_{j}(c))^{2}Tb_{j}g_{j}(c_{0j})\lesssim(Tb_{j})^{-1}|c-c_{0j}|^{2}b_{j}^{-2}.$
Consider another kind of region
$\mathcal{A}_{2}:=\Big{\{}t:X_{jt}\in[c-b_{j},c_{0j}-b_{j})\Big{\}}.$
In region $\mathcal{A}_{2},$ different from $\mathcal{A}_{1},$ only $w_{jt,b}(c)$ is non-zero. The variance of noise in this region follows
$\mathrm{Var}\Big{(}\sum_{t\in\mathcal{A}_{2}}(w_{jt,b}(c)-w_{jt,b}(c_{0j}))e_{jt}\Big{)}\lesssim(Tb_{j})^{-2}g_{j}(c_{0j})^{-2}|c-c_{0j}|Tg_{j}(c_{0j})\lesssim(Tb_{j})^{-2}|c-c_{0j}|T.$
The rest regions can be similarly bounded. A similar GA can be constructed as in the proof of Theorem 3.1, which can be used to handle the maximum over $j$. Hence we have
$\max_{1\leq j\leq N}|\mathrm{I}_{3j}|=O_{\mathbb{P}}\Big{(}(Tb_{j})^{-1/2}\mathrm{log}^{1/2}(N)|c-c_{0j}|/b_{j}+(Tb_{j})^{-1}\mathrm{log}^{1/2}(N)|c-c_{0j}|^{1/2}T^{1/2}\Big{)}.$
Since $\hat{\gamma}_{j}(\hat{c})$ shall be the largest, for $c=\hat{c},$
$\gamma_{j}|c-c_{0j}|/b_{j}\lesssim(Tb_{j})^{-1/2}\mathrm{log}^{1/2}(N)|c-c_{0j}|/b_{j}+(Tb_{j})^{-1}\mathrm{log}^{1/2}(N)|c-c_{0j}|^{1/2}T^{1/2},$
which can only be achieved if the following bound uniformly over $j$,
$|c-c_{0j}|=O_{\mathbb{P}}(\mathrm{log}(N)/(\gamma_{j}^{2}T)).$
B.6 Proofs for the Unknown Threshold Case
Proof.
(Proof of Theorem A.4) Recall that $\hat{e}_{jt}=e_{jt}+f_{j}(X_{jt})-\hat{f}_{j,b^{*}}(X_{jt})$ for the observations with $|X_{jt}-c_{i}|\leq b^{*}$ where $b^{*}\ll\underline{b}$. For $c_{i}$ satisfying $\min_{1\leq i\leq K}|c_{i}-c_{0j}|\leq b^{*},$ we have
$\hat{\sigma}_{e,j}^{2}(c_{i})=$ $\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j},|X_{jt}-c_{0j}|\leq b^{*}\}}\phi_{\acute{a}}(\hat{e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}+\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|>b^{*},|X_{jt}-c_{i}|\leq b_{j}\}}\phi_{\acute{a}}(\hat{e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}$
$=$ $:\mathrm{I}_{1ij}+\mathrm{I}_{2ij}.$
By a similar argument as in Lemma B.2, we have for $x\gtrsim\sqrt{T\mathrm{log}(NK)},$
$\mathbb{P}\Big{(}\max_{1\leq i\leq K,1\leq j\leq N}\Big{|}\sum_{t=1}^{T}\big{(}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}-\mathbb{E}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}\big{)}\Big{|}>x\Big{)}$
$\lesssim$ $T^{(p/2-\alpha p)\vee 1}\mathrm{log}(NK)^{p/2}/x^{p}+e^{-x^{2}/T}.$ (SA-24)
Same argument can be applied on $\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j},|X_{jt}-c_{0j}|\leq b^{*}\}}$. For part $\mathrm{I}_{1ij},$ since $|\phi_{\acute{a}}|\leq\acute{a}$, by the above bounds we have
$\max_{1\leq i\leq K,1\leq j\leq N}|\mathrm{I}_{1ij}|\leq\max_{1\leq i\leq K,1\leq j\leq N}\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j},|X_{jt}-c_{0j}|\leq b^{*}\}}\acute{a}}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b})$
For $\mathrm{I}_{2ij}$ part, since $\max_{i,j}|f_{j}(X_{jt})-\hat{f}_{j,b^{*}}(X_{jt})|=O(b^{*2}),$ we have $\max_{i,j}|\phi_{\acute{a}}({e}_{jt}^{2})-\phi_{\acute{a}}(\hat{e}_{jt}^{2})|=O(b^{*2})$ and
$\max_{1\leq i\leq K,1\leq j\leq N}|\mathrm{I}_{2ij}|$
$\lesssim$ $b^{*2}+\max_{1\leq i\leq K,1\leq j\leq N}\Big{|}\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{0j}|>b^{*},|X_{jt}-c_{i}|\leq b_{j}\}}\phi_{\acute{a}}({e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}-\mathbb{E}(\phi_{\acute{a}}({e}_{jt}^{2})|X_{jt}=c_{i})\Big{|}$
$+\max_{1\leq i\leq K,1\leq j\leq N}\big{|}\mathbb{E}(\phi_{\acute{a}}({e}_{jt}^{2})|X_{jt}=c_{i})-\mathbb{E}({e}_{jt}^{2}|X_{jt}=c_{i})\big{|}=:b^{*2}+\mathrm{II}_{1}+\mathrm{II}_{2}.$
Similarly to (B.6), we have $\mathrm{II}_{1}=O_{\mathbb{P}}((Tb_{j})^{-1/2}\mathrm{log}^{1/2}(NK))$. For the $\mathrm{II}_{2}$ part, by Markov’s inequality we have $\mathrm{II}_{2}=O(\acute{a}^{-p/2+1}).$
Proof.
(Proof of Theorem A.5) The proof of Theorem A.5 follows directly from the proof of Theorem A.3, thus we omit the proof. ∎
B.7 Proofs for the Derivatives
First we introduce a few technical lemmas.
Lemma B.11.
(Properties for weights in derivatives) Under Assumptions 3.1-3.3 and SA-1-SA-3, the followings hold:
[…]
(ii) For $l=0,1,2,$ on set $\mathcal{A}_{T}$, with $\mathcal{A}_{T}$ defined in Lemma B.2,
$\max_{1\leq j\leq N}\Big{|}w_{jt,b}^{+[1]}-\frac{K((X_{jt}-c_{j0})/b_{j})}{g_{j}(c_{j0})Tb^{2}_{j}}\frac{K_{1}^{+}-K_{0}^{+}(X_{jt}-c_{j0})/b_{j}}{-K_{2}^{+}K_{0}^{+}+K_{1}^{+2}}\mathbf{1}_{\{X_{jt}\geq c_{j0}\}}\Big{|}$
$\lesssim$ $\sqrt{\mathrm{log}(TN)/(T^{3}\underline{b}^{5})}+1/(T\underline{b}),$ (SA-25)
and moreover,
$\sum_{t=1}^{T}|w_{jt,b}^{+}|=O(1/\underline{b})\quad\text{and}\quad\sum_{t=1}^{T}|w_{jt,b}^{-}|=O(1/\underline{b})$ (SA-26)
uniformly over $j$.
[…]
Lemma B.12 ($m$-dependent approximation for derivative case).
For some $1\ll m\leq T/2,$
  • (i)
    assume that the conditions in Theorem A.6 $(i)$ hold, we have
    $\mathbb{P}\big{(}|I_{\epsilon}^{[1]}-I_{\epsilon,m}^{[1]}|\geq z\big{)}\lesssim N\mathrm{log}^{q}(NT)(T\underline{b})^{1-q/2}m^{-q\beta}z^{-q}+Ne^{-z^{2}m^{2\beta}}+\mathbb{P}(\mathcal{A}_{T}^{c}).$
  • (ii)
    assume that the conditions in Theorem A.6 $(ii)$ hold, we have
    $\mathbb{P}\big{(}|I_{\epsilon}^{[1]}-I_{\epsilon,m}^{[1]}|\geq z\big{)}\lesssim N\mathrm{exp}\Big{\{}-c\frac{z^{2}}{m^{-\beta}(\underline{b}T)^{-1/2}z+m^{-2\beta}}\Big{\}}+\mathbb{P}(\mathcal{A}_{T}^{c}),$
where $\mathcal{A}_{T}$ is defined in Lemma B.2.
Proof of Lemma B.12. We only show case $(i)$, for the case $(ii)$ the result can be similary derived. Let
$r_{j}^{[1]}(X_{jt},U_{jt})=\sigma_{j}(X_{jt},U_{jt})/v_{j}^{[1]}$ (SA-27)
and
$\xi_{jk}^{[1]}=(Tb_{j}^{3})^{1/2}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}^{[1]}r_{j}^{[1]}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\eta_{k}.$ (SA-28)
Similar to (SA-22), we have
$|I_{\epsilon}^{[1]}-I_{\epsilon,m}^{[1]}|\mathbf{1}_{\mathcal{A}_{T}}\leq\max_{1\leq j\leq N}\Big{|}\sum_{k\leq T-m}\xi_{jk}^{[1]}\Big{|}.$
Then by (SA-27), we have
$\sigma^{[1]2}$ $=\max_{1\leq j\leq N}\sum_{k\leq T-m}\mathbb{E}(\xi_{jk}^{[1]2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$
$\lesssim\max_{1\leq j\leq N}Tb_{j}^{3}\sum_{k\leq T-m}\Big{(}\sum_{t=1\vee(k+m)}^{T}|w_{jt,b}^{[1]}r_{j}^{[1]}(X_{jt},U_{jt})A_{t-k,j,\cdot}|_{2}\Big{)}^{2}$
$\lesssim m^{-2\beta}.$ (SA-29)
By Lemma B.3, for any $z>0,$
$\mathbb{P}\Big{(}|I_{\epsilon}^{[1]}-I_{\epsilon,m}^{[1]}|\mathbf{1}_{\mathcal{A}_{T}}\geq z\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}$
$\lesssim$ $\sum_{k\leq T-m}\sum_{1\leq j\leq N}\mathbb{P}\Big{(}|\xi_{jk}^{[1]}|\geq u|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}+2N\mathrm{exp}\Big{(}-\frac{z^{2}}{zu+m^{-2\beta}}\Big{)}.$ (SA-30)
By Markov’s inequality, we have
$\sum_{k\leq T-m}\mathbb{P}\Big{(}|\xi_{jk}^{[1]}|\geq u\Big{)}$ $\leq\sum_{k\leq T-m}u^{-q}\mathbb{E}\big{(}\mathbb{E}(|\xi_{jk}^{[1]}|^{q}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\big{)}$
$\lesssim u^{-q}\sum_{k\leq T-m}(Tb_{j}^{3})^{q/2}\mathbb{E}\Big{|}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}^{[1]}r_{j}^{[1]}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\Big{|}_{2}^{q}$
$\lesssim(T\underline{b})^{1-q/2}m^{-q\beta}u^{-q}.$ (SA-31)
[…]
Lemma B.13.
Let $T/2\geq m\rightarrow\infty.$
$(i)$ Suppose that the conditions in Theorem A.6 $(i)$ hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(I_{\epsilon,m}^{[1]}\leq u)-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j}^{3})^{1/2}\gamma_{j}^{[1]}/v_{j}^{[1]}\big{|}\leq u\Big{)}\Big{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)},$
where the constant in $\lesssim$ is independent of $N,T.$
$(ii)$ Suppose that the conditions in Theorem A.6 $(ii)$ hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}(I_{\epsilon,m}^{[1]}\leq u)-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j}^{3})^{1/2}\gamma^{[1]}_{j}/v_{j}^{[1]}\big{|}\leq u\Big{)}\Big{|}$
$\leq$ $(\underline{b}T)^{-1/6}\mathrm{log}(NT)^{7/6}+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)},$
where the constant in $\lesssim$ is independent of $N,T.$
Proof of Lemma B.13. Note that
$I_{\epsilon,m}^{[1]}\mathbf{1}_{\mathcal{A}_{T}}$ $=\max_{1\leq j\leq N}\frac{(Tb_{j}^{3})^{1/2}}{v_{j}^{[1]}}\Big{|}\sum_{t=1}^{T}w_{jt,b}^{[1]}\big{[}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}+\varepsilon_{j}(X_{jt},U_{jt})+f_{j}(X_{jt})\big{]}\Big{|}$
$=\max_{1\leq j\leq N}\frac{(Tb_{j}^{3})^{1/2}}{v_{j}^{[1]}}\Big{|}\sum_{k=2-m}^{T}\xi_{jk,1}^{[1]}+\sum_{t=1}^{T}\xi_{jt,2}^{[1]}+\gamma_{j}^{[1]}+O(\bar{b})\Big{|},$ (SA-32)
where $\xi_{jk,1}^{[1]}$ and $\xi_{jk,2}^{[1]}$ are the same as $\xi_{jk,1}$ and $\xi_{jk,2}$ in (SA-3) respectively. By Lemma B.11, $b_{j}w_{jt,b}^{[1]}$ has similar performance as $w_{jt,b}$, hence similar to (SA-8), we have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}I_{\epsilon,m}^{[1]}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j}^{3})^{1/2}\gamma_{j}^{[1]}/v_{j}^{[1]}\big{|}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$ (SA-33)
[…]
Then we show the proofs of the main theorems in this subsections.
Proof of Theorem A.6.
$(i)$ Consider the notation
$I_{\epsilon}^{[1]}=\max_{1\leq j\leq N}(Tb_{j}^{3})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+[1]}-w_{jt,b}^{-[1]})\big{(}\varepsilon_{j}(X_{jt},U_{jt})+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}\big{)}+\gamma_{j}^{[1]}\Big{|}/v_{j}^{[1]}.$ (SA-34)
Let $I_{\epsilon,m}^{[1]}$ be $I_{\epsilon}^{[1]}$ with $\epsilon_{jt}$ replaced by $\epsilon_{jt,m}.$ Similarly, let $I_{z}^{[1]}$ (resp. $I_{z,m}^{[1]}$) be $I_{\epsilon}$ (resp. $I_{\epsilon,m}$) with $\eta_{t}$ therein replaced by $z_{t}$, where $\left\{z_{t}\right\}_{t\in\mathbb{Z}}$ are i.i.d. Gaussian vectors with zero mean and identity covariance matrix.
Similar to the proof of Theorem 3.1, for $m=T/2$, by (SA-11) we have
$\sup_{u\in\mathbb{R}}\Big{[}\mathbb{P}\big{(}\mathcal{I}^{[1]}\leq u)-\mathbb{P}(I_{z}^{[1]}\leq u\big{)}\Big{]}$
$\leq$ $\mathbb{P}\big{(}|\mathcal{I}^{[1]}-I_{\epsilon,m}^{[1]}|\geq l\big{)}+\mathbb{P}\big{(}|I_{z}^{[1]}-I_{z,m}^{[1]}|\geq l\big{)}+\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(I_{\epsilon,m}^{[1]}\leq u)-\mathbb{P}(I_{z,m}^{[1]}\leq u)\big{|}$
$+\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}\big{(}I_{z}^{[1]}\leq u+2l\big{)}-\mathbb{P}\big{(}I_{z}^{[1]}\leq u\big{)}\big{|}$
$=$ $:\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}+\mathrm{I}_{4}.$ (SA-35)
For part $\mathrm{I}_{1}$, we have $|\mathcal{I}^{[1]}-I_{\epsilon,m}^{[1]}|\leq|\mathcal{I}^{[1]}-I_{\epsilon}^{[1]}|+|I_{\epsilon}^{[1]}-I_{\epsilon,m}^{[1]}|.$ By Assumption SA-2 and Lemma B.11, we have
$\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}w_{jt,b}^{+[1]}f_{j}(X_{jt})-\partial_{+}f_{j}(c_{j0})\Big{|}$
$=$ $\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}w_{jt,b}^{+[1]}\big{[}f_{j}(c_{j0})+\partial_{+}f_{j}(c_{j0})(X_{jt}-c_{j0})+O((X_{jt}-c_{j0})^{2})\big{]}-\partial_{+}f_{j}(c_{j0})\Big{|}$
$=$ $O(\bar{b}).$
Thus
$|\mathcal{I}^{[1]}-I_{\epsilon}^{[1]}|\leq\max_{1\leq j\leq N}(Tb_{j}^{3})^{1/2}\Big{|}\sum_{t=1}^{T}(w_{jt,b}^{+[1]}-w_{jt,b}^{-[1]})f_{j}(X_{jt})\Big{|}/v_{j}^{[1]}=O(T^{1/2}\bar{b}^{5/2}).$ (SA-36)
We define $l^{\prime}=N^{1/q}T^{-\beta}\sqrt{\mathrm{log}(NT)}$. Then by Lemma B.12 with $m=T/2$, we have
$\mathbb{P}\big{(}|I_{\epsilon}^{[1]}-I_{\epsilon,m}^{[1]}|\geq l^{\prime}\big{)}=O\big{\{}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{\}}.$ (SA-37)
Hence for $l=l^{\prime}+cT^{1/2}\bar{b}^{5/2}$, where $c>0$ is some constant large enough, we have
$\mathrm{I}_{1}=O\big{\{}\mathrm{log}(NT)^{q/2}(T\underline{b})^{-q/2+1}+T^{-(\alpha p)\wedge(p/2-1)}\big{\}}.$
Similar argument leads to the same bound for part $\mathrm{I}_{2}$.
For part $\mathrm{I}_{3}$, by Lemma B.13 we have
$\mathrm{I}_{3}\lesssim(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+T^{-\beta})^{1/3}\mathrm{log}^{2/3}(N)+T^{-(\alpha p)\wedge(p/2-1)}.$
For part $\mathrm{I}_{4}$, we use Lemma B.5 to obtain
$\mathrm{I}_{4}\leq\sup_{u\in\mathbb{R}}\mathbb{P}\big{(}|I_{z}^{[1]}-u\big{|}\leq 2l\big{)}\lesssim l\mathrm{log}(N)^{1/2}.$
We complete the proof by combining the $\mathrm{I}_{1}$-$\mathrm{I}_{4}$ parts and a similar argument for the other side of the inequality.
[…]
Proof of Theorems A.7A.9 Theorems A.7, A.8 and A.9 are derived from Theorems 3.2, A.2 and A.3 respectively using arguments similar to those in the proof of Theorem A.6.
Appendix C Proofs of the Technical Lemmas in Appendix B
In this section, we prove the technical lemmas in Section B.
C.1 Proofs of Lemma B.1 in Section B.1.1
Proof of Lemma B.1.
(i). Recall ~$\tilde{\mathcal{F}}_{t}$ in (3.1). Define
$w_{jt,b}=(w_{jt,b}^{+}-w_{jt,b}^{-})\mathbf{1}_{\mathcal{A}_{T}}\text{ and }r_{j}(X_{jt},U_{jt})=\sigma_{j}(X_{jt},U_{jt})/v_{j}.$ (SA-1)
Note that $I_{\epsilon,m}$ on $\mathcal{A}_{T}$ can be rewritten as
$I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}$ $=\max_{1\leq j\leq N}\frac{(Tb_{j})^{1/2}}{v_{j}}\Big{|}\sum_{t=1}^{T}w_{jt,b}\big{[}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}+\varepsilon_{j}(X_{jt},U_{jt})\big{]}+\gamma_{j}\Big{|}$
$=\max_{1\leq j\leq N}\frac{(Tb_{j})^{1/2}}{v_{j}}\Big{|}\sum_{k=2-m}^{T}\xi_{jk,1}+\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\Big{|},$ (SA-2)
where
$\xi_{jk,1}=\Big{(}\sum_{t=1\vee k}^{T\wedge(k+m-1)}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\Big{)}\eta_{k}\quad\text{and}\quad\xi_{jt,2}=w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt}).$ (SA-3)
In the following we shall show that for $B_{T}=c\underline{b}^{-1/2}$ with some constant $c>0$ being large enough:
  1. (a)
    with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for some constant $c^{\prime}>0,$ we have
    $\min_{1\leq j\leq N}(Tb_{j})\sum_{k=2-m}^{T}\mathbb{E}\big{(}\xi_{jk,1}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}\geq c^{\prime};$
  2. (b)
    with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for $l=1,2,$ we have
    $\max_{1\leq j\leq N}T^{1+l}b_{j}^{(2+l)/2}\sum_{k=2-m}^{T}\mathbb{E}(|\xi_{jk,1}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\leq B_{T}^{l}.$
Since $|w_{jt,b}|\lesssim(Tb_{j})^{-1}$, we have $\max_{k}\mathbb{E}(\max_{j}|\xi_{jk,1}|^{q}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\lesssim(Tb_{j})^{-q},$ and thus
$\mathbb{E}\left(\max_{1\leq j\leq N}(Tb_{j}^{1/2}|\xi_{jk,1}|/B_{T})^{q}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\right)\leq 2,$
for all $k.$ Let $\tilde{Z}_{j,1}$, $1\leq j\leq N$, be Gaussian distributed random variables with the same conditional covariance structure as
$(Tb_{j})^{1/2}\sum_{k=2-m}^{T}\xi_{jk,1}/v_{j}$
given $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$. For any fixed $j$, conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T},$ $\xi_{jk,1}$’s are independent for different $k$. Thus combining the above with (a) and (b), by Proposition 2.1 in Chernozhukov et al., 2017b , we have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}\tilde{Z}_{j,1}+(Tb_{j})^{1/2}v_{j}^{-1}\big{(}\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\big{)}\big{|}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT),$
where the conditional variance of $\tilde{Z}_{j,1}$ follows,
$\mathbb{E}(\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}))$
$=$ $(Tb_{j})\mathbb{E}\Big{(}\sum_{t=1}^{T}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt,m}/v_{j}\Big{)}^{2}$
$=$ $\mathbb{E}(v_{j,1}^{2})+(Tb_{j})\sum_{1\leq t_{1}\neq t_{2}\leq T}\mathbb{E}\Big{(}w_{jt_{1},b}\sigma_{j}(X_{jt_{1}},U_{jt_{1}})w_{jt_{2},b}\sigma_{j}(X_{jt_{2}},U_{jt_{2}})/v_{j}^{2}\Big{)}\mathbb{E}\big{(}\epsilon_{jt_{1},m}\epsilon_{jt_{2},m}\big{)}+O(m^{-\beta})$
$=$ $\mathbb{E}(v_{j,1}^{2})+O(b_{j}+m^{-\beta}),$
with
$v_{j,1}^{2}=(Tb_{j})\sum_{t=1}^{T}w_{jt,b}^{2}\mathrm{Var}\big{(}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}|X_{jt}\big{)}/v_{j}^{2}.$ (SA-4)
For any $j_{1}\neq j_{2},$ we have $\mathbb{E}(\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j_{2},1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}))=O(\bar{b}).$ Consider the random variable
$Z_{j,1}=(Tb_{j})^{1/2}v_{j}^{-1}\sum_{t=1}^{T}\tilde{\xi}_{jt,1},$ (SA-5)
where $\tilde{\xi}_{jt,1}$ are i.i.d. from $N(0,v_{j,1}^{2}v_{j}^{2}/(T^{2}b_{j}))$ and independent from ~$\tilde{\mathcal{F}}_{T}$ and $\epsilon_{jt}$.
By a similar argument as in the proof of part (a) below, with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$ the maximum difference of $\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ (resp. $\mathrm{Var}(Z_{j,1}|\mathcal{F}_{T})$) and $\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j2,1}|$ $\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ (resp. $\mathrm{Cov}(Z_{j_{1},1},Z_{j2,1}|\mathcal{F}_{T})$) from their expectation are bounded by $\bar{b}+\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$ Hence by Lemma B.6, we further have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\Big{|}Z_{j,1}+(Tb_{j})^{1/2}v_{j}^{-1}\big{(}\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\big{)}\Big{|}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$
Then by the construction of $Z_{j,1}$ in (SA-5), we have
$Z_{j,1}+(Tb_{j})^{1/2}v_{j}^{-1}\Big{(}\sum_{t=1}^{T}\xi_{jt,2}+\gamma_{j}\Big{)}=(Tb_{j})^{1/2}v_{j}^{-1}\Big{(}\sum_{t=1}^{T}(\tilde{\xi}_{jt,1}+\xi_{jt,2})+\gamma_{j}\Big{)}.$
Given $\mathcal{F}_{T}$, $\tilde{\xi}_{jt,1}+\xi_{jt,2}$ are independent for different $t.$ Recall the definition of $v_{j}^{2}$ in (2.14), we have
$(Tb_{j})v_{j}^{-2}\sum_{t=1}^{T}\mathrm{Var}\big{(}\tilde{\xi}_{jt,1}+\xi_{jt,2}|\mathcal{F}_{T}\big{)}=1.$ (SA-6)
Since $|\varepsilon_{j}(\cdot,\cdot)|$ is upper bounded, $\max_{t}\mathbb{E}(|\tilde{\xi}_{jt,1}|+|\xi_{jt,2}||\mathcal{F}_{T})=O((b_{j}T)^{-1}).$ With probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for $l=1,2,$
$\max_{1\leq j\leq N}T^{1+l}b_{j}^{(2+l)/2}\sum_{t=1}^{T}\mathbb{E}\Big{(}|(\tilde{\xi}_{jt,1}+\xi_{jt,2})/v_{j}|^{2+l}|\mathcal{F}_{T}\Big{)}$
$\lesssim$ $\max_{1\leq j\leq N}Tb_{j}^{(2-l)/2}\sum_{t=1}^{T}\mathbb{E}\Big{(}((\tilde{\xi}_{jt,1}+\xi_{jt,2})/v_{j})^{2}|\mathcal{F}_{T}\Big{)}\leq B_{T}^{l}.$ (SA-7)
For all $k$, we have
$\mathbb{E}\Big{(}\max_{1\leq j\leq N}(b_{j}^{1/2}Tv_{j}^{-1}|\tilde{\xi}_{jt,1}+\xi_{jt,2}|/B_{T})^{q}\big{|}\mathcal{F}_{T}\Big{)}\leq 2.$
By combining the above results with (SA-6) and (SA-7), Proposition 2.1 in Chernozhukov et al., 2017b , and Lemma B.6, we have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}I_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}+(Tb_{j})^{1/2}v_{j}^{-1}\gamma_{j}\big{|}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$ (SA-8)
Hence we complete the proof.
Proof of part (a). Since $\mathbb{E}(\eta_{jk}^{2})=1,$ we have
$\sum_{k=2-m}^{T}\mathbb{E}(\xi_{jk,1}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ $\geq\sum_{k=1}^{T-m}\Big{|}\sum_{t=k}^{k+m-1}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}\Big{|}_{2}^{2}$
$=\sum_{k=1}^{T-m}\Big{|}\sum_{t=k}^{k+m-1}(M_{jt,1}+D_{jt,1})A_{t-k,j,\cdot}\Big{|}_{2}^{2},$ (SA-9)
where
$D_{jt,1}=\mathbb{E}\big{(}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})|\mathcal{F}_{t-1}\big{)}\quad\text{and}\quad M_{jt,1}=w_{jt,b}\sigma_{j}(X_{jt},U_{jk})-D_{jt,1}.$ (SA-10)
For the simplicity of notation, denote
$Y_{jk}=\Big{|}\sum_{t=k}^{k+m-1}M_{jt,1}A_{t-k,j,\cdot}\Big{|}_{2}^{2}.$ (SA-11)
Since $|D_{jt,1}|\lesssim 1/T,$ (SA-9) can be further lower bounded by
$\sum_{k=2-m}^{T}\mathbb{E}(\xi_{jk,1}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\geq\sum_{k=1}^{T-m}\Big{(}Y_{jk}^{1/2}-O(1/T)\Big{)}^{2}\geq\sum_{k=1}^{T-m}Y_{jk}-O\Big{(}\Big{(}\sum_{k=1}^{T-m}Y_{jk}/T\Big{)}^{1/2}\Big{)}.$ (SA-12)
We shall bound the difference between $Y_{jk}$ and $\mathbb{E}Y_{jk}$. Note that $\mathbb{E}_{0}Y_{jk}$ can be rewritten as
$\sum_{k=1}^{T-m}\mathbb{E}_{0}Y_{jk}$ $=\sum_{t=1}^{T}\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}(M_{jt,1}L_{k,t,1}+\mathbb{E}_{0}L_{k,t,2})$
$=\sum_{t=1}^{T}\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\big{[}M_{jt,1}L_{k,t,1}+L_{k,t,2}-\mathbb{E}(L_{k,t,2}|\mathcal{F}_{t-1})\big{]}$
$+\sum_{t=1}^{T}\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\big{[}\mathbb{E}(L_{k,t,2}|\mathcal{F}_{t-1})-\mathbb{E}(L_{k,t,2})\big{]}=:\mathrm{I}_{j1}+\mathrm{I}_{j2},$
where
$L_{k,t,1}=2\sum_{s=k}^{t-1}M_{js,1}A_{t-k,j,\cdot}^{\top}A_{s-k,j,\cdot}\quad\text{and}\quad L_{k,t,2}=M_{jt,1}^{2}|A_{t-k,j,\cdot}|_{2}^{2}.$
Let
$\varsigma_{jt}:=\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\big{[}M_{jt,1}L_{k,t,1}+L_{k,t,2}-\mathbb{E}(L_{k,t,2}|\mathcal{F}_{t-1})\big{]}.$
Then $\mathrm{I}_{j1}=\sum_{t=1}^{T}\varsigma_{jt}$, the summation of martingale differences. Since $|M_{jt,1}|=O((Tb_{j})^{-1})$, we have $|\varsigma_{jt}|=O((Tb_{j})^{-2})$ and $\mathbb{E}(\varsigma_{jt}^{2}|\mathcal{F}_{t-1})=O(b_{j}(Tb_{j})^{-4})$. Thus by Freedman’s inequality,
$\mathbb{P}\big{(}b_{j}^{-1/2}(Tb_{j})^{2}|\mathrm{I}_{j1}|\geq z\big{)}\lesssim\mathrm{exp}\Big{\{}-\frac{z^{2}}{b_{j}^{-1/2}z+T}\Big{\}}.$ (SA-13)
Hence with probability greater than $1-T^{-p}$ we have
$\max_{1\leq j\leq N}(Tb_{j})|\mathrm{I}_{j1}|\lesssim\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$ (SA-14)
For $\mathrm{I}_{j2}$ part, let $\tilde{\sigma}_{j}^{2}(x)=\mathbb{E}(\sigma_{j}^{2}(X_{jt},U_{jk})|X_{jt}=x)$. By Assumption SA-2,
[…]
By Theorem 6.2 in Zhang and Wu, (2017), we have
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}(b_{j}T^{2})|\mathrm{I}_{j2}|\geq z+O(Tb_{j})\Big{)}\lesssim z^{-p}T^{(p/2-\alpha p)\vee 1}\mathrm{log}(N)^{p/2}+e^{-z^{2}/T}.$
Hence with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$ we have
$\max_{1\leq j\leq N}(Tb_{j})|\mathrm{I}_{j2}|\lesssim\sqrt{\mathrm{log}(NT)/T}+\bar{b}.$ (SA-16)
For the expectation part, since $M_{jt,1}$ are martingale differences, for $m\leq T/2,$
$\frac{Tb_{j}}{v_{j}^{2}}\sum_{k=1}^{T-m}\mathbb{E}(Y_{jk})=\frac{Tb_{j}}{v_{j}^{2}}\sum_{k=1}^{T-m}\sum_{t=k}^{k+m-1}\mathbb{E}(M_{jt,1}^{2})|A_{t-k,j,\cdot}|_{2}^{2}\asymp c,$
where $c>0$ is some constant independent of $j$. We complete the proof by combining above with (SA-14) and (SA-16).
Proof of part (b). Recall the definition of $Y_{jk}$ in (SA-11). Since $|Y_{jk}|\lesssim(Tb_{j})^{-2}$, we have
$\sum_{k=2-m}^{T}\mathbb{E}(|\xi_{jk,1}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ $\lesssim\sum_{k=2-m}^{T}\Big{|}\sum_{t=1\vee k}^{T\wedge(k+m-1)}w_{jt,b}\sigma_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}\Big{|}_{2}^{2+l}$
$\lesssim\sum_{k=2-m}^{T}Y_{jk}^{1+l/2}+\sum_{k=2-m}^{T}\Big{(}\sum_{t=1\vee k}^{T\wedge(k+m-1)}|D_{jt,2}||A_{t-k,j,\cdot}|_{2}\Big{)}^{2+l}$
$\lesssim(Tb_{j})^{-l}\sum_{k=2-m}^{T}Y_{jk}+T^{-1-l}.$ (SA-17)
By a similar argument in $(a)$, with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, we have
$\Big{|}\sum_{k=2-m}^{T}Y_{jk}\Big{|}\lesssim(Tb_{j})^{-1}.$ (SA-18)
Then part $(b)$ follows from (SA-17) and (SA-18).
(ii) Since
$\Big{|}\sum_{t=1\vee k}^{T\wedge(k+m-1)}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\Big{|}_{2}=O(1/(b_{j}T)),$
we have
$\mathbb{E}\Big{(}\mathrm{exp}\big{\{}T\sqrt{b_{j}}|\tilde{\xi}_{jt,1}+\xi_{jt,2}|/B_{T}\big{\}}\big{|}\mathcal{F}_{T}\Big{)}\leq 2,$
[…]
C.2 Proofs of Lemmas in Section B.1.3
Proof of Lemma B.2. $(i)$ The results can be derived by elementary calculus.
$(ii)$ Define $\mathbb{E}_{0}(X)=X-\mathbb{E}(X)$. without loss of generality and for notation simplicity, we assume $c_{j0}=0.$ For random variable $X,$ denote $X^{+}=X\mathbf{1}_{\{X\geq 0\}}.$ We shall show that $S_{jl,b}^{+}$ is close to its expectation. To this end, denote
$M_{jt}=X_{jt}^{+l}K(X_{jt}^{+}/b_{j})-\mathbb{E}(X_{jt}^{+l}K(X_{jt}^{+}/b_{j})|\mathcal{F}_{t-1})$
and $D_{jt}=\mathbb{E}(X_{jt}^{+l}K(X_{jt}^{+}/b_{j})|\mathcal{F}_{t-1})-\mathbb{E}(X_{jt}^{+l}K(X_{jt}^{+}/b_{j})).$
Then the difference between $S_{jl,b}^{+}$ and its expectation can be decomposed into two parts
$\mathbb{E}_{0}S_{jl,b}^{+}=\sum_{t=1}^{T}M_{jt}+\sum_{t=1}^{T}D_{jt}=\mathrm{I}_{j1}+\mathrm{I}_{j2}.$
Note that for any $j,$ $M_{jt}$, $t\geq 1,$ are martingale differences with respect to $\mathcal{F}_{t}$. Moreover, we have $M_{jt}\leq|K|_{\infty}b_{j}^{l}$ and
[…]
uniformly over $j$. Thus by Freedman’s inequality (Freedman, (1975)),
$\mathbb{P}\big{(}|\mathrm{I}_{j1}|\geq z\big{)}\lesssim\mathrm{exp}\big{\{}-z^{2}/(b_{j}^{l}z+Tb_{j}^{2l+1})\big{\}},$
where the constants in $\lesssim$ or $O(\cdot)$ here and all the followings of this proof are independent of $N,T,j,b_{j}.$ That is with probability greater than $1-(TN)^{-p}$,
$\max_{1\leq j\leq N}|\mathrm{I}_{j1}/(Tb_{j}^{l+1})|\lesssim\sqrt{\mathrm{log}(TN)/(T\underline{b})}.$ (SA-19)
For part $\mathrm{I}_{j2}$, by Taylor’s expansion, we have
[…]
uniformly over $j$. By Theorem 6.2 in Zhang and Wu, (2017), for $z\gtrsim\sqrt{T\mathrm{log}(N)},$
$\mathbb{P}\bigg{(}\max_{1\leq j\leq N}\Big{|}\sum_{t=1}^{T}\mathbb{E}_{0}g_{j,t}(0|\mathcal{F}_{t-1})\Big{|}\geq z\bigg{)}\lesssim T^{(p/2-\alpha p)\vee 1}\mathrm{log}(N)^{p/2}/z^{p}+e^{-z^{2}/T}.$
Hence with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$ we have
$\max_{1\leq j\leq N}|\mathrm{I}_{j2}/(Tb_{j}^{l+1})|\lesssim\bar{b}^{l+1}\sqrt{T\mathrm{log}(TN)}+\bar{b}.$ (SA-20)
Combining (SA-19) and (SA-20), with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, we have
$\max_{1\leq j\leq N}|\mathbb{E}_{0}S_{jl,b}^{+}|/(Tb_{j}^{l+1})\lesssim\sqrt{\mathrm{log}(TN)/(T\underline{b})}+\bar{b}.$
Lastly, we need to approximate the expectation part. Note that
[…]
uniformly over $j$. Hence (SA-6) holds. Inequality (SA-7) directly comes from (SA-6).
[…]
Proof of Lemma B.7. We shall only work on the set $\mathcal{A}_{T}$. Let $w_{jt,b}$ be as defined in (SA-1). Then $I_{\epsilon}$ can be rewritten as
$I_{\epsilon}\mathbf{1}_{\mathcal{A}_{T}}=\max_{1\leq j\leq N}\frac{(Tb_{j})^{1/2}}{v_{j}}\Big{|}\sum_{t=1}^{T}w_{jt,b}\big{[}\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}+\varepsilon_{j}(X_{jt},U_{jt})\big{]}+\gamma_{j}\Big{|}.$ (SA-21)
Recall the definition of $\epsilon_{jt,m}$ in (SA-3) and $r_{j}(X_{jt},U_{jt})$ in (SA-1), then
$|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\leq$ $\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})(\epsilon_{jt}-\epsilon_{jt,m})\Big{|}$
$=$ $\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})\sum_{k\leq t-m}A_{t-k,j,\cdot}^{\top}\eta_{k}\Big{|}$
$=$ $\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{k\leq T-m}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\eta_{k}\Big{|},$ (SA-22)
where $A_{t,j,\cdot}$ is the $j$th row of matrix $A_{t}$. Let
$\xi_{jk}=(Tb_{j})^{1/2}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\eta_{k}.$
Then conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$, $\xi_{jk}$ are independent for different $k,$ and
$|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\leq\max_{1\leq j\leq N}\Big{|}\sum_{k\leq T-m}\xi_{jk}\Big{|}.$
Let
$\sigma^{2}$ $=\max_{1\leq j\leq N}\sum_{k\leq T-m}\mathbb{E}(\xi_{jk}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$
$\lesssim\max_{1\leq j\leq N}Tb_{j}\sum_{k\leq T-m}\Big{(}\sum_{t=1\vee(k+m)}^{T}|w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}|_{2}\Big{)}^{2}.$ (SA-23)
Since kernel function $K(\cdot)$ is bounded, by Lemma B.2, $\max_{t,j}(Tb_{j})|w_{jt,b}|=O(1)$ and $\max_{j}\sum_{t=1}^{T}|w_{jt,b}|=O(1),$ where the constants in $O(\cdot)$ and $\lesssim$ here and throughout this proof are independent of $N,T,b,j$. Hence we have
$\max_{j,k}(Tb_{j})\sum_{t=1\vee(k+m)}^{T}|w_{jt,b}A_{t-k,j,\cdot}|_{2}/v_{j}\lesssim m^{-\beta},$ (SA-24)
and
$\max_{j}\sum_{k\leq T-m}\sum_{t=1\vee(k+m)}^{T}|w_{jt,b}||A_{t-k,j,\cdot}|_{2}/v_{j}=\max_{j}\sum_{t=1}^{T}\sum_{k\leq t-m}|w_{jt,b}||A_{t-k,j,\cdot}|_{2}/v_{j}\lesssim m^{-\beta}.$ (SA-25)
It follows that
$\sigma^{2}\lesssim m^{-2\beta}.$ (SA-26)
By Lemma B.3, for any $z>0,$
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\geq z\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$
$\lesssim$ $\sum_{k\leq T-m}\sum_{1\leq j\leq N}\mathbb{P}\Big{(}|\xi_{jk}|\geq u|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}+2N\mathrm{exp}\Big{(}-\frac{z^{2}}{zu+m^{-2\beta}}\Big{)},$ (SA-27)
By Markov’s inequality, we have
$\sum_{k\leq T-m}\mathbb{P}\Big{(}|\xi_{jk}|\geq u\Big{)}$ $\leq\sum_{k\leq T-m}u^{-q}\mathbb{E}\big{(}\mathbb{E}(|\xi_{jk}|^{q}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\big{)}$
$\lesssim u^{-q}\sum_{k\leq T-m}(Tb_{j})^{q/2}\mathbb{E}\Big{|}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}^{\top}\Big{|}_{2}^{q}$
$\lesssim(T\underline{b})^{1-q/2}m^{-q\beta}u^{-q}.$ (SA-28)
[…]
Proof of Lemma B.8. By (SA-22), $|I_{\epsilon}-I_{\epsilon,m}|$ can be bounded by
$|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\leq\max_{1\leq j\leq N}\Big{|}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\Big{|},$
where $c_{jk}=(Tb_{j})^{1/2}\sum_{t=1\vee(k+m)}^{T}w_{jt,b}r_{j}(X_{jt},U_{jt})A_{t-k,j,\cdot}.$ Then by (SA-24) and (SA-25), we have
$\max_{j,k}|c_{jk}|_{2}\lesssim m^{-\beta}(\underline{b}T)^{-1/2}\quad\text{and}\quad\max_{j}\sum_{k}|c_{jk}|_{2}^{2}\lesssim m^{-2\beta}.$ (SA-29)
Conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$, $c_{jk}^{\top}\eta_{k}$ are independent for different $k$, therefore for $\lambda|c_{jk}|_{2}\leq\lambda_{0}$, where $\lambda_{0}$ is defined in Assumption 3.1. By Markov’s inequality
$\mathbb{P}\Big{(}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\geq z\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}\leq e^{-\lambda z}\prod_{k\leq T-m}\mathbb{E}\big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}.$ (SA-30)
Since $\mathbb{E}(\eta_{k}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})=0$, we have
$\mathbb{E}\big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$ $=1+\mathbb{E}\Big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}-1-\lambda c_{jk}^{\top}\eta_{k}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}$
$\leq 1+\mathbb{E}\bigg{(}\frac{e^{|\lambda c_{jk}^{\top}\eta_{k}|}-1-|\lambda c_{jk}^{\top}\eta_{k}|}{\lambda^{2}|c_{jk}|_{2}^{2}}\Big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\bigg{)}\lambda^{2}|c_{jk}|_{2}^{2},$
in view of $e^{x}-x\leq e^{|x|}-|x|$ for any $x.$ Note that for any fixed $x>0,$ $(e^{tx}-tx-1)/t^{2}$ is increasing on $t\in(0,\infty)$. Therefore
$\mathbb{E}\big{(}e^{\lambda c_{jk}^{\top}\eta_{k}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$ $\leq 1+\mathbb{E}\bigg{(}\frac{e^{\lambda_{0}|c_{jk}^{\top}\eta_{k}|/|c_{jk}|_{2}}-1-\lambda_{0}|c_{jk}^{\top}\eta_{k}|/|c_{jk}|_{2}}{\lambda_{0}^{2}}\Big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\bigg{)}\lambda^{2}|c_{jk}|_{2}^{2}$
$\leq 1+a_{0}^{\prime}\lambda^{2}|c_{jk}|_{2}^{2}$
$\leq e^{a_{0}^{\prime}\lambda^{2}|c_{jk}|_{2}^{2}},$
where the last inequality is due to $1+x\leq e^{x}$ and $a_{0}^{\prime}=a_{0}/\lambda_{0}^{2}$ with $a_{0}$ defined in Assumption 3.1. Combining the above results with (SA-30), we have
$\mathbb{P}\Big{(}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\geq z\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}\leq e^{-\lambda z}\mathbb{E}\Big{(}e^{a_{0}^{\prime}\lambda^{2}\sum_{k\leq T-m}|c_{jk}|_{2}^{2}}\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}\leq e^{-\lambda z}e^{c\lambda^{2}m^{-2\beta}},$
where the last inequality is due to (SA-29) and $c>0$ is some constant. The same argument can be applied for $\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}<-z,$ then we have
$\mathbb{P}\big{(}|I_{\epsilon}-I_{\epsilon,m}|\mathbf{1}_{\mathcal{A}_{T}}\geq z\big{)}\leq 2Ne^{-\lambda z+c\lambda^{2}m^{-2\beta}}\lesssim N\mathrm{exp}\Big{\{}-c^{\prime}\frac{z^{2}}{m^{-\beta}(\underline{b}T)^{-1/2}z+m^{-2\beta}}\Big{\}},$
[…]
C.3 Proofs of the Technical Lemmas in Section B.2
Proof of Lemma B.9. $(i)$ Recall the definition of $a_{jt,b}$ in (SA-14). We have
$\sum_{t=1}^{T}a_{jt,b}(\epsilon_{jt}-\epsilon_{jt,m})=\sum_{k=-\infty}^{T-m}\sum_{t=1\vee(k+m)}^{T}a_{jt,b}A_{t-k,j,\cdot}^{\top}\eta_{k}.$
Denote
$\xi_{jk}=(Tb_{j})^{1/2}\sum_{t=1\vee(k+m)}^{T}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}^{\top}\eta_{k}/\tilde{v}_{j}.$
Then $\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}$ can be rewritten into
$|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\leq\max_{1\leq j\leq N}\Big{|}\sum_{k\leq T-m}\xi_{jk}\Big{|}.$
Given $\mathcal{F}_{T},\tilde{\mathcal{F}}_{T},$ $\xi_{jk}$ are independent for different $k.$ By the same argument as used to obtain (SA-27), we have
$\mathbb{P}\big{(}|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\geq z|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$
$\leq$ $\sum_{k\leq T-m}\mathbb{P}\Big{(}\max_{1\leq j\leq N}|\xi_{jk}|\geq u\big{|}\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\Big{)}+2Ne^{-z^{2}/(2zu+2v)},$ (SA-31)
where $v$ is the upper bound for
$\sigma^{2}=$ $\max_{1\leq j\leq N}\sum_{k\leq T-m}\mathbb{E}\big{(}\xi_{jk}^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}$
$=$ $\max_{1\leq j\leq N}\sum_{k\leq T-m}Tb_{j}\Big{|}\sum_{t=1\vee(k+m)}^{T}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}\Big{|}_{2}^{2}/\tilde{v}_{j}^{2}$
$\leq$ $\max_{1\leq j\leq N}Tb_{j}\sum_{k\leq T-m}\Big{(}\sum_{t=1\vee(k+m)}^{T}\Big{(}|a_{jt,b}||A_{t-k,j,\cdot}|_{2}+\sum_{l=1}^{N}|a_{lt,b}||A_{t-k,l,\cdot}|_{2}/N\Big{)}\Big{)}^{2}/\tilde{v}_{j}^{2}.$
As in the proof of (SA-26), we have
$\sigma^{2}\lesssim(\bar{b}/\underline{b})m^{-2\beta}\lesssim m^{-2\beta}.$ (SA-32)
Since $\sum_{k\geq m}|A_{k,j,\cdot}|_{2}\lesssim m^{-\beta},$ by the same argument as used to derive (SA-28), we have
$\sum_{k\leq T-m}\sum_{1\leq j\leq N}\mathbb{P}\Big{(}|\xi_{jk}|\geq u\Big{)}\lesssim N(T\underline{b})^{-q/2+1}m^{-q\beta}u^{-q}.$ (SA-33)
Take $u=z/\mathrm{log}(NT)$. Substituting (SA-32) and (SA-33) into (SA-31), we complete the proof of part $(i)$.
$(ii)$ Note that $|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|$ can be bounded by
$|\mathcal{Q}_{\epsilon}-\mathcal{Q}_{\epsilon,m}|\leq\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{k\leq T-m}c_{jk}^{\top}\eta_{k}\Big{|}/\tilde{v}_{j},$
where
$c_{jk}=\sum_{t=1\vee(k+m)}^{T}\Big{(}a_{jt,b}A_{t-k,j,i}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,i}/N\Big{)}.$
[…]
Proof of Lemma B.10 $(i)$ Recall the definition of $a_{jt,b}$ in (SA-14), then $\mathcal{Q}_{\epsilon,m}$ can be rewritten as
$\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}$ $=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}\Big{(}a_{jt,b}\epsilon_{jt,m}-\sum_{l=1}^{N}a_{lt,b}\epsilon_{lt,m}/N+w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})$
$\quad-\sum_{l=1}^{N}w_{lt,b}\varepsilon_{l}(X_{lt},U_{lt})/N\Big{)}+(\gamma_{j}-\bar{\gamma})\Big{|}/\tilde{v}_{j}$
$=\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{k=2-m}^{T}c_{jk}^{\top}\eta_{k}+\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\overline{{\gamma}})\Big{|}/\tilde{v}_{j},$ (SA-34)
where
$c_{jk}=\sum_{t=1\vee k}^{T\wedge(k+m-1)}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}$
and
$\xi_{jt}=w_{jt,b}\varepsilon_{j}(X_{jt},U_{jt})-\sum_{l=1}^{N}w_{lt,b}\varepsilon_{l}(X_{lt},U_{lt})/N.$
Decompose $a_{jt,b}$ into two parts,
$M_{jt}=a_{jt,b}-\mathbb{E}(a_{jt,b}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})\quad\text{and}\quad D_{jt}=\mathbb{E}(a_{jt,b}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1}).$ (SA-35)
Let $\tilde{\sigma}_{j}(X_{jt})=\mathbb{E}(\sigma_{j}(X_{jt},U_{jt})|X_{jt}).$ Since density function $g_{j}$ is bounded,
$|D_{jt}|$ $\lesssim\big{|}\mathbb{E}(\tilde{\sigma}_{j}(X_{jt})K(X_{jt}/b_{j})/(b_{j}T)|\mathcal{F}_{t-1})\big{|}$
$\lesssim\Big{|}\int T^{-1}K(x)\tilde{\sigma}_{j}(b_{j}x)g_{j}(b_{j}x|\mathcal{F}_{t-1})dx\Big{|}\lesssim T^{-1},$
where the constants in $\lesssim$ and $O(\cdot)$ here and all the followings of this proof are independent of $j,t,N,T.$ In the following we shall firstly show that with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$,
$\min_{1\leq j\leq N}(Tb_{j})\sum_{k=2-m}^{T}\mathbb{E}\big{(}(c_{jk}^{\top}\eta_{k})^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}\geq c,$ (SA-36)
where $c>0$ is some constant. Note that
$\sum_{k=2-m}^{T}\mathbb{E}\big{(}(c_{jk}^{\top}\eta_{k})^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}\big{)}\geq\sum_{k=1}^{T-m}\Big{|}\sum_{t=k}^{k+m-1}\Big{(}a_{jt,b}A_{t-k,j,\cdot}-\sum_{l=1}^{N}a_{lt,b}A_{t-k,l,\cdot}/N\Big{)}\Big{|}_{2}^{2}.$ (SA-37)
Denote $M_{jt}$ and $D_{jt}$ as in (SA-35), and let
$Y_{jk}=\Big{|}\sum_{t=k}^{k+m-1}\Big{(}M_{jt}A_{t-k,j,\cdot}-\sum_{l=1}^{N}M_{lt}A_{t-k,l,\cdot}/N\Big{)}\Big{|}_{2}^{2}.$
Since $|D_{jt}|\lesssim 1/T,$ similarly to (SA-12), (SA-37) can be further lower bounded by
$\sum_{k=2-m}^{T}\mathbb{E}((c_{jk}^{\top}\eta_{k})^{2}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\geq\sum_{k=1}^{T-m}\Big{(}Y_{jk}^{1/2}-O(1/T)\Big{)}^{2}\geq\sum_{k=1}^{T-m}Y_{jk}-O\Big{(}\Big{(}\sum_{k=1}^{T-m}Y_{jk}/T\Big{)}^{1/2}\Big{)}.$
By Assumption 3.3, for $j\neq l$, we have $\mathbb{E}(M_{jt}M_{lt})=O(T^{-2})$ and thus
$\sum_{k=1}^{T-m}\mathbb{E}(Y_{jk})$
$=$ $\sum_{k=1}^{T-m}\sum_{t=k}^{k+m-1}\mathbb{E}\Big{|}M_{jt}A_{t-k,j,\cdot}-\sum_{l=1}^{N}M_{lt}A_{t-k,l,\cdot}/N\Big{|}_{2}^{2}$
$=$ $\sum_{k=1}^{T-m}\sum_{t=k}^{k+m-1}\Big{(}\mathbb{E}(M_{jt}^{2})(1-1/N)^{2}|A_{t-k,j,\cdot}|_{2}^{2}+\sum_{\{l:l\neq j\}}\mathbb{E}(M_{lt}^{2})|A_{t-k,l,\cdot}|_{2}^{2}/N^{2}\Big{)}+O(1/T)$
$\gtrsim$ $(\underline{b}T)^{-1}.$ (SA-38)
We shall bound the difference between $Y_{jk}$ and $\mathbb{E}Y_{jk}$. Let $L_{jtk}=M_{jt}A_{t-k,j,\cdot}-\sum_{l=1}^{N}M_{lt}A_{t-k,l,\cdot}/N$ and denote
$\tilde{\xi}_{jt}=\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\Big{(}|L_{jtk}|_{2}^{2}+2L_{jtk}^{\top}\sum_{k\leq s<t}L_{jsk}\Big{)}.$
Then the difference can be decomposed into two parts,
$\sum_{k=1}^{T-m}\mathbb{E}_{0}Y_{jk}=\sum_{t=1}^{T}(\tilde{\xi}_{jt}-\mathbb{E}(\tilde{\xi}_{jt}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1}))+\sum_{t=1}^{T}(\mathbb{E}(\tilde{\xi}_{jt}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})-\mathbb{E}\tilde{\xi}_{jt})=\mathrm{I}_{j1}+\mathrm{I}_{j2},$
Similarly to the argument in (SA-13), we have $|\tilde{\xi}_{jt}|=O((b_{j}T)^{-2})$, $\sum_{t=1}^{T}\mathbb{E}(\tilde{\xi}_{jt}^{2}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})=O((b_{j}T)^{-3})$ and thus
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}|\mathrm{I}_{j1}|\geq z/(b_{j}T)^{2}\Big{)}\lesssim N\mathrm{exp}\Big{\{}-\frac{z^{2}}{z+b_{j}T}\Big{\}}.$ (SA-39)
Hence with probability greater than $1-T^{-p}$, we have $\max_{1\leq j\leq N}(b_{j}T)|\mathrm{I}_{j1}|=\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$
For $\mathrm{I}_{j2}$ part, similar to (SA-15), we have
$\mathbb{E}(M_{jt}^{2}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})=\int b_{j}(b_{j}T)^{-2}K^{2}(x)\tilde{\sigma}_{j}^{2}(c_{j0})g_{j}(c_{j0}|\mathcal{F}_{t-1})\mathrm{d}x(1+O(b_{j})),$
and that
$\mathbb{E}(\tilde{\xi}_{jt}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1})=\sum_{k=1\vee(t-m+1)}^{(T-m)\wedge t}\mathbb{E}(|L_{jtk}|_{2}^{2}|\mathcal{F}_{t-1},\tilde{\mathcal{F}}_{t-1}).$
Hence by Theorem 6.2 in Zhang and Wu, (2017),
$\mathbb{P}\Big{(}\max_{1\leq j\leq N}(b_{j}T^{2})|\mathrm{I}_{j2}|\geq z+O(\bar{b})\Big{)}\lesssim z^{-p}T^{(p/2-\alpha p)\vee 1}\mathrm{log}(N)^{p/2}+e^{-z^{2}/T}.$
Hence combining $\mathrm{I}_{j1}$ and $\mathrm{I}_{j2},$ with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$
$b_{j}T\Big{|}\sum_{k=1}^{T-m}\mathbb{E}_{0}Y_{jk}\Big{|}\lesssim\sqrt{\mathrm{log}(NT)/(\underline{b}T)}+\bar{b}.$ (SA-40)
This, in conjunction with (SA-38), completes the proof of (SA-36).
We shall then show that with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)})$, for $B_{T}=c\underline{b}^{-1/2},$ some constant $c>0$ large enough, $l=1,2,$ we have
$\max_{1\leq j\leq N}b_{j}^{(2+l)/2}T^{1+l}\sum_{k=2-m}^{T}\mathbb{E}(|c_{jk}^{\top}\eta_{k}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\leq B_{T}^{l}.$ (SA-41)
Since $\eta_{ik}$ are independent for different $i,k$,
$\sum_{k=2-m}^{T}\mathbb{E}(|c_{jk}^{\top}\eta_{k}|^{2+l}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})\lesssim\sum_{k=2-m}^{T}|c_{jk}|_{2}^{2+l}\lesssim(b_{j}T)^{-l}\sum_{k=2-m}^{T}Y_{jk}+T^{-(1+l)}.$ (SA-42)
By a similar argument as used to derive (SA-38), we have $b_{j}T\sum_{k=2-m}^{T}\mathbb{E}Y_{jk}=O(1).$ Hence together with (SA-40), with probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$
$b_{j}T\sum_{k=2-m}^{T}Y_{jk}=O(1).$ (SA-43)
Combining (SA-42) and (SA-43), we complete the proof of (SA-41). Let $\tilde{Z}_{j,1}$, $1\leq j\leq N$, be Gaussian distributed random variables with the same conditional covariance structure as $(Tb_{j})^{1/2}\sum_{k=2-m}^{T}c_{jk}^{\top}\eta_{k}$ given $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$. For any $j$ conditional on $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T},$ $c_{jk}^{\top}\eta_{k}$ are independent for different $k$. Thus by Proposition 2.1 in Chernozhukov et al., 2017b ,
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\Big{|}\tilde{Z}_{j,1}+(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\bar{\gamma})\Big{)}\Big{|}/\tilde{v}_{j}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT).$
Define
$\tilde{v}_{j,1}^{2}=v_{j}^{2}\big{\{}(1-1/N)^{2}v_{j,1}^{2}+\sum_{l=1}v_{l,1}^{2}/N^{2}\big{\}},$
where $v_{j,1}$ defined in (SA-4). Note that
$\mathbb{E}(\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}))=\tilde{v}_{j,1}^{2}+O(b_{j}+m^{-\beta}).$
And for any $j_{1}\neq j_{2}$ we have
$\mathbb{E[}\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j_{2},1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})]=O(\bar{b}).$
Let $Z_{j,1}$ be i.i.d. from $N(0,\tilde{v}_{j,1}^{2})$ and is independent of $\mathcal{F}_{T}$ and ~$\tilde{\mathcal{F}}_{T}$. With probability greater than $1-O(T^{-(\alpha p)\wedge(p/2-1)}),$ the maximum difference of $\mathrm{Var}(\tilde{Z}_{j,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ and $\mathrm{Cov}(\tilde{Z}_{j_{1},1},\tilde{Z}_{j2,1}|\mathcal{F}_{T},\tilde{\mathcal{F}}_{T})$ from their expectation are bounded by $\bar{b}+\sqrt{\mathrm{log}(NT)/(\underline{b}T)}.$ Hence by Lemma B.6, we further have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\Big{|}Z_{j,1}+(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\bar{\gamma})\Big{)}\Big{|}/\tilde{v}_{j}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$
Random variable $Z_{j,1}=(Tb_{j})^{1/2}\sum_{t=1}^{T}\tilde{\xi}_{jt,1},$ where $\tilde{\xi}_{jt,1}$ are i.i.d. from $N(0,\tilde{v}_{j,1}^{2}/(T^{2}b_{j}))$ and independent from $\mathcal{F}_{T},\tilde{\mathcal{F}}_{T}$. Then
$Z_{j,1}+(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}\xi_{jt}+(\gamma_{j}-\bar{\gamma})\Big{)}=(Tb_{j})^{1/2}\Big{(}\sum_{t=1}^{T}(\tilde{\xi}_{jt,1}+\xi_{jt})+(\gamma_{j}-\bar{\gamma})\Big{)}.$
Define $Z_{j}^{\diamond}$s as independent Gaussian random variable with mean 00 and variance $\tilde{v}_{j}^{2}$. Given $\mathcal{F}_{T}$, $\tilde{\xi}_{jt,1}+\xi_{jt,2}$ are independent for different $t.$ Similar to the argument in the proof of (SA-8), we have
$\sup_{u\in\mathbb{R}}\bigg{|}\mathbb{P}\big{(}\mathcal{Q}_{\epsilon,m}\mathbf{1}_{\mathcal{A}_{T}}\leq u\big{)}-\mathbb{P}\Big{(}\max_{1\leq j\leq N}\big{|}Z_{j}^{\diamond}+(Tb_{j})^{1/2}(\gamma_{j}-\bar{\gamma})\big{|}/\tilde{v}_{j}\leq u\Big{)}\bigg{|}$
$\lesssim$ $(\underline{b}T)^{-1/6}\mathrm{log}^{7/6}(NT)+(T^{2/q}/(\underline{b}T))^{1/3}\mathrm{log}(NT)+(\bar{b}+m^{-\beta})^{1/3}\mathrm{log}^{2/3}(N).$ (SA-44)
Proof of $(ii)$. Since $\max_{j,k}|c_{jk}|_{2}=O(1/(\underline{b}T)),$ for $B_{T}=c\underline{b}^{-1/2}$ some constant $c>0$ large enough, we have $Tb_{j}^{1/2}|c_{jk}|_{2}/B_{T}\leq\lambda_{0}\mathrm{log}_{(a_{0}\vee 2)}(2)$ and thus
$\mathbb{E}\big{(}\mathrm{exp}\{Tb_{j}^{1/2}|c_{jk}^{\top}\eta_{k}|/B_{T}\}\big{)}\leq 2.$
[…]
Appendix D Additional Algorithms
We shall list the two additional algorithms, i.e. Algorithm SA-2 and Algorithm SA-3. The derivative case follows a similar approach to the mean case. Thus, we include the algorithm for the derivative case below, presented in a bracelet format.
Algorithm SA-2 Testing for homogeneity, i.e., $\gamma_{1}=\ldots=\gamma_{N}$ ( $\gamma_{1}^{[1]}=\ldots=\gamma_{N}^{[1]}$).
1:for $j=1,\ldots,N$ do
2:     Estimate $\gamma_{j}$ for each group/individual $j$ using equation (2.2).
3:     Estimate the variance $v_{j}^{2}$ ($v_{j}^{[1]2}$) of $(Tb_{j})^{1/2}\hat{\gamma}_{j}$($(Tb_{j}^{3})^{1/2}\hat{\gamma}_{j}^{[1]}$) by $\hat{v}_{j}^{2}=Tb_{j}^{3}\sum_{t=1}^{T}\left(w_{jt,b}^{+}-w_{jt,b}^{-}\right)^{2}\hat{\sigma}_{e,j}^{2}$ ( $\hat{v}_{j}^{[1]2}=Tb_{j}^{3}\sum_{t=1}^{T}\left(w_{jt,b}^{+[1]}-w_{jt,b}^{-[1]}\right)^{2}\hat{\sigma}_{e,j}^{2}$), where $\hat{\sigma}_{e,j}^{2}$ is a consistent estimator for the error variance using the observations close to the threshold $c_{0j}$.
4:end for
5:Calculate the feasible test statistic $\hat{\mathcal{Q}}:=\mbox{max}_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}-\frac{1}{N}\sum_{i=1}^{N}\hat{\gamma}_{i}|/\hat{\tilde{v}}_{j}$ ($\hat{\mathcal{Q}}^{[1]}:=\mbox{max}_{1\leq j\leq N}(Tb_{j}^{3})^{1/2}|\hat{\gamma}_{j}^{[1]}-\frac{1}{N}\sum_{i=1}^{N}\hat{\gamma}_{i}^{[1]}|/\hat{\tilde{v}}_{j}^{[1]}$), where $\hat{\tilde{v}}_{j}=\left(\hat{v}_{j}^{2}\left(1-\frac{1}{N}\right)^{2}+\frac{1}{N^{2}}\sum_{i\neq j}^{N}\hat{v}_{i}^{2}\right)^{1/2}$ ($\hat{\tilde{v}}_{j}^{[1]}=\left(\hat{v}_{j}^{[1]2}\left(1-\frac{1}{N}\right)^{2}+\frac{1}{N^{2}}\sum_{i\neq j}^{N}\hat{v}_{i}^{[1]2}\right)^{1/2}$).
6:Reject the null hypothesis if $\hat{\mathcal{Q}}(\hat{\mathcal{Q}}^{[1]})>q_{\alpha}$, where $q_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}|Z_{j}|$ with $Z_{j}$’s being independent standard normal variables.
Similarly for the algorithm below in derivative, we shall replace $\hat{\gamma}_{j}(c_{i})$ by $\hat{\gamma}_{j}^{[1]}(c_{i})$, $\hat{v}_{j}^{2}(c_{i})$ by $\hat{v}_{j}^{[1]2}(c_{i})$ and $Tb_{j}$ by $Tb_{j}^{3}$. The newly estimated threshold estimator can be denoted as $\hat{c}_{j}^{[1]}$.
Algorithm SA-3 Testing for existence of jumps: i.e., $\gamma_{j}=0$ for all $j$ with unknown threshold.
1:for $j=1,\ldots,N$ do
2:     For a given grid of $[c_{1},c_{2},\cdots,c_{K}]$, estimate $\gamma_{j}(c_{i})$, for each threshold location $c_{i}$, ($i=1,\cdots,K$), using equation (2.2).
3:     Estimate the variance $v_{j}^{2}(c_{i})$ of $(Tb_{j})^{1/2}\hat{\gamma}_{j}(c_{i})$ by $\hat{v}_{j}^{2}(c_{i})=Tb_{j}\sum_{t=1}^{T}\left(w_{jt,b}^{+}(c_{i})-w_{jt,b}^{-}(c_{i})\right)^{2}\hat{\sigma}_{e,j}^{2}(c_{i})$ for each $i$, where $\hat{\sigma}_{e,j}^{2}(c_{i})$ is estimated by (SA-4).
4:end for
5:Calculate the feasible test statistic ^$\hat{\mathcal{I}}^{C}:=\max_{1\leq i\leq K}\max_{1\leq j\leq N}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$.
6:Reject the null hypothesis if ^$\hat{\mathcal{I}}^{C}>q^{C}_{\alpha}$, where $q^{C}_{\alpha}$ is the $(1-\alpha)$ quantile of the Gaussian random variable $\max_{1\leq j\leq N}\max_{1\leq l\leq K}|Z_{jl}|$ with $Z_{jl}$’s being independent standard normal variables.
7:Return the $\hat{c}_{j}$ which corresponds to the maximum absolute value of $\max_{1\leq i\leq K}(Tb_{j})^{1/2}|\hat{\gamma}_{j}(c_{i})/\hat{v}_{j}(c_{i})|$.
Appendix E Additional Simulation Settings and Results
E.1 Data Generating Processes
Throughout the simulation study, we consider the following setup:
$Y_{jt}=\cos(X_{jt})+\sin(U_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq 0\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt},$ (SA-1)
where the latent variable is generated by $U_{jt}=\tilde{U}_{jt}X_{jt}$ and $\tilde{U}_{jt}\sim i.i.d.\ U[-1,1]$. To study the performance of our uniform testing procedure, we consider three different data generating processes (DGPs) with different specifications of the covariate variable $X_{jt}$, the error term $\epsilon_{jt}$ and the volatility function $\sigma_{j}(\cdot,\cdot)$. For the first one, we consider the independent setting. For the latter two, we include cross-sectional and temporal dependences. We consider either homoskedastic settings (DGP2) in which $\sigma_{j}(\cdot,\cdot)=1$ and heteroskedastic settings (DGP1, DGP3) in which we set
$\sigma_{j}(x,u)=1+\left(\frac{3}{8}-\frac{\left|x\right|}{4}\right)\left(\frac{3}{2}\right)^{2u}.$ (SA-2)
  1. DGP1: i.i.d. setting with heteroskedasticity. $X_{jt}\sim i.i.d.\ U[-1,1]$, $\epsilon_{jt}\sim i.i.d.\ N(0,1)$.
  2. DGP2: Factor structure and homoskedasticity. Both the running variable and error term have a factor structure, $\epsilon_{jt}=\lambda_{\epsilon,j}f_{\epsilon,t}+u_{\epsilon,jt}$ and $X_{jt}=1/4\left(\lambda_{x,j}f_{x,t}+u_{x,jt}\right)$, where $f_{\epsilon,t}$, $u_{\epsilon,jt}$, $f_{x,t}$ and $u_{x,jt}$ are each generated from an MA($\infty$) process with algebraic decay parameter, $\beta=1.5$. The factor loadings, $\lambda_{\epsilon,j}$ and $\lambda_{x,j}$, are i.i.d. draws from a standard normal distribution.
  3. DGP3: Factor structure and heteroskedasticity. The setting is the same as in DGP2 except that the error term is heteroskedastic with the conditional standard deviation described in (SA-2).
Additionally, we examine three additional cases, referred to as DGP4–6 in the Supplementary Material E, which account for varying degrees of cross-sectional dependence and a possible time effect. All DGPs, except for the first one, exhibit strong cross-sectional and weak temporal correlation in the covariate variable and the error term. We are interested in testing the existence of the jump effect, namely, $H_{0}^{\left(1\right)}:\gamma_{j}=0$ for all $j\in[N]$. For each DGP, we study both the size and local power performance of our testing procedure. When studying the size performance, we set all $\gamma_{j}=0$. When examining the local power, according to Theorem A.3 in the Supplement and to the selection of the bandwidth, we set a fraction of the threshold effects to be $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$, and the remaining $\gamma_{j}$’s to be zero. We consider two different choices for the cross-sectional dimension: $N=10,$ 100. When $N=10$, the above fraction is set to $20\%$, whereas in the high-dimensional setting with $N=100,$ the fraction is decreased to $10\%$.
E.2 Additional Simulation Results
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP1 10 200 0.113 0.059 0.008 0.395 0.293 0.142
10 400 0.133 0.060 0.014 0.428 0.316 0.153
10 800 0.115 0.054 0.005 0.466 0.333 0.172
100 200 0.099 0.051 0.014 0.884 0.797 0.596
100 400 0.115 0.054 0.011 0.926 0.874 0.699
100 800 0.099 0.050 0.008 0.948 0.922 0.786
DGP2 10 200 0.105 0.054 0.013 0.539 0.443 0.269
10 400 0.122 0.061 0.012 0.583 0.489 0.316
10 800 0.115 0.062 0.010 0.647 0.548 0.377
100 200 0.101 0.042 0.007 0.981 0.962 0.892
100 400 0.107 0.043 0.010 0.982 0.976 0.944
100 800 0.096 0.053 0.010 0.996 0.991 0.964
DGP3 10 200 0.123 0.055 0.007 0.371 0.275 0.105
10 400 0.121 0.058 0.017 0.393 0.299 0.152
10 800 0.122 0.066 0.014 0.465 0.380 0.220
100 200 0.115 0.056 0.019 0.848 0.780 0.594
100 400 0.111 0.061 0.011 0.909 0.849 0.689
100 800 0.116 0.064 0.012 0.952 0.902 0.775
Table A1: Empirical size and power results for the uniform test for the existence of threshold effects for DGP1–3.
The finite sample size and power properties for the test of homogeneous effects are also examined for DGPs 1-3, as shown in Table A2. The findings are consistent with those of the test for the existence of jump effects, with the empirical size closely matching the nominal size across all the scenarios. However, for DGP3, the test procedure shows a slight tendency to over-reject the null, potentially due to the presence of heteroskedasticity in the error terms. The local power results align with the earlier findings from the significance test for the existence of jump effects.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP1 10 200 0.114 0.070 0.010 0.315 0.226 0.092
10 400 0.114 0.065 0.014 0.374 0.267 0.118
10 800 0.119 0.056 0.010 0.416 0.286 0.113
100 200 0.104 0.058 0.018 0.845 0.741 0.519
100 400 0.118 0.062 0.014 0.897 0.824 0.603
100 800 0.120 0.062 0.013 0.933 0.882 0.709
DGP2 10 200 0.115 0.067 0.011 0.440 0.329 0.160
10 400 0.111 0.059 0.013 0.483 0.382 0.188
10 800 0.107 0.059 0.010 0.563 0.445 0.272
100 200 0.111 0.052 0.012 0.954 0.922 0.816
100 400 0.104 0.051 0.013 0.984 0.960 0.898
100 800 0.115 0.055 0.009 0.995 0.992 0.951
DGP3 10 200 0.115 0.070 0.012 0.315 0.216 0.081
10 400 0.123 0.074 0.012 0.353 0.242 0.094
10 800 0.113 0.059 0.015 0.386 0.269 0.122
100 200 0.125 0.054 0.011 0.801 0.712 0.498
100 400 0.123 0.060 0.009 0.854 0.787 0.572
100 800 0.128 0.058 0.009 0.923 0.873 0.701
Table A2: Empirical size and power results for the uniform test of heterogeneity for DGP1–3.
Now we consider three additional simulation settings with strong cross-sectional and weak temporal dependence.
  1. DGP4:
    Strong cross-sectional dependence. The setting is similar to that in DGP3, but now the factor term in the running variable and error dominate the idiosyncratic part via a nonzero mean and a relatively larger variance,
    $\epsilon_{jt}$ $=\left(\lambda_{\epsilon,j}+2\right)f_{\epsilon,t}+u_{\epsilon,jt}/8,$
    $X_{jt}$ $=\frac{1}{4}\left((\lambda_{x,j}+2)f_{x,t}+u_{x,jt}/8\right).$
  2. DGP5:
    Semi-strong cross-sectional dependence. The setting is similar to that in DGP4, with some more weight put on the idiosyncratic part of the error term and the running variable,
    $\epsilon_{jt}$ $=\left(\lambda_{\epsilon,j}+2\right)f_{\epsilon,t}+u_{\epsilon,jt}/4$
    $X_{jt}$ $=\frac{1}{4}\left((\lambda_{x,j}+2)f_{x,t}+u_{x,jt}/4\right).$
  3. DGP6:
    Semi-strong cross-sectional dependence and time effect. The setting is similar to that in DGP5, with an additional time effect for the error term,
    $\epsilon_{jt}=\left(\lambda_{\epsilon,j}+2\right)f_{\epsilon,t}+\nu_{t}+u_{\epsilon,jt}/4,$
    where $\nu_{t}\sim i.i.d.\ N(0,0.25)$.
We study the size and power properties of the test for the existence of jump effects for DGP4 – DGP6 in Table A3. These DGPs are the ones with the strong cross-sectional dependence we consider in our simulation study. For DGP4, we observe an empirical size slightly below the nominal size for the high-dimensional setting with $N=100$ which is due to the increase in the cross-sectional dependence in the covariate and the error term. The sizes for DGP5 are almost perfect, while the inclusion of an additional time effect in DGP6 increases the empirical size slightly. This time effect adds more noise to the system, so it is not surprising that the empirical power of DGP6 is lower than that of the other DGPs we consider.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP4 10 200 0.104 0.051 0.012 0.311 0.246 0.157
10 400 0.100 0.045 0.008 0.334 0.273 0.193
10 800 0.098 0.048 0.014 0.380 0.314 0.209
100 200 0.084 0.042 0.006 0.755 0.711 0.625
100 400 0.079 0.039 0.009 0.798 0.757 0.671
100 800 0.090 0.047 0.014 0.828 0.792 0.728
DGP5 10 200 0.108 0.058 0.010 0.320 0.240 0.141
10 400 0.114 0.052 0.018 0.368 0.289 0.199
10 800 0.100 0.051 0.011 0.397 0.312 0.199
100 200 0.107 0.053 0.012 0.753 0.702 0.581
100 400 0.106 0.056 0.009 0.816 0.780 0.685
100 800 0.102 0.058 0.009 0.858 0.807 0.713
DGP6 10 200 0.100 0.056 0.017 0.216 0.137 0.053
10 400 0.119 0.059 0.015 0.259 0.174 0.068
10 800 0.108 0.060 0.009 0.270 0.185 0.076
100 200 0.122 0.071 0.013 0.477 0.391 0.229
100 400 0.110 0.056 0.010 0.564 0.468 0.283
100 800 0.117 0.049 0.007 0.647 0.546 0.376
Table A3: Empirical results on the size and power of the uniform test for the existence of jump effects for DGP4–6.
While the preceding results are concerned with the known threshold case, we now exam the size and power properties of our uniform testing procedure in the case of unknown thresholds. We consider the same model specification as in the known threshold case. Therefore, the true threshold is $c_{0j}=0$ for all $j$. For the practical implementation of our test procedure, we consider the following grids of threshold locations. Due to different ranges in the covariate distribution, we consider the grid $[-0.6,-0.3,0,0.3,0.6]$ for DGP1 and $[-0.3,-0.15,0,0.15,0.3]$ for DGP2 and 3. The results in Table A4 reveal that the test has still good size properties, with the empirical size being slightly smaller than nominal in some settings and slightly larger in others. As expected, the power of the test decreases compared to the case with a known threshold. However, as $T$ increases, the power still approaches one.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP1 10 400 0.091 0.038 0.005 0.184 0.110 0.031
10 800 0.120 0.060 0.008 0.323 0.221 0.098
100 400 0.086 0.039 0.006 0.463 0.339 0.186
100 800 0.099 0.048 0.012 0.883 0.813 0.649
DGP2 10 400 0.118 0.058 0.013 0.444 0.349 0.195
10 800 0.102 0.055 0.008 0.503 0.414 0.263
100 400 0.107 0.045 0.006 0.955 0.941 0.869
100 800 0.117 0.057 0.011 0.986 0.972 0.935
DGP3 10 400 0.100 0.045 0.006 0.316 0.231 0.134
10 800 0.116 0.071 0.017 0.534 0.458 0.355
100 400 0.124 0.058 0.009 0.624 0.535 0.388
100 800 0.124 0.068 0.010 0.877 0.842 0.751
Table A4: Empirical size and power results for the uniform test of the existence of jump effects under unknown threshold locations for DGP1–3.
Apart from the size and power properties of our uniform testing procedure, we also check the estimation accuracy of $\hat{c}_{j}$ ( estimator of the unknown threshold location $c_{0j}$). For this purpose, we set all $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}5B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$. We consider a grid of possible threshold locations, i.e. $[-0.3,-0.29,\ldots,0.3]$. The results in Table A5 reveal that the maximum and averaged error over $j$ can be reduced effectively with increasing $T$.
$N$ $T$ $1/N\sum_{j}|\hat{c}_{j}-c_{0j}|$ $\max_{j}|\hat{c}_{j}-c_{0j}|$
DGP1 10 200 0.0241 0.1399
10 400 0.0140 0.0979
10 800 0.0087 0.0668
10 1600 0.0063 0.0528
50 200 0.0156 0.2260
50 400 0.0069 0.1520
50 800 0.0035 0.1008
50 1600 0.0021 0.0679
DGP2 10 200 0.0085 0.0635
10 400 0.0058 0.0486
10 800 0.0038 0.0332
10 1600 0.0029 0.0272
50 200 0.0048 0.1177
50 400 0.0024 0.0811
50 800 0.0015 0.0564
50 1600 0.0010 0.0381
Table A5: Estimation accuracy of the $\hat{c}_{j}$’s for DGP1–2.
In the following, we compare our uniform testing procedure to that of a pooled parametric threshold model. To be precise, as a comparison we assume the following linear panel threshold model (Hansen, , 2000) (see also Section G.3 for details),
$Y_{jt}=\beta_{0}+\beta_{1}X_{jt}+\mathbf{1}_{\{X_{jt}\geq c_{0}\}}\left(\beta_{0}^{+}+\beta_{1}^{+}X_{jt}\right)+e_{jt}.$
The parameter of interest is $\gamma:=\beta_{0}^{+}+\beta_{1}^{+}c_{0}$ and we test the null hypothesis $H_{0}:\gamma=0$. We consider sup-Wald type statistics discussed by Andrews, (1993). For this comparison, we consider the following linear data generating process.
  1. DGP7:
    Linear model. The smooth part of the conditional mean function is now linear.
    $Y_{jt}=0.2X_{jt}+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\sigma(X_{jt},U_{jt})\epsilon_{jt}.$
    The generation of $X_{jt}$, $U_{jt}$, and $\sigma(\cdot,\cdot)$ is identical to DGP3. A proportion between 00 and $1$ of the coefficients $\gamma_{j}$ are drawn from a standard normal distribution, the remaining coefficients are set to zero.
The comparison results are presented in Table A6. The results confirm that our uniform testing procedure has similar size properties as a parametric test but has much larger power under the alternative. This is true for the sparse case with a proportion of only $10\%$ non-zero signals, as well as for the relatively dense case with a proportion of $20\%$. The reasons for this are signal-dilution on the one hand and signal cancellation on the other hand. Regarding the former, our uniform testing procedure is robust towards settings with sparse signal, in which a pooled test suffers from dilution of signals. And regarding the latter, our uniform testing procedure does not have the problem of signal cancellation in case the signals for different $j$’s have opposite signs and cancel each other out.
Uniform test Pooled parametric
Proportion $N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
0% 50 400 0.085 0.036 0.007 0.126 0.080 0.017
50 800 0.117 0.054 0.013 0.106 0.058 0.011
100 400 0.123 0.055 0.014 0.116 0.067 0.020
100 800 0.106 0.060 0.011 0.113 0.055 0.019
10% 50 400 0.451 0.368 0.249 0.251 0.158 0.066
50 800 0.743 0.682 0.576 0.300 0.215 0.115
100 400 0.608 0.526 0.380 0.216 0.128 0.051
100 800 0.894 0.858 0.753 0.307 0.231 0.120
20% 50 400 0.660 0.583 0.430 0.325 0.225 0.122
50 800 0.896 0.865 0.780 0.392 0.316 0.210
100 400 0.842 0.792 0.661 0.298 0.225 0.110
100 800 0.983 0.967 0.938 0.407 0.327 0.211
Table A6: Empirical size and power results for the uniform test of the existence of jump effects under unknown threshold locations for DGP7. Comparison with a sup-Wald test based on a linear panel threshold model.
We obtain similar results in the comparison with a pooled nonparametric test in Table A7. The results are based on DGP3 and the non-zero coefficients are generated from a standard normal distribution. For the comparison, we pool the observations along the individual dimension. Again, the uniform testing procedure is more powerful than the pooled testing procedure in all settings with existence of some non-zero coefficients.
Uniform test Pooled nonparametric
Proportion $N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
0% 50 400 0.120 0.061 0.012 0.099 0.054 0.009
50 800 0.117 0.061 0.011 0.083 0.044 0.004
100 400 0.122 0.067 0.019 0.094 0.050 0.011
100 800 0.117 0.053 0.018 0.106 0.055 0.014
10% 50 400 0.460 0.376 0.235 0.124 0.072 0.013
50 800 0.693 0.633 0.522 0.144 0.076 0.029
100 400 0.651 0.569 0.414 0.140 0.082 0.021
100 800 0.862 0.823 0.729 0.157 0.095 0.028
20% 50 400 0.653 0.575 0.441 0.148 0.086 0.026
50 800 0.915 0.885 0.790 0.206 0.144 0.061
100 400 0.807 0.737 0.591 0.144 0.097 0.027
100 800 0.983 0.971 0.932 0.215 0.124 0.055
Table A7: Empirical size and power results for the uniform test of the existence of jump effects under unknown threshold locations for DGP3. Comparison with pooled nonparametric test.
Finally, we examine the size and power properties of the test of existence of a threshold effect in the first derivative of the conditional expectations. We restrict our simulations on the known threshold case. Under the null, we set $f_{j}(x,u)=5x$, whereas we consider $f_{j}(x,u)=5|x|$ under the alternative hypothesis. The share of individuals’ DGPs generated under the alternative is $20\%$ for both $N=10$ and $N=100$. For this setting, we oversmooth the bandwidth. The results for DGP1–DGP3 are presented in Table A8. The size properties are reasonably close the nominal size. Similarly, the power approaches one with increasing sample size and dimension.
Size Power
$N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
DGP1 10 400 0.102 0.052 0.007 0.923 0.831 0.527
10 800 0.113 0.046 0.011 0.998 0.995 0.941
100 400 0.120 0.058 0.008 0.999 0.998 0.909
100 800 0.115 0.060 0.010 1.000 1.000 1.000
DGP2 10 400 0.109 0.050 0.014 0.718 0.590 0.327
10 800 0.096 0.048 0.009 0.956 0.915 0.716
100 400 0.100 0.040 0.009 0.978 0.935 0.710
100 800 0.138 0.054 0.012 1.000 1.000 0.997
DGP3 10 400 0.108 0.064 0.014 0.710 0.564 0.303
10 800 0.128 0.067 0.013 0.946 0.903 0.703
100 400 0.130 0.054 0.010 0.988 0.939 0.693
100 800 0.126 0.057 0.016 1.000 1.000 0.999
Table A8: Empirical size and power results for the uniform test of the existence of derivative jump effects under known threshold locations for DGP1–3.
Appendix F Empirical Applications
In this section, we present further estimation results for our two empirical applications. First, we study possible jumps in the news impact curve for constituents of the S&P 500. As a second application, we study the impact of the incumbency effect in U.S. House elections.
F.1 Application I: Jumps in the News Impact Curve
Symbol $c_{0j}$ $100\cdot\hat{\gamma}_{j}(c_{0j})$ $100\cdot(T_{j}b_{j})^{-1/2}\hat{v}_{j}(c_{0j})$ $(T_{j}b_{j})^{1/2}\hat{\gamma}_{j}(c_{0j})/\hat{v}_{j}(c_{0j})$
ALL –0.010 –0.240 0.043 –5.535
AXP –0.010 –0.223 0.048 –4.702
CTAS –0.010 –0.189 0.041 –4.632
ORLY –0.010 –0.211 0.044 –4.750
SJM –0.010 –0.191 0.040 –4.742
NXPI –0.005 –0.529 0.078 –6.793
HON 0.005 0.136 0.028 4.763
T 0.010 0.162 0.033 4.857
MCK 0.010 0.211 0.042 5.000
TRV 0.010 0.156 0.031 4.978
WM 0.010 0.181 0.034 5.290
$\hat{\mathcal{I}}$ 6.793
Table A9: Uniform test for the existence of jump effects. This table reports the selected threshold locations, the estimated coefficients, the standard errors, and the corresponding individual test statistic. Only the significant results at the $1\%$ level are reported. The last line reports the test statistic of the uniform testing procedure, with critical values $4.610$ at $1\%$ level.
F.2 Application II: Party Incumbency Effects on U.S. House Elections
Year $\hat{\gamma}_{j}$ $(T_{j}b_{j})^{-1/2}\hat{v}_{j}$ $(T_{j}b_{j})^{1/2}\hat{\gamma}_{j}/\hat{v}_{j}$ Obs EObs
1978 0.174 0.043 4.103 412 127
1984 0.163 0.030 5.515 413 154
1986 0.098 0.041 2.418 408 86
1988 0.161 0.042 3.830 395 79
1994 0.117 0.035 3.325 387 102
1996 0.037 0.034 1.087 394 112
1998 0.112 0.035 3.167 369 132
2004 0.045 0.066 0.676 359 65
2006 0.063 0.052 1.222 403 102
2008 0.063 0.029 2.179 405 142
2014 0.075 0.026 2.912 380 103
2016 0.036 0.040 0.902 387 42
2018 0.037 0.032 1.171 403 37
$\hat{\mathcal{I}}$ 5.515
Table A10: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations (Obs) and the effective number of observations (EObs). The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.406$, $2.657$ and $3.166$, respectively.
Year $\hat{\gamma}_{j}-\hat{\gamma}_{0.5}$ $(T_{j}b_{j})^{-1/2}\hat{\tilde{v}}_{j}$ $(T_{j}b_{j})^{1/2}(\hat{\gamma}_{j}-\hat{\gamma}_{0.5})/\hat{\tilde{v}}_{j}$ Obs EObs
1978 0.100 0.041 2.451 412 127
1984 0.089 0.029 3.011 413 154
1986 0.024 0.039 0.606 408 86
1988 0.086 0.040 2.140 395 79
1994 0.042 0.034 1.229 387 102
1996 –0.038 0.033 –1.146 394 112
1998 0.037 0.034 1.078 369 132
2004 –0.030 0.062 –0.481 359 65
2006 –0.011 0.049 –0.234 403 102
2008 –0.012 0.029 –0.414 405 142
2014 0.000 0.026 0.000 380 103
2016 –0.038 0.039 –0.986 387 42
2018 –0.038 0.031 –1.198 403 37
$\hat{\mathcal{Q}}$ 3.011
Table A11: Uniform test for heterogeneity. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure. Critical values at $10\%$, $5\%$ and $1\%$ level are $2.406$, $2.657$ and $3.166$.
State State_ic $\hat{\gamma}_{j}$ $(T_{j}b_{j})^{-1/2}\hat{v}_{j}$ $(T_{j}b_{j})^{1/2}\hat{\gamma}_{j}/\hat{v}_{j}$ Obs EObs
CONNECTICUT 1 0.142 0.048 2.949 72 41
MASSACHUSETTS 3 0.251 0.184 1.362 129 22
NEW JERSEY 12 0.092 0.047 1.953 171 55
NEW YORK 13 0.053 0.026 2.032 387 153
PENNSYLVANIA 14 0.096 0.036 2.699 237 100
ILLINOIS 21 0.114 0.043 2.658 250 85
INDIANA 22 0.080 0.031 2.546 125 62
MICHIGAN 23 0.067 0.032 2.086 206 79
OHIO 24 0.133 0.047 2.815 230 69
WISCONSIN 25 0.128 0.087 1.472 97 31
IOWA 31 0.125 0.038 3.247 64 50
MISSOURI 34 0.258 0.050 5.117 115 37
VIRGINIA 40 0.223 0.069 3.209 102 54
ALABAMA 41 0.212 0.068 3.124 74 21
FLORIDA 43 0.039 0.066 0.593 221 85
GEORGIA 44 0.101 0.075 1.340 126 40
LOUISIANA 45 –0.403 0.386 –1.044 80 16
MISSISSIPPI 46 0.101 0.121 0.836 51 13
NORTH CAROLINA 47 0.027 0.039 0.684 147 77
SOUTH CAROLINA 48 0.352 0.106 3.324 65 25
TEXAS 49 0.016 0.060 0.271 344 117
KENTUCKY 51 0.067 0.086 0.786 70 33
MARYLAND 52 0.126 0.096 1.314 105 29
TENNESSEE 54 0.049 0.128 0.381 101 20
ARIZONA 61 0.131 0.071 1.836 75 34
COLORADO 62 0.070 0.053 1.322 78 44
CALIFORNIA 71 0.088 0.027 3.252 639 189
OREGON 72 0.046 0.042 1.086 64 33
WASHINGTON 73 0.036 0.037 0.981 114 55
$\hat{\mathcal{I}}$ 5.117
Table A12: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.685$, $2.917$ and $3.392,$ respectively. Due to the low effective sample size, the bandwidth is chosen uniformly the same over all states.
Figure F.4: Local linear fit of a significant threshold effect for California (left panel) and an insignificant effect for Pennsylvania (right panel).
State State_ic $\hat{\gamma}_{j}-\hat{\gamma}_{0.5}$ $(T_{j}b_{j})^{-1/2}\hat{\tilde{v}}_{j}$ $(T_{j}b_{j})^{1/2}(\hat{\gamma}_{j}-\hat{\gamma}_{0.5})/\hat{\tilde{v}}_{j}$ Obs EObs
CONNECTICUT 1 0.046 0.050 0.913 72 41
MASSACHUSETTS 3 0.155 0.179 0.866 129 22
NEW JERSEY 12 –0.004 0.049 –0.077 171 55
NEW YORK 13 –0.043 0.032 –1.347 387 153
PENNSYLVANIA 14 0.000 0.039 0.000 237 100
ILLINOIS 21 0.017 0.045 0.384 250 85
INDIANA 22 –0.016 0.036 –0.453 125 62
MICHIGAN 23 –0.029 0.036 –0.789 206 79
OHIO 24 0.037 0.049 0.744 230 69
WISCONSIN 25 0.031 0.086 0.366 97 31
IOWA 31 0.028 0.042 0.684 64 50
MISSOURI 34 0.161 0.052 3.096 115 37
VIRGINIA 40 0.127 0.070 1.818 102 54
ALABAMA 41 0.116 0.068 1.698 74 21
FLORIDA 43 –0.057 0.066 –0.86 221 85
GEORGIA 44 0.005 0.075 0.064 126 40
LOUISIANA 45 –0.499 0.373 –1.339 80 16
MISSISSIPPI 46 0.005 0.118 0.040 51 13
NORTH CAROLINA 47 –0.069 0.043 –1.627 147 77
SOUTH CAROLINA 48 0.256 0.104 2.463 65 25
TEXAS 49 –0.080 0.061 –1.301 344 117
KENTUCKY 51 –0.029 0.085 –0.342 70 33
MARYLAND 52 0.029 0.094 0.313 105 29
TENNESSEE 54 –0.048 0.125 –0.381 101 20
ARIZONA 61 0.034 0.071 0.484 75 34
COLORADO 62 –0.026 0.055 –0.475 78 44
CALIFORNIA 71 –0.009 0.032 –0.269 639 189
OREGON 72 –0.050 0.045 –1.114 64 33
WASHINGTON 73 –0.060 0.040 –1.496 114 55
$\hat{\mathcal{Q}}$ 3.096
Table A13: Uniform test for heterogeneity. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.685$, $2.917$ and $3.392,$ respectively. Due to the low effective sample size, the bandwidth is chosen uniformly over all states.
Appendix G Addtional Remarks
G.1 Key Intuition on the Dependence Structure
Let $\epsilon_{t}=(\epsilon_{1t},\ldots,\epsilon_{Nt})^{\top}.$ The critical question is how the dependency structure in $\epsilon_{t}$ affects the construction of our test statistics. In Section 3 we allow $\left\{\epsilon_{jt}\right\}$ to exhibit some fairly general spatial and temporal dependency structures. One example is a moving average type of model along the $t$-dimension, i.e., $\epsilon_{t}=A_{0}\eta_{t}+A_{1}\eta_{t-1},$ where $A_{0}$ and $A_{1}$ are $N\times N$ matrices, $\eta_{t}=\left(\eta_{1t},\ldots,\eta_{Nt}\right)^{\prime}$ with $\eta_{jt}$’s being i.i.d. across $\left(j,t\right).$ If both $A_{0}$ and $A_{1}$ are diagonal, there is autocorrelation over $t$ but no cross-sectional correlation over $j$. But we allow for general cross-sectional dependence among $\epsilon_{jt}$’s so that neither $A_{0}$nor $A_{1}$ is restricted to be diagonal. Another example is an autoregressive structure along the $t$-dimension: $\epsilon_{t}=A_{0}\epsilon_{t-1}+\eta_{t},$ where, again, $A_{0}$ may or may not be diagonal.
Suppose that there is a cross-sectional dependence in the error term $\epsilon_{jt}$. Then the variance-covariance (VC) matrix of $\epsilon_{t}$ is non-diagonal. Does this have any implication on the choice of the critical value for the simultaneous inferences? In the usual GA setup, a simulated VC matrix is needed; see, e.g., Chernozhukov et al., (2019). Nevertheless, in our algorithm, we do not see any attempt to estimate the covariance of $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$, even though general cross-sectional dependency structures are allowed in $\epsilon_{t}$. This is mainly due to the localized nature of our kernel-type estimators $\hat{\gamma}_{j}$’s. In fact, we can show that under some mild conditions, as long as the joint density of $X_{it}$ and $X_{jt}$ is not degenerate for any $\left(i,j\right)$ pair (see Assumption 3.3 in Section 3), the covariance between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ will be smaller order compared to the variance. Therefore it is fine to standardize just by the standard deviation of $\hat{\gamma}_{j}$ or its consistent estimate without worrying about the covariance.
To understand why the localized statistics have the above favorable properties, we provide some intuitions by focusing on a simplified setting, where both the confounding variable $U_{jt}$ and the individual/group fixed effect $\alpha_{j}$ are absent from the structural model in (1.1), and $\sigma_{j}(X_{jt},U_{jt})=1.$ In this case, $e_{jt}=\epsilon_{jt}.$ For simplicity, we also assume for the moment that $\left(X_{it},\epsilon_{it}\right)$ is i.i.d. over time $t,$ and $X_{it}$’s are independent of $\epsilon_{it}$’s. Heuristically, the weighted statistic $\mbox{max}_{j}(bT)^{1/2}\sum_{t=1}^{T}(w_{jt,b}^{+}\epsilon_{jt}-w_{jt,b}^{-}\epsilon_{jt})$, is of the same order as a weighted statistic adopting a uniform kernel, viz., $\mbox{max}_{j}(bT)^{-1/2}$ $\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{jt}$. The high dimensional GA ensures that we can approximate $\mbox{max}_{j}(bT)^{-1/2}$ $\sum_{t=1}^{T}$ $\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{jt}$ by a maximum of Gaussian random vector with the same VC structure as $\{(bT)^{-1/2}$ $\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{jt}\}_{j=1}^{N}$. Thus the population counterpart of the covariance function is
$(Tb)^{-2}\sum_{t=1}^{T}\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}\epsilon_{it}\epsilon_{jt}]=(Tb)^{-2}\sum_{t=1}^{T}\mathbb{E}(\epsilon_{it}\epsilon_{jt})\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}].$
We focus on the case with $\mathbb{E}(\epsilon_{it}\epsilon_{jt})\neq 0$ as cross-sectional dependence is our main concern here. Obviously, $\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}]=\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}]=O(b)$ if $i=j$, and $\mathbb{E}[\mathbf{1}_{\{|X_{it}|\leq b\}}\mathbf{1}_{\{|X_{jt}|\leq b\}}]=O(b^{2})$ otherwise. Therefore the $i\neq j$ covariance terms are of smaller order than the $i=j$ variance terms and can be ignored asymptotically. In general, such order calculations hold true even when there exists strong cross-sectional correlation among $\left\{(X_{jt},\epsilon_{jt})\right\}_{j=1}^{N}$ as long as we do not have perfect correlation between $X_{it}$ and $X_{jt}$. So we will exclude the case of linear dependence, e.g., $X_{jt}=a_{0}+a_{1}X_{it}$ for some constants $a_{0}$ and $a_{1}\neq 0.$ We will also exclude the case of asymptotic linear dependence: $X_{jt}=a_{1}X_{it}+\eta_{it}$ with Var$(\eta_{it})\rightarrow 0$, which means that the variable $X_{it}$ and $X_{jt}$ are asymptotically linearly dependent. In our simulation study in Section 4, we can even allow strong cross-sectional dependence in $\left\{(X_{jt},\epsilon_{jt})\right\}_{j=1}^{N}$ that is generated via factor structures. As expected, as the cross-sectional dependence becomes stronger, we may lose some information in comparison with the independent case and a large sample size may be needed to achieve the same level of finite sample accuracy as the independent case.
The above explanation is in line with the kernel estimation in the time series literature. It is well known in the nonparametric time series literature that the asymptotic variance of the kernel estimator shares the same asymptotic variance as in the independent case, which means that the covariance terms in the variance calculation are asymptotically negligible under some weak dependence condition. See, e.g., Theorem 5.1 in Fan and Yao, (2003) and Theorems 18.2 and 18.4 in Li and Racine, (2007). Comparatively, our settings are slightly different, since we do not sum over the cross-sectional dimension $j$ and therefore we do not need to impose decay of dependence in the cross-sectional dimension. Even so, we see the covariance between $\hat{\gamma}_{i}$ and $\hat{\gamma}_{j}$ is strictly dominated by the variance of $\hat{\gamma}_{j},$ regardless of the strong or weak cross-sectional dependence structure among $\left\{(X_{jt},\epsilon_{jt})\right\}_{j=1}^{N}.$
G.2 Remarks on Additional Covariates and Fixed Effects
Remark 5.
(Additive time and individual effects) Our model can account for both individual fixed effects $\alpha_{j}$ and time fixed effects $\nu_{t}$. Specifically, the working model (2.1) becomes
$Y_{jt}=\alpha_{j}+\nu_{t}+f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}.$ (SA-1)
If $\mathbb{E}(\nu_{t}|X_{jt})=d(X_{jt})$ for some function $d(\cdot)$ under a covariance-stationarity condition along the $t$-dimension, then by defining $\tilde{e}_{jt}=e_{jt}+[\nu_{t}-\mathbb{E}\left(\nu_{t}|X_{jt}\right)]$ and $\tilde{f}_{j}(X_{jt})=\alpha_{j}+\mathbb{E}\left(\nu_{t}|X_{jt}\right)+f_{j}(X_{jt}),$ the above model can be reformulated as follows:
$Y_{jt}=\tilde{f}_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\tilde{e}_{jt}.$ (SA-2)
In the absence of covariance-stationarity, the estimation would become much more involved and we have to leave it for future research.            $\Box$
Remark 6.
(Adding more covariates to fit the threshold model) Suppose that we want to include more covariates $Z_{jt}\in\mathbb{R}^{d}$ and $\theta\in\mathbb{R}^{d}$, we can impose a partial linear structure. Then the main model in (2.1) can be of the following form:
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}+Z_{jt}^{\top}\theta.$ (SA-3)
Our methodology can be extended to the model above by adapting the approach used in Li et al., (2022).            $\Box$
G.3 Details on the Comparison
In this section, we match our model with its parametric counterpart. Recall our full model in (1.1), consider a more general case
$Y_{jt}=h_{j}(X_{jt},U_{jt})+\tau_{j}(X_{jt},U_{jt})\mathbf{1}_{\{Q_{jt}\geq c_{0j}\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt},$ (SA-4)
where $Q_{jt}$ is some threshold variable. Suppose $Q_{jt}$ is a function of covariate vector $(X_{jt},U_{jt}^{\top})^{\top}$, namely $Q_{jt}=T_{Q}(X_{jt},U_{jt})$. We can rewrite $\tau_{j}(X_{jt},U_{jt})=\tau_{j}^{*}(Q_{jt},V_{jt})$ with $V_{jt}$ uncorrelated with $Q_{jt}$. Thus it suffices to consider our model (1.1). In another totally special case, if $Q_{jt}$ is independent of $X_{jt}$ and $U_{jt}$, we shall fit nonparametric functions $\tau_{j}(\cdot,\cdot)$ and $h_{j}(\cdot,\cdot)$ with separablely samples according to the threshold $Q_{jt}$.
The linear version of the above model without linearize $\sigma_{j}(\cdot,\cdot)$ is
$Y_{jt}=X_{jt}\beta_{j,1}+U_{jt}^{\top}\theta_{j,1}+[X_{jt}\beta_{j,2}+U_{jt}^{\top}\theta_{j,2}]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt}.$ (SA-5)
Assume $U_{jt}=\gamma_{j}+X_{jt}\xi_{j}$. Then by definition
$\tilde{\tau}_{j}(X_{jt})=\mathbb{E}[X_{jt}\beta_{j,2}+U_{jt}^{\top}\theta_{j,2}|X_{jt}]=X_{jt}(\beta_{j,2}+\xi_{j})+\gamma_{j}\theta_{j,2}.$
Similarly $\tilde{h}_{j}(X_{jt})=X_{jt}(\beta_{j,1}+\xi_{j})+\gamma_{j}\theta_{j,1}$. Denote $\tilde{\beta}_{j,1}=\beta_{j,1}+\xi_{j}$, $\tilde{\beta}_{j,2}=\beta_{j,2}+\xi_{j}$, $\gamma_{j,1}=\gamma_{j}\theta_{j,1}$ and $\gamma_{j,2}=\gamma_{j}\theta_{j,2}$, then model (SA-5) can be rewritten into
$Y_{jt}=X_{jt}\tilde{\beta}_{j,1}+\gamma_{j,0}+[X_{jt}\tilde{\beta}_{j,2}+\gamma_{j,1}]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}.$ (SA-6)
Then the jump effect is $\gamma_{j,1}$ and the change of derivative corresponding to $\tilde{\beta}_{j,2}$.
For the parametric model, usually in the literature heterogeneity is ignored. And a pooled ordinary least square regression estimator for $\tilde{\beta}_{j,.}$ is considered and used to test the overall heterogeneity effect, see Lee et al., (2011). Thus we shall see that under the null hypothesis, our statistics and a pooled statistics will both have the correct asymptotic distribution. However, under the alternative, the signals can easily be canceled out and leads to a trivial low power setup.